### v1

# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

Abstract
This paper presents a novel framework that leverages the Kullback-Leibler divergence (KLD) to analyze and optimize performance trade-offs in integrated sensing and communication (ISAC) systems. We consider a multiple-input-multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using shared antenna deployment. The proposed KLD-based approach provides a unified performance measure encompassing both UE error rate and target detection capability. We apply this approach on two well-known communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), and evaluate the effect on the radar subsystem. Furthermore, two optimization problems are formulated and solved. The first one optimizes the KLD of the radar subsystem for given constraints on the communication KLD, whereas the second one focuses on communication waveform KLD-based optimization and constrained radar KLD. These optimization problems are solved using a projected gradient method with an adaptive penalty for the radar waveform and a gradient-assisted interior point method for the communication waveform. Through theoretical derivations and extensive simulations, it is demonstrated that our approach can be a powerful tool for characterizing and optimizing the performance trade-offs of ISAC under various configurations. The results also show significant improvements in both sensing and communication performance by the KLD-optimized system compared to well-known benchmarks, such as conventional MRT and ZF for the communication subsystem, and the conventional identity covariance design for the radar subsystem. These findings support the holistic design and optimization of ISAC in next-generation wireless networks.
Index Terms:
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission (MRT), beamforming, optimization, Kullback–Leibler divergence (KLD).
I Introduction
The rapid evolution of wireless communication networks, particularly with the advent of fifth-generation (5G) technology, has dramatically expanded the landscape of interconnected devices and services. This technological leap has enabled a wide array of applications that depend heavily on advanced sensing capabilities. From autonomous vehicles and unmanned aerial vehicles (UAVs) to the Internet-of-Things (IoT) ecosystems, these applications are pushing the boundaries of traditional communication systems, necessitating the integration of sophisticated sensing functionalities [1, 2, 3, 4]. As we stand on the threshold of the sixth-generation (6G) era, network operators and service providers face unprecedented challenges in not only delivering robust communication services but also in integrating advanced sensing-based applications such as detection, localization, tracking, navigation, and environmental surveillance [5, 6, 7].
Integrated Sensing and Communications (ISAC) systems have garnered significant interest in the research and industrial communities to meet these emerging demands. ISAC systems aim to synergistically merge sensing and communication functionalities within the same infrastructure, thereby optimizing the use of available resources at base stations (BSs) for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various architectural approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilize resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, potentially offering higher resource efficiency at the cost of the increased complexity and the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC systems has been accelerated by the advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems. These systems allow for the generation of highly directional beams and support simultaneous service to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy, usually done by exploiting spatial diversity by setting the antenna separation to a minimum of half wavelength [16]. The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate, while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimization of ISAC systems.
I-A Literature review
To address these challenges mentioned above, there has been a growing interest in developing unified performance measures that can simultaneously capture both sensing and communication capabilities. One such measure that has gained traction is the Kullback-Leibler divergence (KLD), also known as the relative entropy. KLD has a well-established history in information theory and has been extensively used in the analysis and design of sensing systems, particularly in the context of MIMO radars [18, 19, 20]. It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another. While the application of KLD in sensing systems is well-established, its potential in characterizing communication system performance remains open. Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems. For instance, in our previous works [21, 22], we employed KLD to analyze performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimizing network resources for both separated and shared deployment scenarios. Moreover, further insights about the achievable KLD trade-off in ISAC systems is explored in [24], and more designs have been discovered. Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalized multi-user-multi-target (MUMT) ISAC system using the promising KLD.
Further efforts have explored various optimization approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimization based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25]. This approach minimizes the CRB of angle estimation for radar while maximizing the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously. However, these methods often focus on optimizing individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimization has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26]. While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in scenarios where these functionalities are deeply integrated. Waveform optimization has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27]. This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics. However, this work does not consider a unified optimization framework that could simultaneously address the needs of both radar and communication subsystems.
I-B Motivation and contribution
Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems. Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilized for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities and maximises the system’s overall resource efficiency [14]. This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterize and optimize the sensing-communication performance trade-offs in shared antenna deployment ISAC systems. This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems. By considering the shared deployment scenario and proposing specific optimization techniques for both radar and communication waveforms, we provide a comprehensive approach for system optimization that reflects the true potential of ISAC systems in future wireless networks. Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems that can effectively balance the needs of both radar and communication functionalities. Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources. The key contributions of our work are as follows:
  • Unified Performance Measure: We derive the Kullback-Leibler Divergence (KLD) specifically for the integrated operation of communication and radar systems within ISAC. This performance measure unifies the communication error rate with radar target detection capabilities into a single objective function. By deriving this KLD-based metric for communication and radar components, we optimise ISAC system design holistically, addressing the trade-offs and enhancing overall system performance.
  • Inherent Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilizing communication interference as an additional source of information for target detection.
  • KLD-based Optimized ISAC: We propose two novel optimization techniques: radar and communication waveform KLD-based optimization. Both maximize their respective KLD measures subject to the minimum KLD of both subsystems and power constraints. The non-convex radar optimization employs a projected gradient method with an adaptive penalty, while the communication optimization uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimization across both subsystems within system constraints.
  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyze their impact on the overall system performance within our KLD-based framework. This analysis provides practical insights for system designers seeking to optimize ISAC performance in shared deployment scenarios with a complexity-performance trade-off.
Our results demonstrate that the proposed KLD-based optimization techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design. The radar waveform optimization shows substantial improvements in both target detection and communication performance, while the communication waveform optimization primarily benefits the communication subsystem with modest radar performance gains. Both techniques exhibit robust performance across varying SNR levels, with the radar optimization demonstrating particularly stable computational efficiency. Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources. These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
I-C Paper organization
The paper is structured as follows. Section II presents the system model. Sections III and IV analyze the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated. Section VI introduces the second KLD-based optimisation technique based on communication KLD sum with radar KLD constraint. Section VII displays the numerical results, and finally, Section VIII concludes the work.
$Notation$: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors. Superscripts ${(\cdot)}^{*}$, ${(\cdot)}^{T}$, and ${(\cdot)}^{H}$, denote the conjugate, transpose, and Hermitian transpose, respectively. Subscripts ${(\cdot)}_{\mathrm{c}}$ and ${(\cdot)}_{\mathrm{r}}$ relate to the communication, and radar subsystems, respectively.
II System Model
This paper is concerned in an ISAC system that comprises an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication user equipments (UEs) in the downlink direction using the shared deployment scenario. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The portions of the power that are allocated to the radar and communication subsystems are respectively denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimized systems, which will be introduced and analyzed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},$ (1)
where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. Notably, the transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received version of $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might impose interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal for sensing targets as well. This dual-use approach enhances radar performance by utilizing the communication signal as an additional source of information. Moreover, the radar waveform design in our work is based on the combined transmit signal covariance matrix for the $t$-th beam such that the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, where $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
II-A Communication System
At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a certain normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$. The received signal at the $k$-th UE can be formulated as follows,
$y_{\mathrm{c},k,l}=\overbrace{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}}^{\text{Desired $k$-th UE signal}}+\overbrace{\omega_{\mathrm{MN},k}}^{\begin{subarray}{c}\text{Inter-user}\\ \text{interference}\end{subarray}}+\overbrace{\eta_{k}}^{\begin{subarray}{c}\text{Radar interference}\\ \text{and noise}\end{subarray}},$ (2)
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, the variable $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ represents the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent, and $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ id the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. In the next two subsections, we revisit the conventional MRT and ZF designs which are typically used to precode the communication data symbols at the BS. These precoding schemes will serve as benchmarks for our designs.
II-A1 MRT beamforming
In MRT precoding, the amount of received power at the UEs is maximized without taking into account the imposed interference. This approach usually lead to acceptable performance at low SNR values but its performance deteriorates at large SNRs. The normalized MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\>\mathbf{H}^{*}/\|\mathbf{H}\|_{F}=\mathbf{P}\>[\mathbf{h}_{1}^{*},\ldots,\mathbf{h}_{K}^{*}]/\|\mathbf{H}\|_{F},$ (3)
where $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
II-A2 ZF beamforming
On the contrary, ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalized ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (4)
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$ is the ZF beamformer before normalisation.
II-B Radar system
The radar system proposed in this study is designed for high adaptability, allowing for real-time adjustments in detection capabilities on a per-frame basis, where each frame is composed of $L$ snapshots. A key parameter of this system is $T$, representing the count of radar beams emitted during a specific detection frame. By harnessing MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [30, 31]. Consequently, $T$ signifies the upper limit of detectable targets within a single frame.
This radar system’s design is robust, as it can explore multiple angular-range-Doppler bins over successive frames, thereby increasing the chance of identifying more targets. Furthermore, its flexibility in varying the number of emitted beams allows for the simultaneous detection of a larger set of targets, providing the ability to fine-tune the system’s detection capability according to different operational demands. However, it is crucial to acknowledge that the value of $T$ is naturally constrained by the available number of antennas and the total number of UEs being serviced by the BS, highlighting the critical need for efficient waveform design, and resource allocation strategies in ISAC systems.
Our focus aligns with scenarios where targets are spatially separated, ensuring each target is confined to a distinct radar bin, as explored in prior studies [32, 33]. Additionally, the literature offers a range of algorithms developed to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34]. A straightforward method to count detected targets involves applying the detection mechanism across all radar bins spanning the angular, range, and Doppler dimensions. The total radar return signal at the $l$-th snapshot is shown as follows,
$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},$ (5)
where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix of the BS-Target-BS route for the $t$-th target, following a Rayleigh distribution [35]. The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot. To analyze the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin. Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target. The radar precoding matrix $\mathbf{W}_{r,l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$. The received signal for the $t$-th target at the BS, under the binary hypotheses $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ signifies target presence and $\mathcal{H}_{0}$ denotes target absence, can be mathematically represented as follows,
$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}\>q+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}\>q+\mathbf{n}_{t,l},$ (6)
The $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
III The KLD of communication system: ZF and MRT precoding
In this section, we derive the KLD of the communication system for two typical scenarios, namely, the MRT, and the ZF beamforming schemes using the shared deployment antenna configuration. The analysis is carried out for the normalized precoders introduced in Section II-A.
In order to, derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise components. For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right),$ (7)
The application of this formula to the communication system requires the determination of the mean vectors and covariance matrices for each precoding scheme.
III-A MRT beamforming
For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components. The examination starts with the IUI term, which is the second term in (2), which can be written as,
$\omega_{\mathrm{MN},k}=d_{\mathrm{c},k}^{-\zeta/2}\mathbf{h}_{k}^{T}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}=\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\|\mathbf{H}\|_{F}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (8)
Since that the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d. complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors. Therefore, central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a significant number of antennas $N$. On the other hand, the normalization term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$. Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable. Unfortunately, this ratio distribution does not have a simple closed-form expression. However, it is worth noting that we can approximate its behavior for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$. Consequently, using this approximation, we can write (8) as,
$\omega_{\mathrm{MN},k}\approx\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\sqrt{NK\sigma_{h}^{2}}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (9)
Now, $\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking the CLT, for large $KN$, this sum approaches a complex normal distribution with zero mean and a variance of,
$\sigma_{\omega,\mathrm{MN}}^{2}\approx\frac{d_{\mathrm{c},k}^{-\zeta}}{NK\sigma_{h}^{2}}\sum_{i\neq k}N\sigma_{h}^{4}p_{i}=\frac{d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}}{K}\sum_{i\neq k}p_{i},$ (10)
The accuracy of this approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation and shown in Fig.1, which compares it with simulated density functions under different values of $N$. It can be also shown that the distribution of $\eta_{k}$ can be well approximated by a complex Gaussian distribution using the CLT, particularly for large $N$. This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (11)
where $\mathbf{R}_{w,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{\tilde{w}}_{\mathrm{r},t,l}\mathbf{\tilde{w}}_{\mathrm{r},t,l}^{H}$. Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$. In order to find the KLD for the MRT case with matrix normalization, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2). This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
$|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}=|\mathbf{h}_{k}^{T}\mathbf{h}_{k}^{*}|^{2}/\|\mathbf{H}\|_{F}^{2}=\|\mathbf{h}_{k}\|^{4}/\|\mathbf{H}\|_{F}^{2}.$ (12)
The expected value of this term can be approximated using the properties of the chi-squared distribution and the approximation $\mathbb{E}[1/X]\approx 1/\mathbb{E}[X]$ for a random variable $X$,
$\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]\approx p_{k}\mathbb{E}[\|\mathbf{h}_{k}\|^{4}]/\mathbb{E}[\|\mathbf{H}\|_{F}^{2}]=p_{k}(N+1)\sigma_{h}^{2}/K.$ (13)
After deriving the statistical properties of the signal and interference components, the KLD formula can be applied. Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, these can be substituted into equation (7), then the KLD can be found as [21, Corollary 1],
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]}{(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2},$ (14)
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ which depends on the transmitted data constellation, and thus $\lambda$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$. Finally, substituting (13) in (14), then the KLD for MRT precoding with matrix normalization can be found as,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N+1)\sigma_{h}^{2}}{K\>(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2}.$ (15)
Figure 1: The density functions of the approximated random variable $\omega_{MN}$.
III-B ZF beamforming
With ZF precoding, the received signal at the $k$-th user can be rewritten as follows,
$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l},$ (16)
where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalized ZF precoding matrix. It should be noticed that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder completely eliminates the IUI. The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are shown as follows,
$\boldsymbol{\mu}_{k,j}$ $=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]s_{\mathrm{c},k,l}^{(j)}$
$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[|\mathbf{\tilde{W}}_{\mathrm{c}}|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (17)
$\ \Sigma_{n}$ $=\Sigma_{m}=\sigma_{\eta}^{2},$ (18)
where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalization of the ZF precoding matrix. Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, it can be shown that $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$. Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$. By substituting (17), and (18) in (7), then the KLD can be reduced to,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}}{M(M-1)\sigma_{\eta}^{2}\ln 2}\mathbb{E}\left[\frac{1}{\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}}\right],$ (19)
where the expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
[…]
Substituting this result into the expression for $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$ in (19), the final expression for the average KLD in the case of ZF precoder with matrix normalization can be reduced to,
$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N-K)/M(M-1)\sigma_{\eta}^{2}\ln 2.$ (21)
III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the bamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimization problem for which $\mathbf{W}_{\mathrm{c}}$ can be designed based on the channel side information matrix. Accordingly, by considering the received signal representation in (2), the conditional mean vectors and variance can be obtained as follows,
$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\$ (22)
$\operatorname{Var}(y_{\mathrm{c},k,l}^{(j)})$ $=\mathbb{E}[|y_{\mathrm{c},k,l}^{(j)}-\boldsymbol{\mu}_{k,n}|^{2}]$
$=d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},\!\quad j\in\{n,m\},$ (23)
where the assumptions of zero-mean, temporal whiteness, and wide-sense stationarity of the communication and radar signals, as well as their mutual orthogonality, have been utilized, it is important noting that the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Therefore, the covariance is,
$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$ (24)
Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}\times$
$\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ and $\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}=\lambda/(M(M-1))$, substituting it in the KLD expression which becomes,
$\begin{split}\mathrm{KLD}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (26)
It should be noted that the second term in this equation, represents the form of the SINR, where the interference is the sum of IUI and radar interference.
IV Radar system analysis
IV-A Composite detection and response matrix estimation
The main objective for typical radar systems is to detect targets and estimate the matrix response associated with each one. After separating different signals associated with different targets, such problem can be formulated for each target as a composite detection-estimation problem as follows,
$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{q}).$ (27)
This joint optimization problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present. Now, considering the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each of these hypotheses can be shown as a multivariate normal density, which can be written as
$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
×$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (28)
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalized likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses. The GLRT is derived as follows,
$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (29)
However, the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$. Therefore, the target response matrix is typically estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain
$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1}).$ (30)
Since the noise $\mathbf{n}_{t,l}$ is Gaussian and the $\mathbf{H}_{t}$ is normally distributed, the likelihood function is Gaussian, and the maximization leads to the least squares solution. Therefore, this maximization reduces to the following form,
$\hat{\mathbf{H}}_{t}$ $=\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}-\mathbf{H}_{t}\mathbf{x}_{t}\right\|^{2}=\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t}^{H}\left(\mathbf{x}_{t}\mathbf{x}_{t}^{H}\right)^{-1}.$ (31)
This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain the following,
$\Lambda(\mathbf{y}_{\mathrm{r},t,l})=$
$\frac{\pi^{N}\sigma_{n}^{2N}\exp\left(-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\right)}{\pi^{N}\det(\hat{\mathbf{R}}_{1,t})\exp\left(-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\right)}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{0}$ (32)
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold that is a design parameter, for example, according to Neyman-Pearson lemma it can be determined to obtain a certain false alarm rate. Taking the logarithm and rearranging terms, we obtain the following expression, which now clearly includes the estimated target response matrix $\hat{\mathbf{H}}_{t}$,
$\ln\Lambda(\mathbf{y}_{\mathrm{r},t,l})$ $=\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}-N\ln\sigma_{n}^{2}+\ln\det(\hat{\mathbf{R}}_{1,t})$
$\!-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\ln\tau_{0}$ (33)
The terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ are constant with respect to $\mathbf{y}_{\mathrm{r},t,l}$, so they can be absorbed into the threshold, which further simplifies our GLRT as follows,
$\dot{\Lambda}(\mathbf{y}_{\mathrm{r},t,l})=\mathbf{y}_{\mathrm{r},t,l}^{H}\left(\mathbf{I}_{N}+\hat{\mathbf{R}}_{1,t}^{-1}\right)\mathbf{y}_{\mathrm{r},t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{1}$ (34)
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. This final expression correctly reflects the comparison between the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. In order to utilize the whole $L$ snapshots whose received signals are independent, we average as follows,
$\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}.$ (35)
where $\tau_{2}$ is the detection threshold that accounts for all snapshots and could be designed according to the Neyman-Pearson lemma as illustrated above.
IV-B The KLD of radar system
[…]
Leveraging this asymptotic property and utilizing the $L$ number of snapshots, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (28), and (7) as follows,
$\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=$ $\frac{1}{\ln{2}}\Big{(}\ln\left(\mathrm{det}(\mathbf{R}_{2,t})\right)+\mathrm{Tr}\left(\mathbf{R}_{2,t}^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)$
$-N\>\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\Big{)},$ (36)
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. It should be noted that $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
V Radar Waveform Optimisation
Here, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem can be shown as follows,
$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (37a)
s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (37b)
$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (37c)
$\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{r}},$ (37d)
$P_{\mathrm{c}}=P_{\mathrm{T}}-P_{\mathrm{r}},$ (37e)
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (36). After substituting (36) in (37a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimization problem focuses on the radar waveform while considering the whole ISAC system performance. The constraint on radar power (37d) is explicit because the radar waveform is our primary optimization variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (37c) on communication performance. While the communication power is not directly optimized, it is indirectly influenced by the radar power allocation. Specifically, the communication subsystem utilizes its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information. This approach allows us to focus on the interplay between subsystems which is mainly governed by waveform design and interference management while maintaining system-wide performance.
$\mathcal{P}_{2}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}$ $\!\!\!\!\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\ln{2}}\Bigg{(}\ln\left(\mathrm{det}(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N})\right)$ (38)
$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)-N\>\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\Bigg{)}$
s.t. (37b), (37c), (37d), and (37e)
To Tackle the optimisation problem in (38), we utilize the Projected Gradient method with a penalty function [36]. This method is chosen for its suitability to the non-linear, non-convex optimization problem [37]. This approach incorporates complex KLD constraints via the penalty function, while efficiently handles the radar power constraint through projection. It offers a balance between the solution quality and the computational efficiency, crucial for potential real-time ISAC applications. The method’s scalability with problem dimensionality further justifies the selection of this method for the radar waveform optimization task in the ISAC setup illustrated in the problem formulation. The optimization variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor. We define the objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ and a penalty function $p(\mathbf{\dot{W}}_{\mathrm{r}})$ to handle the constraints in (37b) and (37c),
$f(\mathbf{\dot{W}}_{\mathrm{r}})=\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}},$ (39)
$\begin{split}p(\mathbf{\dot{W}}_{\mathrm{r}})=&\sum_{t=1}^{T}\left(\max(0,\underline{A}_{t}-\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}})^{2}\right)\\ &+\sum_{k=1}^{K}\left(\max(0,\underline{B}_{k}-\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}}})^{2}\right).\end{split}$ (40)
The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix I.A. The final equation is given by,
$\begin{split}[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=&\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}\Big{(}\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ &-\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big{)}_{i}.\end{split}$ (41)
The gradient of the penalty function $\nabla p(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$ is also a gradient tensor, computed as,
$\begin{split}[\nabla p(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=&-2(\underline{A}_{t}-\mathrm{KLD}_{\mathrm{r},t})[\nabla\mathrm{KLD}_{\mathrm{r},t}]_{i,l}\cdot\mathbb{I}_{1}\\ &-\sum_{k=1}^{K}2(\underline{B}_{k}-\mathrm{KLD}_{\mathrm{c},k})[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}\cdot\mathbb{I}_{2},\end{split}$ (42)
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{r}}$ and $\mathrm{KLD_{c}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A. The final forms are given by,
$[\nabla\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}]_{i,l}=$ $\frac{2\>d_{\mathrm{r},t}^{\zeta}}{L\ln{2}}\Big{(}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$
$-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big{)}_{i},$ (43)
$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}=$ $-\frac{2\>d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (44)
The update rule for the projected gradient method with penalty ensures that each iteration moves the solution towards optimality while maintaining feasibility, and can be shown as follows,
$\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}+\alpha_{n}\mathbf{G}^{(n)}\right),$ (45)
where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, which balances the importance of the objective function and the constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (37d) is always satisfied. The projection mechanism is shown in the following equation,
$\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}})=\begin{cases}\mathbf{\dot{W}}_{\mathrm{r}}\sqrt{\frac{P_{\mathrm{r}}L}{\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}}},&\text{if }\frac{1}{L}\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}>P_{\mathrm{r}}\\ \mathbf{\dot{W}}_{\mathrm{r}},&\text{otherwise}\end{cases}$ (46)
The step size $\alpha_{n}$ is determined using a backtracking line search to ensure convergence. We find the smallest non-negative integer $m$ such that,
$\begin{split}&f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\geq\\ &f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})+c\alpha_{n}\|\mathbf{G}^{(n)}\|_{F}^{2},\end{split}$ (47)
where $c\in(0,1)$. This condition, known as the Armijo condition, ensures a sufficient decrease in the penalized objective function [38]. The step size is then set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, where $\beta\in(0,1)$ is a reduction factor. The penalty parameter $\rho_{n}$ is updated as follows,
$\rho_{n+1}=\begin{cases}\gamma\rho_{n},&\text{if }p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})>0\\ \rho_{n},&\text{otherwise}\end{cases}$ (48)
where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied.
Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimization
0:  Initial point $\mathbf{\dot{W}}_{\mathrm{r}}^{(0)}$, constants $\alpha_{0}$, $\rho_{0}$, $\beta$, $\gamma$, $c$, $P_{\mathrm{r}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\varepsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{\dot{W}}_{\mathrm{r}}^{*}$
1:  Initialize $n=0$
2:  while $n<\mathrm{max\_iter}$ do
3:     Compute gradients $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ and $\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$
4:     Compute total gradient $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$
5:     Perform backtracking line search to find $\alpha_{n}$
6:     Update $\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}=\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}+\alpha_{n}\mathbf{G}^{(n)})$
7:     Compute $f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})$ and $p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})$
8:     Update $\rho_{n+1}$ according to the penalty update rule
9:     if $\|\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\|_{F}<\varepsilon$ then
10:        break
11:     end if
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{\dot{W}}_{\mathrm{r}}^{*}=\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}$
Algorithm 1 implements the projected gradient method with penalty for the radar waveform optimization problem. It iteratively updates the waveform $\mathbf{\dot{W}}_{\mathrm{r}}$ by moving in the direction of the total gradient $\mathbf{G}^{(n)}$, which combines the gradients of the objective function gradient and the penalty function. The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. This projection is crucial as it maintains the feasibility of the solution with respect to the power constraint throughout the optimization process. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
$\begin{split}\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\leq&\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})\\ &+c\alpha_{n}\langle\nabla\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}),\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\end{split}$ (49)
where $\phi(\mathbf{\dot{W}}_{\mathrm{r}})=f(\mathbf{\dot{W}}_{\mathrm{r}})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}})$ is the penalized objective function. The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation. This adaptive scheme allows the algorithm to balance between optimizing the objective function and satisfying the constraints. As the optimization progresses, if constraints are consistently violated, the increasing penalty parameter will put more emphasis on constraint satisfaction.
VI Communication optimisation
The communication subsystem optimization employs the gradient-assisted interior point method [39], contrasting with the projected gradient method with penalty used for radar optimization, due to the fundamental difference in optimization variable structure. While radar optimization involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, communication beamforming optimizes over a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions for the matrix optimization problem. The communication waveform is optimised while using conventional identity covariance matrix design for the radar subsystem with a covariance matrix set to $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be shown as follows,
$\mathcal{P}_{3}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (50a)
s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (50b)
$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (50c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (50d)
$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.$ (50e)
It is important to note that the radar power is uniformly distributed among the potential targets, as the radar covariance matrix ensures an omnidirectional beam flow. To solve this optimization problem, we need to derive the gradient of the objective function and the constraints. Let’s define the objective function as,
$f(\mathbf{W}_{\mathrm{c}})=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}.$ (51)
Now, to compute the gradient of the objective function, we need to derive the gradients of each objective function $\mathbf{W}_{\mathrm{c}}$ components.Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$. Using the chain rule and the properties of matrix derivatives, the gradients can be expressed as,
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\quad\text{for all }j$ (52)
$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ =$=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (53)
where $C=\frac{\lambda}{2M(M-1)\ln 2}$, and $b_{k}=d_{\mathrm{c},k}^{\zeta}\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$. Finally, we can express the gradient of the objective function as,
$\begin{split}\nabla_{\mathbf{w}_{\mathrm{c},k}}f(\mathbf{W}_{\mathrm{c}})=\frac{d_{\mathrm{c},k}^{\zeta}C}{K}\sum_{k=1}^{K}\Bigg{(}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}}\\ -\sum_{j\neq k}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}}\Bigg{)}.\end{split}$ (54)
Next, we need to consider the constraints of our optimization problem. The power constraint is given by,
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.$ (55)
The gradient of this constraint is derived straightforwardly as follows,
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.$ (56)
The gradient of the radar KLD constraint with respect to $\mathbf{W}_{\mathrm{c}}$ is derived as follows,
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).$ (57)
Algorithm 2 Interior Point Method for Communication Optimization
0:  Initial point $\mathbf{W}_{c}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{W}_{c}^{*}$
1:  Initialize $n=0$, $\mu=\mu_{0}$
2:  while $n<\mathrm{max\_iter}$ do
3:     Compute $f(\mathbf{W}_{c}^{(n)})$, $\nabla f(\mathbf{W}_{c}^{(n)})$, $g_{i}(\mathbf{W}_{c}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{c}^{(n)})$
4:     Compute KKT conditions
5:     if KKT conditions satisfied within $\epsilon$ then
6:        break
7:     end if
8:     Solve Newton system for search direction $\Delta\mathbf{W}_{c}$
9:     Perform line search to find step size $\alpha$
10:     Update $\mathbf{W}_{c}^{(n+1)}=\mathbf{W}_{c}^{(n)}+\alpha\Delta\mathbf{W}_{c}$
11:     Update $\mu=\gamma\mu$
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{W}_{c}^{*}=\mathbf{W}_{c}^{(n)}$
Solving the optimization problem defined in equations (50a)-(50d) involves employing the Interior Point Method (IPM) as outlined in Algorithm 2. This method efficiently handles constrained optimization using key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions, and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is reduced as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This process continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximizes average communication KLD while satisfying radar and power constraints.
VII Numerical Results
In this section, we present the simulation results for both our conventional benchmarks and the two optimisation techniques introduced previously. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The total number of antennas at BS is $N=20$, a number of $L=100$ snapshots, the maximum number of potential radar targets is $T=3$, and the number of UEs is $K=3$. The pathloss exponent is $\eta=3$ to model the effect of large-scale fading, and the distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters, respectively, and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters, respectively. The channel variance is $\sigma_{h}^{2}=1$.
VII-A Trade-off for the conventional benchmarks
The conventional identity covariance matrix design is used for the radar subsystem with a covariance matrix is set to $\mathbf{R}_{w}=\mathbf{I}_{N}$, whereas the communication subsystem uses MRT and ZF precoding introduced earlier.
Figure 2: The performance of the communication subsystem for both beamforming schemes, where the Bit Error Rate (BER) and the achiveable $\mathrm{KLD}_{\mathrm{c}}$ are shown versus $P_{\mathrm{R}}/{N}_{o}$, with $P_{\mathrm{r}}=0.5$.
Fig. 2 presents the performance of the communication subsystem for both ZF and MRT beamforming schemes. In Fig. 2.a, the MRT scheme’s performance is depicted, with the KLD performance on the right $y$-axis and the BER performance on the left. As can be observed from the figure, the KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ until around $10$ dB, where a noticeable upper-bound occurs primarily caused by radar interference, with IUI contributing to a less extent. Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, indicating an improvement in the communication subsystem performance, but also exhibits the same error floor at around $10$ dB. Fig. 2.b illustrates the ZF scheme’s performance, exhibiting similar trends in the KLD and BER curves. However, compared to the MRT scheme, the ZF scheme shows an increase in KLD values and a decrease in BER values for the same received SNR, indicating that the primary source of interference is indeed caused by radar interference as IUI is cancelled using ZF precoding. Consequently, both the ZF and MRT schemes exhibit a drastic error floor in BER, or equivalently an upper bound for KLD, due to the unmitigated radar interference at $P_{\mathrm{c}}=0.5$. Moreover, this figure clearly shows the effectiveness of using KLD to characterize the communication system.
Figure 3: The receiver operating characteristic (ROC) curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
Fig. 3 shows the receiver operating characteristic (ROC) curves depicting the performance of the radar subsystem for the three potential targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios, for the same ISAC system in Fig. 2 with ZF precoding for communication subsystem assumed. The ROC curves are obtained by plotting the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target. In Fig. 3, where $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for all three targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are relatively close to each other, indicating similar detection performance at low signal-to-noise ratios. In Fig. 3, with $P_{\mathrm{R}}/{N}_{o}=10$ dB, the ROC curves for the three targets start to diverge, where $\mathrm{T}_{3}$ exhibits the best performance, followed by $\mathrm{T}_{1}$ followed by $\mathrm{T}_{2}$. This suggests that as the signal-to-noise ratio increases, the radar subsystem’s ability to detect the targets and distinguish between them improves. Finally, in Fig. 3, in which $P_{\mathrm{R}}/{N}_{o}=20$ dB, the separation between the ROC curves becomes even more pronounced. As can be observed in 3, at $P_{FA}=10^{-3}$, $P_{D}$ is $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, where $\mathrm{T}_{3}$ maintains its superior performance, while $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant improvements in their detection probabilities compared to the lower $P_{\mathrm{R}}/{N}_{o}$ scenarios. This demonstrates that at high signal-to-noise ratios, the radar subsystem can effectively detect and differentiate between the three targets, with $\mathrm{T}_{3}$ being the most easily detectable as it is the closest target to the BS.
Figure 4: a) The performance of the radar subsystem where the achievable $\mathrm{KLD_{\mathrm{r}}}$ is shown versus $P_{\mathrm{R}}/{N}_{o}$, across different values of $P_{\mathrm{r}}$. b) The KLD performance trade-off between the radar and communication subsystems for ZF and MRT
In Fig.4.a, the performance of the radar subsystem is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus the $P_{\mathrm{R}}/{N}_{o}$ across different values of $P_{\mathrm{r}}$. It is shown that as $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ for each target also increases for all values of $P_{\mathrm{r}}$. This indicates that the radar subsystem’s performance improves with higher $P_{\mathrm{R}}/{N}_{o}$. The performance curves for different $P_{\mathrm{r}}$ values are relatively close to each other, suggesting that the system’s performance is less sensitive to changes in $P_{\mathrm{r}}$ because the communication subsystem signal is also incorporated in the target detection, using our improved system model.
Fig. 4.b presents the trade-off performance between radar and communication subsystems using the ZF and MRT for the communication subsystem and the conventional identity covariance design for the radar subsystem. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation between both subsystems, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value. The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ outperforms the $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ slightly, as both of them exhibits saturation at higher $P_{\mathrm{R}}/N_{o}$ due to the destructive interference from the radar subsystem mainly, while the $\mathrm{KLD_{r}}$ shows a steady increase when moving from left to right on the curve. It can be observed from the figure that a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits can be achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same amount of $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ is achieved at around $25$ dB, which implies an SNR gain of about $13$ dB in favour of ZF.
VII-B Radar Waveform KLD-based Optimisation
The radar waveform KLD-based optimization algorithm is configured with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The penalty parameter is initialized at $\rho_{0}=1$, with a penalty increase parameter of $\gamma=1.5$, and an initial step size of $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
Figure 5: The performance of both subsystems using optimised radar waveform. a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 5 illustrates the performance of the radar and communication subsystems using the optimized radar waveform. Fig. 5.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$. As can be seen from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance at higher power-to-noise ratios. It can be also observed that target T3 consistently achieves the highest $\mathrm{KLD_{r}}$, followed by T1 and T2, reflecting their respective distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for T1, T2, and T3 respectively. These values increase to about $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB, showing improvement in radar performance. The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB appears to enter a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, suggesting a consistent pattern of performance enhancement in the radar system.
Fig. 5.b depicts the communication subsystem performance through $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$. The $\mathrm{KLD_{c}}$ for $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB exhibits a minimal improvement, indicative of system infeasibility in this low power regime. This minimal growth at the low SNR region is attributed to the system’s inability to meet the minimum KLD requirements for all UEs simultaneously. UE3 consistently achieves the highest $\mathrm{KLD_{c}}$, followed by UE1 and UE2, which aligns with their respective distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{55.14,20.09,186.09\}$ for UE1, UE2, and UE3 respectively. These values increase substantially to about $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB for all UEs suggests that at this point, the system has reached its optimization feasibility point. Subsequently, the communication performance is enhanced significantly at higher SNR and after reaching feasibility as the radar interference is effectively reduced. These results demonstrate the intricate relationship between radar detection capabilities, and communication system performance in this dual-function system. The analysis shows how the system transitions from infeasibility to high performance as power increases, with both subsystems benefiting once the feasibility threshold is surpassed. The consistent ordering of performance among targets and UEs reflects the impact of their respective distances from the BS on system performance.
Figure 6: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 6.a shows the receiver operating characteristic (ROC) curves depicting the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. At $P_{\mathrm{R}}/{N}_{o}=20$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.909,0.804,0.935\}$. This demonstrates that the optimisation improves detection capabilities comparing it with non-optimisation in Fig. 3. Fig. 6.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The performance of the communication subsystem improves significantly with radar waveform optimization. In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 2. However, with optimization, $P_{e}$ drops below $10^{-5}$ by $13$ dB across all UEs, due to the communication subsystem’s sensitivity to radar interference, which is better mitigated through optimized waveforms.
Figure 7: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 7.a illustrates the trade-off performance between radar and communication subsystems using optimized radar waveform. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value. Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values lead to improved performance in both subsystems. This visualization provides insights into system behaviour across different $P_{\mathrm{R}}/N_{o}$ levels, highlighting the optimization technique’s robustness in balancing radar and communication performance, we can observe that the optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig. 4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
Fig. 7.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$. The plot reveals that the computational complexity of the optimization algorithm remains relatively stable across different $P_{\mathrm{R}}/N_{o}$ values. The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes. This suggests that the optimization technique maintains consistent computational efficiency regardless of the $P_{\mathrm{R}}/N_{o}$ values, which is advantageous for practical implementation in varying channel conditions.
VII-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimization algorithm is configured the same as the radar one with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialized at $\tilde{\mu_{0}}=1$, with a barrier reduction factor of $\tilde{\gamma}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
Fig. 8 illustrates the performance of the radar and communication subsystems when using the optimized communication waveform. Fig. 8.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets in each power allocation scenario, demonstrating enhanced radar performance at higher power-to-noise ratios. For each power allocation, target $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$, indicating that the radar performance is influenced by their respective distances or scattering characteristics. Specifically, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively. These values increase to about $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The linearity observed in $\mathrm{KLD_{r}}$ improvement beyond 15 dB is consistent across the different $P_{\mathrm{r}}$ values, suggesting stable radar system performance at higher $P_{\mathrm{R}}/N_{o}$ values, despite the suboptimal radar waveform.
Fig. 8.b depicts the communication subsystem performance as represented by $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$. $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$, within each power allocation scenario. For $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, respectively. These values increase to around $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the ISAC system’s effort to balance the radar subsystem thresholds with communication subsystem enhancements across different power allocation scenarios.
Figure 8: The performance of both subsystems using optimised communication waveform. a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 9.a shows the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}o$ scenarios. The ROC curves plot the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target.At $P_{\mathrm{R}}/{N}_{o}=10$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.217,0.109,0.273\}$. The results demonstrate that communication optimization slightly improves detection capabilities at lower $P_{\mathrm{R}}/{N}_{o}$ compared to non-optimization, as seen in Fig. 3.
Fig. 9.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The performance of the communication subsystem improves significantly with optimized communication waveform. In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 2. However, with optimization, $P_{e}$ drops below $10^{-4}$ by $15$ dB across all UEs, optimized waveforms still provides a gain in comparison to the non-optimal method, however the gain is lower than optimised radar waveform due to the radar interference.
Figure 9: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 10.a illustrates the trade-off between radar and communication subsystems using the optimized communication waveform. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value. Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values enhance performance in both subsystems. This visualization highlights the robustness of the optimization technique in balancing radar and communication performance across different $P_{\mathrm{R}}/N_{o}$ levels, we can observe that this optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig. 4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
Fig. 10.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$. The plot reveals that the computational complexity of the optimization algorithm stabilizes after $15$ dB $P_{\mathrm{R}}/N_{o}$ values. This is because, at lower $P_{\mathrm{R}}/N_{o}$, the ISAC system cannot maintain the minimum threshold for all targets and UEs, placing the problem in the infeasible region. Once in the feasible region, the average CPU time fluctuates slightly between $0.02$ and $0.04$ seconds.
Figure 10: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
VIII Conclusion
In conclusion, this paper has proposed and validated a novel KLD-based framework for optimizing ISAC systems. The two considered optimization techniques, namely, KLD-based radar waveform optimization and KLD-based communication precoding optimization, demonstrated significant improvements over non-optimized scenarios. Radar waveform KLD-based optimization yielded substantial enhancements in both target detection and communication performance, as evidenced by improved ROC curves and reduced BER across all UEs. Communication KLD-based precoding optimization, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs. Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimization demonstrating stable computational efficiency. Moreover, the trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances overall ISAC system performance. These findings provide valuable insights for designing and optimizing shared deployment ISAC systems, offering a promising approach to balancing sensing and communication requirements in next-generation wireless networks. The proposed KLD-based framework contributes to the holistic design of ISAC systems, paving the way for more efficient and versatile wireless communications and sensing technologies.
Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimization problem.
A-A Gradients for Radar-Waveform Optimisation
The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (39), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (36). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}=\mathbf{R}_{2,t}^{-1}$ (58)
For the trace term, applying $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ yields,
$\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}(\sigma_{n}^{2}\mathbf{I}_{N}))}{\partial\mathbf{R}_{2,t}}=-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}$ (59)
Applying the chain rule to (36), and substituting (58) and (59), we get,
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}$ (60)
where $\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{H}_{t}^{H}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$. Substituting this into (60) and simplifying, we obtain,
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})$ (61)
Finally, the gradient of $f$ with respect to $\mathbf{W}_{\mathrm{r},t}$, derived from (39) and (61), is,
$[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}(\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})_{i}$ (62)
For $\mathrm{KLD}_{\mathrm{c},k}$ gradient calculations, we consider how changes in $\mathbf{W}_{\mathrm{r}}$ affect $\sigma_{\eta}^{2}$, where it is defined as follows,
$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}$ (63)
The derivative of $\sigma_{\eta}^{2}$ with respect to $\mathbf{W}_{\mathrm{r},l}$ is,
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}$ (64)
Computing $\partial\mathrm{KLD}_{\mathrm{c},k}/\partial\sigma_{\eta}^{2}$,
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}$ (65)
Utilising the chain rule and (64), (65), we obtain,
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}$ (66)
Therefore, the gradient of $\mathrm{KLD}_{\mathrm{c},k}$ with respect to $\mathbf{\dot{W}}_{\mathrm{r}}$ is,
$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}$ (67)
Equations (62) and (67) provide the mathematical foundation for the gradients used in our optimization algorithm.

### v2

@@
-Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence
+A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems
 Abstract
-This paper presents a novel framework that leverages the Kullback-Leibler divergence (KLD) to analyze and optimize performance trade-offs in integrated sensing and communication (ISAC) systems.
-We consider a multiple-input-multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using shared antenna deployment.
-The proposed KLD-based approach provides a unified performance measure encompassing both UE error rate and target detection capability.
-We apply this approach on two well-known communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), and evaluate the effect on the radar subsystem.
-Furthermore, two optimization problems are formulated and solved.
-The first one optimizes the KLD of the radar subsystem for given constraints on the communication KLD, whereas the second one focuses on communication waveform KLD-based optimization and constrained radar KLD.
-These optimization problems are solved using a projected gradient method with an adaptive penalty for the radar waveform and a gradient-assisted interior point method for the communication waveform.
-Through theoretical derivations and extensive simulations, it is demonstrated that our approach can be a powerful tool for characterizing and optimizing the performance trade-offs of ISAC under various configurations.
-The results also show significant improvements in both sensing and communication performance by the KLD-optimized system compared to well-known benchmarks, such as conventional MRT and ZF for the communication subsystem, and the conventional identity covariance design for the radar subsystem.
-These findings support the holistic design and optimization of ISAC in next-generation wireless networks.
+This paper introduces a novel framework that leverages the Kullback-Leibler Divergence (KLD) metric to analyse and optimise performance trade-offs in integrated sensing and communication (ISAC) systems.
+We consider a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared antenna deployment.
+We apply this framework to two widely used communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), to assess their impact on the radar subsystem’s performance.
+Additionally, two optimisation problems are formulated: the first optimises the radar subsystem’s KLD under communication constraints, and the second focuses on communication waveform KLD optimisation with constraints on the radar KLD.
+These problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms.
+Through theoretical derivations and extensive simulations, we demonstrate that our KLD approach effectively characterises and optimises the performance trade-offs between sensing and communication in ISAC systems.
+The results show significant improvements in both radar detection and communication performance when compared to traditional MRT and ZF beamforming, and the identity covariance design for radar subsystems.
+These findings promote a more holistic design and optimisation of ISAC for next-generation wireless networks and demonstrate the advantages of KLD-based optimisation in balancing the performance of both sensing and communication.
 Index Terms:
-Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission (MRT), beamforming, optimization, Kullback–Leibler divergence (KLD).
+Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission, beamforming, optimisation, Kullback–Leibler divergence (KLD).
 I Introduction
-The rapid evolution of wireless communication networks, particularly with the advent of fifth-generation (5G) technology, has dramatically expanded the landscape of interconnected devices and services.
-This technological leap has enabled a wide array of applications that depend heavily on advanced sensing capabilities.
-From autonomous vehicles and unmanned aerial vehicles (UAVs) to the Internet-of-Things (IoT) ecosystems, these applications are pushing the boundaries of traditional communication systems, necessitating the integration of sophisticated sensing functionalities [1, 2, 3, 4].
-As we stand on the threshold of the sixth-generation (6G) era, network operators and service providers face unprecedented challenges in not only delivering robust communication services but also in integrating advanced sensing-based applications such as detection, localization, tracking, navigation, and environmental surveillance [5, 6, 7].
-Integrated Sensing and Communications (ISAC) systems have garnered significant interest in the research and industrial communities to meet these emerging demands.
-ISAC systems aim to synergistically merge sensing and communication functionalities within the same infrastructure, thereby optimizing the use of available resources at base stations (BSs) for dual purposes [8, 9, 10, 11, 12, 13].
-This integration can be achieved through various architectural approaches.
-For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilize resources.
-In contrast, the shared deployment strategy employs all antennas for both functionalities, potentially offering higher resource efficiency at the cost of the increased complexity and the need for more sophisticated signal processing techniques[14, 15].
-The evolution of ISAC systems has been accelerated by the advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems.
-These systems allow for the generation of highly directional beams and support simultaneous service to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy, usually done by exploiting spatial diversity by setting the antenna separation to a minimum of half wavelength [16].
+The rapid evolution of wireless communication networks has dramatically expanded the landscape of interconnected devices to include new applications such as autonomous vehicles and unmanned aerial vehicles (UAVs) that are heavily dependent on advanced sensing, pushing the boundaries of communication systems beyond their traditional limits [1, 2, 3, 4].
+As we approach the sixth-generation (6G) era, network operators and service providers aim not only to further enhance their communication services but also to include advanced sensing-based applications such as detection, localisation, tracking, navigation, and environmental surveillance [5, 6, 7].
+Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications.
+ISAC aims to synergistically merge sensing and communication functionalities hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13].
+This integration can be achieved through various approaches.
+For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilise resources.
+In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
+The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems.
+These systems allow for the generation of highly directional beams and support simultaneous services to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy [16].
 The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks.
 However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality.
-Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate, while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17].
-This disparity complicates the holistic assessment and optimization of ISAC systems.
+Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate (BER), while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17].
+This disparity complicates the holistic assessment and optimisation of ISAC systems.
 I-A Literature review
@@
 It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another.
-While the application of KLD in sensing systems is well-established, its potential in characterizing communication system performance remains open.
+While the application of KLD in sensing systems is well-established, its potential in characterising communication system performance remains open.
 Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems.
-For instance, in our previous works [21, 22], we employed KLD to analyze performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimizing network resources for both separated and shared deployment scenarios.
-Moreover, further insights about the achievable KLD trade-off in ISAC systems is explored in [24], and more designs have been discovered.
-Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalized multi-user-multi-target (MUMT) ISAC system using the promising KLD.
-Further efforts have explored various optimization approaches for ISAC systems, each addressing specific aspects of system performance.
-For example, joint radar-communication beamforming optimization based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25].
-This approach minimizes the CRB of angle estimation for radar while maximizing the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously.
-However, these methods often focus on optimizing individual components rather than adopting a unified framework that captures the interplay between different functionalities.
-Beamforming optimization has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26].
-While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in scenarios where these functionalities are deeply integrated.
-Waveform optimization has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27].
+For instance, in our previous works [21, 22], we employed KLD to analyse performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimising network resources for both separated and shared deployment scenarios.
+Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [24], and more designs have been discovered.
+Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target (MUMT) ISAC system using the promising KLD.
+Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance.
+For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25].
+This approach minimises the CRB of angle estimation for radar while maximising the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously.
+However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities.
+Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26].
+While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in deeply integrated scenarios.
+Waveform optimisation has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27].
 This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics.
-However, this work does not consider a unified optimization framework that could simultaneously address the needs of both radar and communication subsystems.
+However, this work does not consider a unified optimisation framework that could simultaneously address the needs of both subsystems.
 I-B Motivation and contribution
 Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems.
-Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilized for radar target detection alongside dedicated radar signals.
-Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities and maximises the system’s overall resource efficiency [14].
+Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals.
+Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities, improving the system’s overall resource efficiency [14].
 This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
-Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterize and optimize the sensing-communication performance trade-offs in shared antenna deployment ISAC systems.
+Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterise and optimise the sensing-communication performance trade-offs in shared antenna deployment ISAC systems.
 This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems.
-By considering the shared deployment scenario and proposing specific optimization techniques for both radar and communication waveforms, we provide a comprehensive approach for system optimization that reflects the true potential of ISAC systems in future wireless networks.
-Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems that can effectively balance the needs of both radar and communication functionalities.
+By considering the shared deployment scenario and proposing specific optimisation techniques for both subsystem waveforms, we provide a comprehensive approach for system optimisation that reflects the true potential of ISAC systems in future wireless networks.
+Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems to effectively balance the needs of both subsystem functionalities.
 Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources.
-The key contributions of our work are as follows:
+The key contributions of this work are:
   • Unified Performance Measure: We derive the Kullback-Leibler Divergence (KLD) specifically for the integrated operation of communication and radar systems within ISAC.
@@
 By deriving this KLD-based metric for communication and radar components, we optimise ISAC system design holistically, addressing the trade-offs and enhancing overall system performance.
-  • Inherent Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation.
-This dual-use approach enhances the radar performance by utilizing communication interference as an additional source of information for target detection.
-  • KLD-based Optimized ISAC: We propose two novel optimization techniques: radar and communication waveform KLD-based optimization.
-Both maximize their respective KLD measures subject to the minimum KLD of both subsystems and power constraints.
-The non-convex radar optimization employs a projected gradient method with an adaptive penalty, while the communication optimization uses a gradient-assisted interior point method.
-These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimization across both subsystems within system constraints.
-  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyze their impact on the overall system performance within our KLD-based framework.
-This analysis provides practical insights for system designers seeking to optimize ISAC performance in shared deployment scenarios with a complexity-performance trade-off.
-Our results demonstrate that the proposed KLD-based optimization techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design.
-The radar waveform optimization shows substantial improvements in both target detection and communication performance, while the communication waveform optimization primarily benefits the communication subsystem with modest radar performance gains.
-Both techniques exhibit robust performance across varying SNR levels, with the radar optimization demonstrating particularly stable computational efficiency.
+  • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation.
+This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection.
+  • KLD-based Optimised ISAC: We propose two novel optimisation techniques: radar and communication waveform KLD-based optimisation.
+Both maximise their respective KLD measures subject to the minimum KLD of both subsystems and power constraints.
+The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method.
+These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimisation across both subsystems within system constraints.
+  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyse their impact on the overall system performance within our KLD-based framework.
+This analysis provides practical insights for optimising ISAC performance in shared deployment scenarios, considering the complexity-performance trade-off.
+Our results demonstrate that the proposed KLD-based optimisation techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design.
+The radar waveform optimisation shows substantial improvements in both target detection and communication performance, while the communication waveform optimisation primarily benefits the communication subsystem with modest radar performance gains.
+Both techniques exhibit robust performance across varying SNR levels, with the radar optimisation demonstrating particularly stable computational efficiency.
 Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources.
 These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
-I-C Paper organization
+I-C Paper organisation
 The paper is structured as follows.
 Section II presents the system model.
-Sections III and IV analyze the communication and radar systems, respectively, including the derivations of KLD for each system.
+Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system.
 In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated.
@@
 II System Model
-This paper is concerned in an ISAC system that comprises an $N$ antenna MIMO-BS.
-The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication user equipments (UEs) in the downlink direction using the shared deployment scenario.
+We consider an $N$ antenna MIMO-BS.
+The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner.
 The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties.
-The portions of the power that are allocated to the radar and communication subsystems are respectively denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$.
+The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$.
 As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16].
-These techniques serve as benchmarks for our subsequent optimized systems, which will be introduced and analyzed later in Sections V and VI.
+These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI.
 The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
@@
 where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem.
-Notably, the transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received version of $\mathbf{x}_{l}$ back to the BS.
-Although the radar signal component received at the UEs might impose interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target.
-Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal for sensing targets as well.
-This dual-use approach enhances radar performance by utilizing the communication signal as an additional source of information.
-Moreover, the radar waveform design in our work is based on the combined transmit signal covariance matrix for the $t$-th beam such that the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, where $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots.
+The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS.
+Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target.
+Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal.
+This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information.
+The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots.
 This formulation allows us to incorporate the communication signals into the radar detection process.
@@
 II-A Communication System
-At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a certain normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$.
-The received signal at the $k$-th UE can be formulated as follows,
+At each $l$ instance, a data symbol $s_{\mathrm{c},k,l}$ intended for the $k$-th UE is drawn from a normalised constellation, i.e, $\mathbb{E}[\left|s_{\mathrm{c},k,l}\right|^{2}]=1$.
+The received signal at the $k$-th UE can be formulated as
 $y_{\mathrm{c},k,l}=\overbrace{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}}^{\text{Desired $k$-th UE signal}}+\overbrace{\omega_{\mathrm{MN},k}}^{\begin{subarray}{c}\text{Inter-user}\\ \text{interference}\end{subarray}}+\overbrace{\eta_{k}}^{\begin{subarray}{c}\text{Radar interference}\\ \text{and noise}\end{subarray}},$ (2)
-where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, the variable $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ represents the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent, and $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ id the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN).
+where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN).
 The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$.
@@
 II-A1 MRT beamforming
-In MRT precoding, the amount of received power at the UEs is maximized without taking into account the imposed interference.
-This approach usually lead to acceptable performance at low SNR values but its performance deteriorates at large SNRs.
-The normalized MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
+In MRT precoding, the amount of received power at the UEs is maximised without taking into account the imposed interference.
+This approach usually leads to acceptable performance at low SNR values but its performance deteriorates at large SNRs.
+The normalised MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
 $\mathbf{W}_{\mathrm{c}}=\mathbf{P}\>\mathbf{H}^{*}/\|\mathbf{H}\|_{F}=\mathbf{P}\>[\mathbf{h}_{1}^{*},\ldots,\mathbf{h}_{K}^{*}]/\|\mathbf{H}\|_{F},$ (3)
@@
 This approach usually results in excellent performance at high SNR values.
-The normalized ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
+The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
 $\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (4)
@@
 II-B Radar system
-The radar system proposed in this study is designed for high adaptability, allowing for real-time adjustments in detection capabilities on a per-frame basis, where each frame is composed of $L$ snapshots.
-A key parameter of this system is $T$, representing the count of radar beams emitted during a specific detection frame.
-By harnessing MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [30, 31].
-Consequently, $T$ signifies the upper limit of detectable targets within a single frame.
-This radar system’s design is robust, as it can explore multiple angular-range-Doppler bins over successive frames, thereby increasing the chance of identifying more targets.
-Furthermore, its flexibility in varying the number of emitted beams allows for the simultaneous detection of a larger set of targets, providing the ability to fine-tune the system’s detection capability according to different operational demands.
-However, it is crucial to acknowledge that the value of $T$ is naturally constrained by the available number of antennas and the total number of UEs being serviced by the BS, highlighting the critical need for efficient waveform design, and resource allocation strategies in ISAC systems.
-Our focus aligns with scenarios where targets are spatially separated, ensuring each target is confined to a distinct radar bin, as explored in prior studies [32, 33].
-Additionally, the literature offers a range of algorithms developed to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34].
-A straightforward method to count detected targets involves applying the detection mechanism across all radar bins spanning the angular, range, and Doppler dimensions.
+The radar system is designed for high adaptability, allowing for real-time adjustments in detection capabilities across $L$ frame snapshots.
+$T$ represents the radar beams emitted during a specific detection frame.
+Using MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [30, 31], where $T$ signifies the upper limit of detectable targets in a single frame.
+The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification.
+Its flexibility in varying emitted beams allows for the simultaneous detection of more targets, to fine-tune the system’s detection capability according to different operational demands.
+However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems.
+Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [32, 33].
+The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34].
 The total radar return signal at the $l$-th snapshot is shown as follows,
@@
 The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
-To analyze the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin.
+To analyse the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin.
 Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target.
@@
 The $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
-III The KLD of communication system: ZF and MRT precoding
-In this section, we derive the KLD of the communication system for two typical scenarios, namely, the MRT, and the ZF beamforming schemes using the shared deployment antenna configuration.
-The analysis is carried out for the normalized precoders introduced in Section II-A.
-In order to, derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise components.
+III The communication KLD: ZF and MRT
+In this section, we derive the KLD of the communication system for the MRT and ZF beamforming using the shared deployment antenna configuration.
+The analysis is carried out for the normalised precoders introduced in Section II-A.
+To derive the KLD, it is necessary to establish the statistical properties of the signal, IUI, radar interference, and noise.
 For a pair of multivariate Gaussian distributed random variables having mean vectors of $\mathbf{\mu}_{m}$ and $\mathbf{\mu}_{n}$ and covariance matrices of $\Sigma_{m}$ and $\Sigma_{n}$, the KLD can be derived as,
-$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right),$ (7)
-The application of this formula to the communication system requires the determination of the mean vectors and covariance matrices for each precoding scheme.
+$\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right).$ (7)
 III-A MRT beamforming
 For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components.
-The examination starts with the IUI term, which is the second term in (2), which can be written as,
+The examination starts with the IUI term, which is the second term in (2), given by,
 $\omega_{\mathrm{MN},k}=d_{\mathrm{c},k}^{-\zeta/2}\mathbf{h}_{k}^{T}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}=\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\|\mathbf{H}\|_{F}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (8)
-Since that the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d.
+Since the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d.
 complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors.
-Therefore, central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a significant number of antennas $N$.
-On the other hand, the normalization term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$.
+Thus, the central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a large number of antennas $N$.
+Meanwhile, the normalisation term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$.
 Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable.
-Unfortunately, this ratio distribution does not have a simple closed-form expression.
-However, it is worth noting that we can approximate its behavior for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$.
-Consequently, using this approximation, we can write (8) as,
+This ratio distribution does not have a simple closed-form expression.
+However, we can approximate its behaviour for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$.
+Using this approximation, we can write (8) as,
 $\omega_{\mathrm{MN},k}\approx\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\sqrt{NK\sigma_{h}^{2}}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (9)
-Now, $\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking the CLT, for large $KN$, this sum approaches a complex normal distribution with zero mean and a variance of,
+$\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking CLT, for large $KN$, the sum approaches a complex normal distribution with zero mean and a variance of
 $\sigma_{\omega,\mathrm{MN}}^{2}\approx\frac{d_{\mathrm{c},k}^{-\zeta}}{NK\sigma_{h}^{2}}\sum_{i\neq k}N\sigma_{h}^{4}p_{i}=\frac{d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}}{K}\sum_{i\neq k}p_{i},$ (10)
-The accuracy of this approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation and shown in Fig.1, which compares it with simulated density functions under different values of $N$.
-It can be also shown that the distribution of $\eta_{k}$ can be well approximated by a complex Gaussian distribution using the CLT, particularly for large $N$.
-This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
+The accuracy of the approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation, which compares it with simulated density functions under different values of $N$.
+It can also be shown that the distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$.
+This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
 $\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (11)
@@
 Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$.
-In order to find the KLD for the MRT case with matrix normalization, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2).
+In order to find the KLD for the MRT case with matrix normalisation, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2).
 This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
@@
 $\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]\approx p_{k}\mathbb{E}[\|\mathbf{h}_{k}\|^{4}]/\mathbb{E}[\|\mathbf{H}\|_{F}^{2}]=p_{k}(N+1)\sigma_{h}^{2}/K.$ (13)
-After deriving the statistical properties of the signal and interference components, the KLD formula can be applied.
-Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, these can be substituted into equation (7), then the KLD can be found as [21, Corollary 1],
+Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, substituting these into equation (7), then the KLD can be found as [21, Corollary 1],
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]}{(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2},$ (14)
-where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ which depends on the transmitted data constellation, and thus $\lambda$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$.
-Finally, substituting (13) in (14), then the KLD for MRT precoding with matrix normalization can be found as,
+where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$.
+Substituting (13) in (14), the KLD for MRT precoding with matrix normalisation is,
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N+1)\sigma_{h}^{2}}{K\>(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2}.$ (15)
-Figure 1: The density functions of the approximated random variable $\omega_{MN}$.
 III-B ZF beamforming
-With ZF precoding, the received signal at the $k$-th user can be rewritten as follows,
+With ZF precoding, the received signal at the $k$-th user is,
 $y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l},$ (16)
-where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalized ZF precoding matrix.
-It should be noticed that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder completely eliminates the IUI.
-The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are shown as follows,
+where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix.
+Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI.
+The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
 $\boldsymbol{\mu}_{k,j}$ $=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]s_{\mathrm{c},k,l}^{(j)}$
@@
 $\ \Sigma_{n}$ $=\Sigma_{m}=\sigma_{\eta}^{2},$ (18)
-where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalization of the ZF precoding matrix.
-Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, it can be shown that $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$.
+where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix.
+Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$.
 Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$.
-By substituting (17), and (18) in (7), then the KLD can be reduced to,
+By substituting (17), and (18) in (7), then the KLD becomes,
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}}{M(M-1)\sigma_{\eta}^{2}\ln 2}\mathbb{E}\left[\frac{1}{\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}}\right],$ (19)
-where the expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
+The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
 […]
-Substituting this result into the expression for $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$ in (19), the final expression for the average KLD in the case of ZF precoder with matrix normalization can be reduced to,
+Substituting this result into (19), we get the final $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$,
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N-K)/M(M-1)\sigma_{\eta}^{2}\ln 2.$ (21)
 III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
-A conditional KLD for communication, in terms of the bamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimization problem for which $\mathbf{W}_{\mathrm{c}}$ can be designed based on the channel side information matrix.
-Accordingly, by considering the received signal representation in (2), the conditional mean vectors and variance can be obtained as follows,
+A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem.
+Considering the received signal representation in (2), the conditional mean vectors and variance are,
 $\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\$ (22)
@@
 $=d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},\!\quad j\in\{n,m\},$ (23)
-where the assumptions of zero-mean, temporal whiteness, and wide-sense stationarity of the communication and radar signals, as well as their mutual orthogonality, have been utilized, it is important noting that the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal.
-Therefore, the covariance is,
+Note that, the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
 $\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$ (24)
-Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
-$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}\times$
-$\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
+Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
+$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
+×$\times\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
 where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ and $\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}=\lambda/(M(M-1))$, substituting it in the KLD expression which becomes,
@@
 IV-A Composite detection and response matrix estimation
-The main objective for typical radar systems is to detect targets and estimate the matrix response associated with each one.
-After separating different signals associated with different targets, such problem can be formulated for each target as a composite detection-estimation problem as follows,
+The main objective of typical radar systems is to detect targets and estimate the matrix response for each one.
+After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
 $\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{q}).$ (27)
-This joint optimization problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present.
-Now, considering the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each of these hypotheses can be shown as a multivariate normal density, which can be written as
+This joint optimisation problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present.
+For the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis can be shown as a multivariate normal density,
 $f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
@@
 where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$.
-The generalized likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses.
-The GLRT is derived as follows,
+The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, derived as,
 $\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (29)
-However, the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$.
-Therefore, the target response matrix is typically estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain
+As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain,
 $\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1}).$ (30)
-Since the noise $\mathbf{n}_{t,l}$ is Gaussian and the $\mathbf{H}_{t}$ is normally distributed, the likelihood function is Gaussian, and the maximization leads to the least squares solution.
-Therefore, this maximization reduces to the following form,
+The likelihood function is Gaussian, and the maximisation leads to the least squares solution,
 $\hat{\mathbf{H}}_{t}$ $=\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}-\mathbf{H}_{t}\mathbf{x}_{t}\right\|^{2}=\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t}^{H}\left(\mathbf{x}_{t}\mathbf{x}_{t}^{H}\right)^{-1}.$ (31)
-This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain the following,
+This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain,
 $\Lambda(\mathbf{y}_{\mathrm{r},t,l})=$
 $\frac{\pi^{N}\sigma_{n}^{2N}\exp\left(-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\right)}{\pi^{N}\det(\hat{\mathbf{R}}_{1,t})\exp\left(-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\right)}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{0}$ (32)
-where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold that is a design parameter, for example, according to Neyman-Pearson lemma it can be determined to obtain a certain false alarm rate.
-Taking the logarithm and rearranging terms, we obtain the following expression, which now clearly includes the estimated target response matrix $\hat{\mathbf{H}}_{t}$,
+where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold.
+Taking the logarithm and rearranging, we obtain,
 $\ln\Lambda(\mathbf{y}_{\mathrm{r},t,l})$ $=\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}-N\ln\sigma_{n}^{2}+\ln\det(\hat{\mathbf{R}}_{1,t})$
@@
 where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$.
-This final expression correctly reflects the comparison between the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$.
-In order to utilize the whole $L$ snapshots whose received signals are independent, we average as follows,
+This expression compares hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$.
+To use all $L$ snapshots with independent signals, we average as,
 $\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}.$ (35)
-where $\tau_{2}$ is the detection threshold that accounts for all snapshots and could be designed according to the Neyman-Pearson lemma as illustrated above.
+where $\tau_{2}$ is the detection threshold considering all snapshots and can be designed per the Neyman-Pearson lemma.
 IV-B The KLD of radar system
 […]
-Leveraging this asymptotic property and utilizing the $L$ number of snapshots, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (28), and (7) as follows,
+Using this property, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (28), and (7) as follows,
 $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=$ $\frac{1}{\ln{2}}\Big{(}\ln\left(\mathrm{det}(\mathbf{R}_{2,t})\right)+\mathrm{Tr}\left(\mathbf{R}_{2,t}^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)$
@@
 where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation.
-It should be noted that $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
+$\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
 V Radar Waveform Optimisation
-Here, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem can be shown as follows,
+In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
 $\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (37a)
@@
 After substituting (36) in (37a), can be written as $\mathcal{P}_{2}$ that is given on page V.
-This optimization problem focuses on the radar waveform while considering the whole ISAC system performance.
-The constraint on radar power (37d) is explicit because the radar waveform is our primary optimization variable.
+This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance.
+The constraint on radar power (37d) is explicit because the radar waveform is our primary optimisation variable.
 The communication power is implicitly considered through the ZF beamforming scheme and the constraint (37c) on communication performance.
-While the communication power is not directly optimized, it is indirectly influenced by the radar power allocation.
-Specifically, the communication subsystem utilizes its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
-This approach allows us to focus on the interplay between subsystems which is mainly governed by waveform design and interference management while maintaining system-wide performance.
+Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
 $\mathcal{P}_{2}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}$ $\!\!\!\!\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\ln{2}}\Bigg{(}\ln\left(\mathrm{det}(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N})\right)$ (38)
@@
 s.t. (37b), (37c), (37d), and (37e)
-To Tackle the optimisation problem in (38), we utilize the Projected Gradient method with a penalty function [36].
-This method is chosen for its suitability to the non-linear, non-convex optimization problem [37].
-This approach incorporates complex KLD constraints via the penalty function, while efficiently handles the radar power constraint through projection.
-It offers a balance between the solution quality and the computational efficiency, crucial for potential real-time ISAC applications.
-The method’s scalability with problem dimensionality further justifies the selection of this method for the radar waveform optimization task in the ISAC setup illustrated in the problem formulation.
-The optimization variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor.
+To solve the optimisation problem in (38), we utilise the Projected Gradient method with a penalty function [36].
+This method is chosen for its suitability to the non-linear, non-convex optimisation problem [37].
+This method uses a penalty function for KLD constraints and handles radar power via projection, balancing the solution quality and computational efficiency for real-time ISAC applications.
+Its scalability makes it suitable for radar waveform optimisation in ISAC.
+The optimisation variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor.
 We define the objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ and a penalty function $p(\mathbf{\dot{W}}_{\mathrm{r}})$ to handle the constraints in (37b) and (37c),
@@
 $\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}+\alpha_{n}\mathbf{G}^{(n)}\right),$ (45)
-where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, which balances the importance of the objective function and the constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (37d) is always satisfied.
-The projection mechanism is shown in the following equation,
+where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, balancing the objective and constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (37d) is always satisfied.
+The projection mechanism is shown as,
 $\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}})=\begin{cases}\mathbf{\dot{W}}_{\mathrm{r}}\sqrt{\frac{P_{\mathrm{r}}L}{\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}}},&\text{if }\frac{1}{L}\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}>P_{\mathrm{r}}\\ \mathbf{\dot{W}}_{\mathrm{r}},&\text{otherwise}\end{cases}$ (46)
@@
 where $c\in(0,1)$.
-This condition, known as the Armijo condition, ensures a sufficient decrease in the penalized objective function [38].
-The step size is then set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, where $\beta\in(0,1)$ is a reduction factor.
+This condition, known as the Armijo condition, ensures a sufficient decrease in the penalised objective function [38].
+The step size is set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, with $\beta\in(0,1)$ as a reduction factor.
 The penalty parameter $\rho_{n}$ is updated as follows,
@@
 where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied.
-Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimization
+Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimisation
 0:  Initial point $\mathbf{\dot{W}}_{\mathrm{r}}^{(0)}$, constants $\alpha_{0}$, $\rho_{0}$, $\beta$, $\gamma$, $c$, $P_{\mathrm{r}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\varepsilon$, and maximum iterations $\mathrm{max\_iter}$
 0:  Optimal solution $\mathbf{\dot{W}}_{\mathrm{r}}^{*}$
-1:  Initialize $n=0$
+1:  Initialise $n=0$
 2:  while $n<\mathrm{max\_iter}$ do
@@
 14:  return  $\mathbf{\dot{W}}_{\mathrm{r}}^{*}=\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}$
-Algorithm 1 implements the projected gradient method with penalty for the radar waveform optimization problem.
-It iteratively updates the waveform $\mathbf{\dot{W}}_{\mathrm{r}}$ by moving in the direction of the total gradient $\mathbf{G}^{(n)}$, which combines the gradients of the objective function gradient and the penalty function.
 The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied.
-This projection is crucial as it maintains the feasibility of the solution with respect to the power constraint throughout the optimization process.
 The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration.
@@
 $\begin{split}\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\leq&\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})\\ &+c\alpha_{n}\langle\nabla\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}),\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\end{split}$ (49)
-where $\phi(\mathbf{\dot{W}}_{\mathrm{r}})=f(\mathbf{\dot{W}}_{\mathrm{r}})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}})$ is the penalized objective function.
-The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation.
-This adaptive scheme allows the algorithm to balance between optimizing the objective function and satisfying the constraints.
-As the optimization progresses, if constraints are consistently violated, the increasing penalty parameter will put more emphasis on constraint satisfaction.
+where $\phi(\mathbf{\dot{W}}_{\mathrm{r}})=f(\mathbf{\dot{W}}_{\mathrm{r}})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}})$ is the penalised objective function.
+The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation, allowing the algorithm to balance optimisation and constraint satisfaction.
+If constraints are consistently violated, the increasing penalty will emphasise constraint adherence.
 VI Communication optimisation
-The communication subsystem optimization employs the gradient-assisted interior point method [39], contrasting with the projected gradient method with penalty used for radar optimization, due to the fundamental difference in optimization variable structure.
-While radar optimization involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, communication beamforming optimizes over a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$.
-This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions for the matrix optimization problem.
-The communication waveform is optimised while using conventional identity covariance matrix design for the radar subsystem with a covariance matrix set to $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$.
-The optimisation problem can be shown as follows,
+The communication subsystem optimisation uses the gradient-assisted IPM [39], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures.
+Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$.
+This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions.
+The communication waveform is optimised while using the conventional identity covariance matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$.
+The optimisation problem can be expressed as follows,
 $\mathcal{P}_{3}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (50a)
@@
 $P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.$ (50e)
-It is important to note that the radar power is uniformly distributed among the potential targets, as the radar covariance matrix ensures an omnidirectional beam flow.
-To solve this optimization problem, we need to derive the gradient of the objective function and the constraints.
-Let’s define the objective function as,
+Note, that the radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow.
+To solve this optimisation problem, we must derive the gradients of the objective function and the constraints.
+Let’s define the objective function as follows,
 $f(\mathbf{W}_{\mathrm{c}})=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}.$ (51)
-Now, to compute the gradient of the objective function, we need to derive the gradients of each objective function $\mathbf{W}_{\mathrm{c}}$ components.Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$.
-Using the chain rule and the properties of matrix derivatives, the gradients can be expressed as,
+To compute the gradient of the objective function, we need to derive the gradients of each component of $\mathbf{W}_{\mathrm{c}}$.
+Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$.
+Using the chain rule and properties of matrix derivatives, the gradients can be expressed as follows,
 $\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\quad\text{for all }j$ (52)
@@
 $\begin{split}\nabla_{\mathbf{w}_{\mathrm{c},k}}f(\mathbf{W}_{\mathrm{c}})=\frac{d_{\mathrm{c},k}^{\zeta}C}{K}\sum_{k=1}^{K}\Bigg{(}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}}\\ -\sum_{j\neq k}\frac{2\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}}\Bigg{)}.\end{split}$ (54)
-Next, we need to consider the constraints of our optimization problem. The power constraint is given by,
+Next, we need to consider the constraints of our optimisation problem. The power constraint is given by,
 $g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.$ (55)
-The gradient of this constraint is derived straightforwardly as follows,
+The gradient of this constraint is derived as follows,
 $\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.$ (56)
@@
 $\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).$ (57)
-Algorithm 2 Interior Point Method for Communication Optimization
+Algorithm 2 Interior Point Method for Communication Optimisation
 0:  Initial point $\mathbf{W}_{c}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
 0:  Optimal solution $\mathbf{W}_{c}^{*}$
-1:  Initialize $n=0$, $\mu=\mu_{0}$
+1:  Initialise $n=0$, $\mu=\mu_{0}$
 2:  while $n<\mathrm{max\_iter}$ do
@@
 14:  return  $\mathbf{W}_{c}^{*}=\mathbf{W}_{c}^{(n)}$
-Solving the optimization problem defined in equations (50a)-(50d) involves employing the Interior Point Method (IPM) as outlined in Algorithm 2.
-This method efficiently handles constrained optimization using key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$.
-The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions, and updates the solution using a Newton system and line search.
-The barrier parameter $\tilde{\mu}$ is reduced as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly.
-This process continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximizes average communication KLD while satisfying radar and power constraints.
+Solving the optimisation problem defined in equations (50a)-(50d) uses the IPM, as detailed in Algorithm 2.
+This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$.
+The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search.
+The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly.
+This continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
 VII Numerical Results
-In this section, we present the simulation results for both our conventional benchmarks and the two optimisation techniques introduced previously.
-The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section.
-The total number of antennas at BS is $N=20$, a number of $L=100$ snapshots, the maximum number of potential radar targets is $T=3$, and the number of UEs is $K=3$.
-The pathloss exponent is $\eta=3$ to model the effect of large-scale fading, and the distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters, respectively, and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters, respectively.
+In this section, we present simulation results for our conventional benchmarks and the two previously introduced optimisation techniques.
+The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout.
+The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets, and $K=3$ UEs.
+The pathloss exponent is $\eta=3$ to model large-scale fading.
+The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters.
 The channel variance is $\sigma_{h}^{2}=1$.
 VII-A Trade-off for the conventional benchmarks
-The conventional identity covariance matrix design is used for the radar subsystem with a covariance matrix is set to $\mathbf{R}_{w}=\mathbf{I}_{N}$, whereas the communication subsystem uses MRT and ZF precoding introduced earlier.
-Figure 2: The performance of the communication subsystem for both beamforming schemes, where the Bit Error Rate (BER) and the achiveable $\mathrm{KLD}_{\mathrm{c}}$ are shown versus $P_{\mathrm{R}}/{N}_{o}$, with $P_{\mathrm{r}}=0.5$.
+The radar subsystem employs a conventional identity covariance matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced MRT and ZF.
+Figure 1: Performance of the communication subsystem for both beamforming schemes, showing BER and achievable $\mathrm{KLD}_{\mathrm{c}}$ versus $P_{\mathrm{R}}/{N}_{o}$ with $P_{\mathrm{r}}=0.5$.
 Fig.
-2 presents the performance of the communication subsystem for both ZF and MRT beamforming schemes.
+1 shows the communication subsystem performance for both ZF and MRT beamforming.
 In Fig.
-2.a, the MRT scheme’s performance is depicted, with the KLD performance on the right $y$-axis and the BER performance on the left.
-As can be observed from the figure, the KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ until around $10$ dB, where a noticeable upper-bound occurs primarily caused by radar interference, with IUI contributing to a less extent.
-Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, indicating an improvement in the communication subsystem performance, but also exhibits the same error floor at around $10$ dB.
+1.a, the MRT scheme’s performance is illustrated, with KLD on the right $y$-axis and BER on the left.
+The KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ up to about $10$ dB, where an upper bound appears due to radar interference, with IUI contributing less significantly.
+Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, reflecting improved communication performance but also reaches an error floor around $10$ dB.
 Fig.
-2.b illustrates the ZF scheme’s performance, exhibiting similar trends in the KLD and BER curves.
-However, compared to the MRT scheme, the ZF scheme shows an increase in KLD values and a decrease in BER values for the same received SNR, indicating that the primary source of interference is indeed caused by radar interference as IUI is cancelled using ZF precoding.
-Consequently, both the ZF and MRT schemes exhibit a drastic error floor in BER, or equivalently an upper bound for KLD, due to the unmitigated radar interference at $P_{\mathrm{c}}=0.5$.
-Moreover, this figure clearly shows the effectiveness of using KLD to characterize the communication system.
-Figure 3: The receiver operating characteristic (ROC) curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
+1.b presents the ZF scheme’s performance, showing similar KLD and BER trends.
+However, the ZF scheme exhibits higher KLD values and lower BER values for the same received SNR compared to MRT, confirming that radar interference is the primary source, as IUI is cancelled using ZF precoding.
+Both schemes display a significant error floor in BER and an upper bound in KLD due to persistent radar interference at $P_{\mathrm{c}}=0.5$.
+This figure effectively demonstrates the utility of KLD in characterising the communication system.
+Figure 2: The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
 Fig.
-3 shows the receiver operating characteristic (ROC) curves depicting the performance of the radar subsystem for the three potential targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios, for the same ISAC system in Fig.
-2 with ZF precoding for communication subsystem assumed.
-The ROC curves are obtained by plotting the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target.
+2 presents the receiver operating characteristic (ROC) curves for the radar subsystem’s performance across three potential targets at varying $P_{\mathrm{R}}/{N}_{o}$ levels, using ZF precoding for the communication subsystem as shown in Fig.
+1.
+The ROC curves plot the probability of detection ($P_{D}$) against the probability of false alarm ($P_{FA}$) for each target.
+At $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are closely aligned, indicating similar detection performance at low SNR.
+However, at $P_{\mathrm{R}}/{N}_{o}=10$ dB, the curves diverge, with $\mathrm{T}_{3}$ showing the best performance, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$.
+This trend continues at $P_{\mathrm{R}}/{N}_{o}=20$ dB, where the separation is even more pronounced.
+Notably, at $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively.
+$\mathrm{T}_{3}$ remains the most detectable, while both $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant detection improvements compared to lower $P_{\mathrm{R}}/{N}_{o}$ scenarios.
+This illustrates that higher SNR enhances the radar subsystem’s capability to detect and differentiate targets effectively.
+Figure 3: a) Radar subsystem performance: achievable $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. b) KLD trade-off between subsystems for ZF and MRT.
 In Fig.
-3, where $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for all three targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are relatively close to each other, indicating similar detection performance at low signal-to-noise ratios.
-In Fig.
-3, with $P_{\mathrm{R}}/{N}_{o}=10$ dB, the ROC curves for the three targets start to diverge, where $\mathrm{T}_{3}$ exhibits the best performance, followed by $\mathrm{T}_{1}$ followed by $\mathrm{T}_{2}$.
-This suggests that as the signal-to-noise ratio increases, the radar subsystem’s ability to detect the targets and distinguish between them improves.
-Finally, in Fig.
-3, in which $P_{\mathrm{R}}/{N}_{o}=20$ dB, the separation between the ROC curves becomes even more pronounced.
-As can be observed in 3, at $P_{FA}=10^{-3}$, $P_{D}$ is $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, where $\mathrm{T}_{3}$ maintains its superior performance, while $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant improvements in their detection probabilities compared to the lower $P_{\mathrm{R}}/{N}_{o}$ scenarios.
-This demonstrates that at high signal-to-noise ratios, the radar subsystem can effectively detect and differentiate between the three targets, with $\mathrm{T}_{3}$ being the most easily detectable as it is the closest target to the BS.
-Figure 4: a) The performance of the radar subsystem where the achievable $\mathrm{KLD_{\mathrm{r}}}$ is shown versus $P_{\mathrm{R}}/{N}_{o}$, across different values of $P_{\mathrm{r}}$.
-b) The KLD performance trade-off between the radar and communication subsystems for ZF and MRT
-In Fig.4.a, the performance of the radar subsystem is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus the $P_{\mathrm{R}}/{N}_{o}$ across different values of $P_{\mathrm{r}}$.
-It is shown that as $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ for each target also increases for all values of $P_{\mathrm{r}}$.
-This indicates that the radar subsystem’s performance improves with higher $P_{\mathrm{R}}/{N}_{o}$.
-The performance curves for different $P_{\mathrm{r}}$ values are relatively close to each other, suggesting that the system’s performance is less sensitive to changes in $P_{\mathrm{r}}$ because the communication subsystem signal is also incorporated in the target detection, using our improved system model.
+3.a, the radar subsystem performance is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$.
+As $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ also improves, indicating enhanced radar performance.
+The performance curves for different $P_{\mathrm{r}}$ values are close, suggesting that the system is less sensitive to $P_{\mathrm{r}}$ changes due to the communication signal’s incorporation in target detection using our improved system model.
 Fig.
-4.b presents the trade-off performance between radar and communication subsystems using the ZF and MRT for the communication subsystem and the conventional identity covariance design for the radar subsystem.
-The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation between both subsystems, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value.
-The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ outperforms the $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ slightly, as both of them exhibits saturation at higher $P_{\mathrm{R}}/N_{o}$ due to the destructive interference from the radar subsystem mainly, while the $\mathrm{KLD_{r}}$ shows a steady increase when moving from left to right on the curve.
-It can be observed from the figure that a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits can be achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same amount of $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ is achieved at around $25$ dB, which implies an SNR gain of about $13$ dB in favour of ZF.
+3.b presents the trade-off performance between both subsystems using ZF and MRT for communication, along with the conventional identity covariance design for radar.
+The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value going higher from left to right.
+The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ slightly outperforms $\mathrm{KLD_{c}}^{\mathrm{MRT}}$, both exhibiting saturation at higher $P_{\mathrm{R}}/N_{o}$ due to radar interference, while $\mathrm{KLD_{r}}$ steadily increases along the curve.
+Notably, a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits is achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ requires around $25$ dB, implying an SNR gain of about $13$ dB in favor of ZF.
 VII-B Radar Waveform KLD-based Optimisation
-The radar waveform KLD-based optimization algorithm is configured with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
-The penalty parameter is initialized at $\rho_{0}=1$, with a penalty increase parameter of $\gamma=1.5$, and an initial step size of $\alpha_{0}=0.1$.
-The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
+The radar waveform KLD-based optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$.
+The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
 Figure 4: The performance of both subsystems using optimised radar waveform.
@@
 Fig.
-5 illustrates the performance of the radar and communication subsystems using the optimized radar waveform.
+4 illustrates the performance of both subsystems with the optimised radar waveform.
+In Fig.
+4.a, the achievable $\mathrm{KLD_{r}}$ for each target versus $P_{\mathrm{R}}/N_{o}$ is shown.
+As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance.
+$\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, indicating consistent performance enhancement in the radar system.
 Fig.
-5.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$.
-As can be seen from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance at higher power-to-noise ratios.
-It can be also observed that target T3 consistently achieves the highest $\mathrm{KLD_{r}}$, followed by T1 and T2, reflecting their respective distances from the BS.
-At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for T1, T2, and T3 respectively.
-These values increase to about $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB, showing improvement in radar performance.
-The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB appears to enter a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, suggesting a consistent pattern of performance enhancement in the radar system.
-Fig.
-5.b depicts the communication subsystem performance through $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$.
-The $\mathrm{KLD_{c}}$ for $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB exhibits a minimal improvement, indicative of system infeasibility in this low power regime.
-This minimal growth at the low SNR region is attributed to the system’s inability to meet the minimum KLD requirements for all UEs simultaneously.
-UE3 consistently achieves the highest $\mathrm{KLD_{c}}$, followed by UE1 and UE2, which aligns with their respective distances from the BS.
-At $P_{\mathrm{R}}/N_{o}=10$ dB, $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{55.14,20.09,186.09\}$ for UE1, UE2, and UE3 respectively.
-These values increase substantially to about $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB for all UEs suggests that at this point, the system has reached its optimization feasibility point.
-Subsequently, the communication performance is enhanced significantly at higher SNR and after reaching feasibility as the radar interference is effectively reduced.
-These results demonstrate the intricate relationship between radar detection capabilities, and communication system performance in this dual-function system.
-The analysis shows how the system transitions from infeasibility to high performance as power increases, with both subsystems benefiting once the feasibility threshold is surpassed.
-The consistent ordering of performance among targets and UEs reflects the impact of their respective distances from the BS on system performance.
+4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$.
+For $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB, $\mathrm{KLD_{c}}$ exhibits minimal improvement, indicating system infeasibility in this low power regime due to the inability to meet minimum KLD requirements for all UEs.
+$\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB suggests the system has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced.
+These results highlight the intricate relationship between radar detection capabilities and communication system performance, illustrating how both subsystems benefit once the feasibility threshold is surpassed, with the performance reflecting the respective distances from the BS.
 Figure 5: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-6.a shows the receiver operating characteristic (ROC) curves depicting the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
-At $P_{\mathrm{R}}/{N}_{o}=20$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.909,0.804,0.935\}$.
-This demonstrates that the optimisation improves detection capabilities comparing it with non-optimisation in Fig.
-3.
+5.a presents the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
+At $P_{\mathrm{R}}/{N}_{o}=20$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.909,0.804,0.935\}$ in the non-optimal case, demonstrating improved detection capabilities through optimisation.
 Fig.
-6.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$.
-The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$.
-$\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
-The performance of the communication subsystem improves significantly with radar waveform optimization.
-In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig.
-2.
-However, with optimization, $P_{e}$ drops below $10^{-5}$ by $13$ dB across all UEs, due to the communication subsystem’s sensitivity to radar interference, which is better mitigated through optimized waveforms.
+5.b shows the BER performance for the three UEs versus $P_{\mathrm{R}}/N_{o}$.
+As $P_{\mathrm{R}}/N_{o}$ increases, BER decreases for all UEs, with $\mathrm{UE}_{3}$ achieving the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
+The communication subsystem significantly benefits from radar waveform optimisation, as evidenced by the non-optimal case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig.
+1, while optimised waveforms reduce $P_{e}$ below $10^{-5}$ by $13$ dB, effectively mitigating radar interference.
 Figure 6: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-7.a illustrates the trade-off performance between radar and communication subsystems using optimized radar waveform.
-The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value.
-Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values lead to improved performance in both subsystems.
-This visualization provides insights into system behaviour across different $P_{\mathrm{R}}/N_{o}$ levels, highlighting the optimization technique’s robustness in balancing radar and communication performance, we can observe that the optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig.
-4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
+6.a illustrates the trade-off performance between both subsystems using the optimised radar waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values.
+As $P_{\mathrm{R}}/N_{o}$ increases, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems.
+This highlights the robustness of the optimisation technique in balancing radar and communication performance.
+In comparison, the non-optimal approach in Fig.
+3.b saturates at around $11$ bits, indicating an increasing and favourable trade-off for the optimal way.
 Fig.
-7.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$.
-The plot reveals that the computational complexity of the optimization algorithm remains relatively stable across different $P_{\mathrm{R}}/N_{o}$ values.
-The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes.
-This suggests that the optimization technique maintains consistent computational efficiency regardless of the $P_{\mathrm{R}}/N_{o}$ values, which is advantageous for practical implementation in varying channel conditions.
+6.b presents the average CPU time of the radar waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$.
+The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes, indicating consistent computational efficiency across varying $P_{\mathrm{R}}/N_{o}$ values, which is beneficial for practical implementation in diverse channel conditions.
 VII-C Communication Waveform KLD-based Optimisation
-The communication waveform KLD-based optimization algorithm is configured the same as the radar one with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
-The barrier parameter is initialized at $\tilde{\mu_{0}}=1$, with a barrier reduction factor of $\tilde{\gamma}=0.1$.
-The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
+The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
+The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$.
+The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
 Fig.
-8 illustrates the performance of the radar and communication subsystems when using the optimized communication waveform.
+7 illustrates the performance of the radar and communication subsystems using the optimised communication waveform.
 Fig.
-8.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$.
-As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets in each power allocation scenario, demonstrating enhanced radar performance at higher power-to-noise ratios.
-For each power allocation, target $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$, indicating that the radar performance is influenced by their respective distances or scattering characteristics.
-Specifically, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively.
-These values increase to about $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The linearity observed in $\mathrm{KLD_{r}}$ improvement beyond 15 dB is consistent across the different $P_{\mathrm{r}}$ values, suggesting stable radar system performance at higher $P_{\mathrm{R}}/N_{o}$ values, despite the suboptimal radar waveform.
+7.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for power allocations of $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$.
+As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$.
+For example, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$, increasing to $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The linear improvement in $\mathrm{KLD_{r}}$ beyond $15$ dB indicates stable radar performance despite the suboptimal waveform.
 Fig.
-8.b depicts the communication subsystem performance as represented by $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$.
-$\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$, within each power allocation scenario.
-For $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, respectively.
-These values increase to around $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the ISAC system’s effort to balance the radar subsystem thresholds with communication subsystem enhancements across different power allocation scenarios.
+7.b depicts the communication subsystem performance with $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$.
+$\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
+At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$, rising to $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem thresholds with communication enhancements across different power allocations.
 Figure 7: The performance of both subsystems using optimised communication waveform.
@@
 Fig.
-9.a shows the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}o$ scenarios.
-The ROC curves plot the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target.At $P_{\mathrm{R}}/{N}_{o}=10$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.217,0.109,0.273\}$.
-The results demonstrate that communication optimization slightly improves detection capabilities at lower $P_{\mathrm{R}}/{N}_{o}$ compared to non-optimization, as seen in Fig.
-3.
+8.a shows the ROC curves for the three targets at different $P_{\mathrm{R}}/N_{o}$ levels.
+At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.217,0.109,0.273\}$ for the non-optimal case.
+These results indicate that communication optimisation enhances detection capabilities at lower $P_{\mathrm{R}}/N_{o}$ compared to non-optimisation, as shown in Fig.
+2.
 Fig.
-9.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$.
-The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$.
+8.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$.
+The BER decreases with increasing $P_{\mathrm{R}}/N_{o}$, reflecting improved communication performance.
 $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
-The performance of the communication subsystem improves significantly with optimized communication waveform.
-In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig.
-2.
-However, with optimization, $P_{e}$ drops below $10^{-4}$ by $15$ dB across all UEs, optimized waveforms still provides a gain in comparison to the non-optimal method, however the gain is lower than optimised radar waveform due to the radar interference.
+In the non-optimal scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB (Fig.
+1), while with optimisation, $P_{e}$ drops below $10^{-4}$ by 15 dB across all UEs.
+Although optimised waveforms provide a performance gain over the non-optimal method, the gain is less pronounced than with optimised radar waveforms due to radar interference.
 Figure 8: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-10.a illustrates the trade-off between radar and communication subsystems using the optimized communication waveform.
-The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value.
-Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values enhance performance in both subsystems.
-This visualization highlights the robustness of the optimization technique in balancing radar and communication performance across different $P_{\mathrm{R}}/N_{o}$ levels, we can observe that this optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig.
-4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
+9.a illustrates the trade-off between both subsystems using the optimised communication waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values.
+As $P_{\mathrm{R}}/N_{o}$ increases from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems.
+This highlights the robustness of the optimisation technique in balancing radar and communication performance.
+In contrast, the non-optimal case in Fig.
+3.b shows $\mathrm{KLD_{c}}$ saturating around 11 bits.
 Fig.
-10.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$.
-The plot reveals that the computational complexity of the optimization algorithm stabilizes after $15$ dB $P_{\mathrm{R}}/N_{o}$ values.
-This is because, at lower $P_{\mathrm{R}}/N_{o}$, the ISAC system cannot maintain the minimum threshold for all targets and UEs, placing the problem in the infeasible region.
-Once in the feasible region, the average CPU time fluctuates slightly between $0.02$ and $0.04$ seconds.
+9.b presents the average CPU time for the communication waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$.
+The computational complexity stabilises after $15$ dB as the ISAC system transitions from an infeasible region, where minimum thresholds cannot be met for all targets and UEs, to a feasible region, with average CPU time fluctuating between $0.02$ and $0.04$ seconds.
 Figure 9: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
 VIII Conclusion
-In conclusion, this paper has proposed and validated a novel KLD-based framework for optimizing ISAC systems.
-The two considered optimization techniques, namely, KLD-based radar waveform optimization and KLD-based communication precoding optimization, demonstrated significant improvements over non-optimized scenarios.
-Radar waveform KLD-based optimization yielded substantial enhancements in both target detection and communication performance, as evidenced by improved ROC curves and reduced BER across all UEs.
-Communication KLD-based precoding optimization, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs.
-Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimization demonstrating stable computational efficiency.
-Moreover, the trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances overall ISAC system performance.
-These findings provide valuable insights for designing and optimizing shared deployment ISAC systems, offering a promising approach to balancing sensing and communication requirements in next-generation wireless networks.
-The proposed KLD-based framework contributes to the holistic design of ISAC systems, paving the way for more efficient and versatile wireless communications and sensing technologies.
+This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems.
+The two considered optimisation techniques, namely, KLD-based radar waveform optimisation and KLD-based communication precoding optimisation, demonstrated significant improvements over non-optimised scenarios.
+Radar waveform KLD-based optimisation yielded substantial enhancements in both target detection and communication performance, as evidenced by the ROC curves and reduced BER across all UEs.
+Communication KLD-based precoding optimisation, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs.
+Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimisation demonstrating stable computational efficiency.
+Moreover, trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances the overall ISAC system performance.
+These findings provide valuable insights for designing and optimising shared deployment ISAC systems, offering a promising approach to balance sensing and communication requirements in future wireless networks.
+The proposed framework provides the means for a holistic ISAC design, paving the way for more efficient and versatile wireless communications and sensing technologies.
 Appendix A Gradient Derivations
-This appendix provides detailed derivations for the gradients used in our optimization problem.
+This appendix provides detailed derivations for the gradients used in our optimisation problem.
 A-A Gradients for Radar-Waveform Optimisation

### v4

@@
 Abstract
-This paper introduces a novel framework that leverages the Kullback-Leibler Divergence (KLD) metric to analyse and optimise performance trade-offs in integrated sensing and communication (ISAC) systems.
-We consider a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared antenna deployment.
-We apply this framework to two widely used communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), to assess their impact on the radar subsystem’s performance.
-Additionally, two optimisation problems are formulated: the first optimises the radar subsystem’s KLD under communication constraints, and the second focuses on communication waveform KLD optimisation with constraints on the radar KLD.
-These problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms.
-Through theoretical derivations and extensive simulations, we demonstrate that our KLD approach effectively characterises and optimises the performance trade-offs between sensing and communication in ISAC systems.
-The results show significant improvements in both radar detection and communication performance when compared to traditional MRT and ZF beamforming, and the identity covariance design for radar subsystems.
-These findings promote a more holistic design and optimisation of ISAC for next-generation wireless networks and demonstrate the advantages of KLD-based optimisation in balancing the performance of both sensing and communication.
+This paper introduces a novel framework aimed at designing integrated waveforms for robust integrated sensing and communication (ISAC) systems.
+The system model consists of a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared-antenna deployment scenario.
+By leveraging Kullback-Leibler divergence (KLD) to holistically characterise both communication and sensing subsystems, three optimisation problems are formulated: (i) radar waveform KLD maximisation under communication constraints, (ii) communication waveform KLD maximisation subject to radar KLD requirements, and (iii) an integrated waveform KLD-based optimisation for ISAC that jointly balances both subsystems.
+The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms.
+The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby strengthening the trade-off between sensing and communication objectives and achieving higher overall performance than the individual radar- or communication-only techniques.
+Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem.
+Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming and identity-covariance radar baselines, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
 Index Terms:
-Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission, beamforming, optimisation, Kullback–Leibler divergence (KLD).
+Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, beamforming, optimisation, Kullback–Leibler divergence (KLD).
 I Introduction
@@
 In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
-The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems.
-These systems allow for the generation of highly directional beams and support simultaneous services to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy [16].
-The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks.
+The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems, which significantly enhance both communication capacity and sensing accuracy [16].
 However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality.
-Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate (BER), while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17].
+Traditionally, communication and sensing subsystems have been assessed using distinct metrics—communication performance through achievable rate, outage probability, and bit error rate (BER), and sensing performance through estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17].
 This disparity complicates the holistic assessment and optimisation of ISAC systems.
+To this end, some unified performance measures have been introduced in the literature, such as the mutual information [18, 19, 20], and Kullback-Leibler divergence, also known as the relative entropy, which we recently introduced [21, 22, 23, 24].
 I-A Literature review
-To address these challenges mentioned above, there has been a growing interest in developing unified performance measures that can simultaneously capture both sensing and communication capabilities.
-One such measure that has gained traction is the Kullback-Leibler divergence (KLD), also known as the relative entropy.
-KLD has a well-established history in information theory and has been extensively used in the analysis and design of sensing systems, particularly in the context of MIMO radars [18, 19, 20].
-It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another.
-While the application of KLD in sensing systems is well-established, its potential in characterising communication system performance remains open.
-Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems.
-For instance, in our previous works [21, 22], we employed KLD to analyse performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimising network resources for both separated and shared deployment scenarios.
-Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [24], and more designs have been discovered.
-Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target (MUMT) ISAC system using the promising KLD.
 Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance.
-For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25].
-This approach minimises the CRB of angle estimation for radar while maximising the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously.
+For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25], where this approach minimises the CRB of angle estimation while maximising the signal-to-interference-plus-noise ratio (SINR) for communication.
 However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities.
-Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26].
-While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in deeply integrated scenarios.
-Waveform optimisation has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27].
-This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics.
-However, this work does not consider a unified optimisation framework that could simultaneously address the needs of both subsystems.
+Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid designs to address hardware complexity issues [26].
+While these approaches offer some practical advantages in terms of implementation, they fail to fully capture the nuanced trade-offs between sensing and communication performance in deeply integrated scenarios.
+Similarly, waveform optimisation has also been extensively studied, with recent work using faster-than-Nyquist approaches to enhance spectral efficiency while maintaining radar functionality as a constraint [27].
+Recent studies have also investigated joint transmit and receive beamforming in full-duplex ISAC systems to manage interference and enhance overall performance; however, these approaches primarily emphasise interference suppression rather than exploitation, thereby neglecting potential performance gains achievable through direct integration of communication signals into radar sensing processes [28].
+Despite these advances, existing work has not adequately addressed the need for a unified optimisation framework that simultaneously serves both subsystems’ requirements and balances their performance trade-offs effectively [29].
+Furthermore, a recent comprehensive survey on interference management in ISAC systems highlights that most current strategies remain oriented toward interference mitigation or cancellation, reinforcing the perspective of communication signals as detrimental rather than beneficial [30].
+Consequently, such designs not only overlook opportunities to leverage inherent signal interactions but also risk introducing excessive complexity, which is especially problematic when real-time processing is required in dynamic environments.
 I-B Motivation and contribution
-Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems.
-Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals.
-Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities, improving the system’s overall resource efficiency [14].
-This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
-Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterise and optimise the sensing-communication performance trade-offs in shared antenna deployment ISAC systems.
-This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems.
-By considering the shared deployment scenario and proposing specific optimisation techniques for both subsystem waveforms, we provide a comprehensive approach for system optimisation that reflects the true potential of ISAC systems in future wireless networks.
-Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems to effectively balance the needs of both subsystem functionalities.
-Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources.
-The key contributions of this work are:
-  • Unified Performance Measure: We derive the Kullback-Leibler Divergence (KLD) specifically for the integrated operation of communication and radar systems within ISAC.
-This performance measure unifies the communication error rate with radar target detection capabilities into a single objective function.
-By deriving this KLD-based metric for communication and radar components, we optimise ISAC system design holistically, addressing the trade-offs and enhancing overall system performance.
+Despite these advances, current approaches often fall short of exploiting the full potential of the integrated system by assuming sensing and communication are competing subsystems rather than considering a synergic design.
+Most notably, there is a need for more comprehensive models that accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals.
+Exploiting the interference from communication signals as a source of information for radar operations, as shown in our paper, enhances radar detection capabilities, improving the system’s overall resource efficiency.
+This dual use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[31, 32].
+Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario.
+Additionally, this work investigates the trade-off between trade-offs between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [22].
+Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model.
+To this end, the key contributions of this work are:
+  • Two Novel Separate Waveform Optimisation Designs: we present radar and communication waveform optimisation using the unified KLD measure to represent the achievable performance for both subsystems on one scale.
+Both formulated problems aim at maximising the KLD measure of a certain subsystem subject to the minimum KLD of the other subsystem and power constraints.
+The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method.
+These techniques efficiently handle ISAC waveform design’s complex nature, enabling balanced performance optimisation across both subsystems within system constraints.
+  • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together.
+This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, leading to enhanced performance in terms of radar detection accuracy and communication reliability compared with the separate waveform methods.
+By capturing and exploiting the synergy between sensing and data transmission, the unified technique achieves higher overall performance in shared-antenna ISAC deployments.
   • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation.
-This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection.
-  • KLD-based Optimised ISAC: We propose two novel optimisation techniques: radar and communication waveform KLD-based optimisation.
-Both maximise their respective KLD measures subject to the minimum KLD of both subsystems and power constraints.
-The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method.
-These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimisation across both subsystems within system constraints.
-  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyse their impact on the overall system performance within our KLD-based framework.
-This analysis provides practical insights for optimising ISAC performance in shared deployment scenarios, considering the complexity-performance trade-off.
-Our results demonstrate that the proposed KLD-based optimisation techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design.
-The radar waveform optimisation shows substantial improvements in both target detection and communication performance, while the communication waveform optimisation primarily benefits the communication subsystem with modest radar performance gains.
-Both techniques exhibit robust performance across varying SNR levels, with the radar optimisation demonstrating particularly stable computational efficiency.
-Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources.
-These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
+This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection, which offers new insights into the underlying synergy between radar and communication waveforms and quantifies how their joint optimisation can enhance overall ISAC capabilities.
+  • Comprehensive Complexity Assessment: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility.
+This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
+Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline.
+The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, with modest radar performance gains.
+Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency.
+The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering even higher overall performance in terms of sensing and communication signals.
+The numerical results highlight how leveraging communication signals for radar detection significantly enhances overall system performance, especially in scenarios with constrained dedicated radar resources.
+These findings emphasise the practical significance of KLD-based designs for shared-antenna ISAC systems, provide important adaptability and insights for balancing sensing and communication requirements in future wireless networks.
 I-C Paper organisation
@@
 Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system.
-In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated.
-Section VI introduces the second KLD-based optimisation technique based on communication KLD sum with radar KLD constraint.
-Section VII displays the numerical results, and finally, Section VIII concludes the work.
+In Section V, the radar waveform optimisation is investigated using the KLD.
+Section VI, the communication waveform optimisation is investigated using the KLD.
+Section VII shows the complexity analysis for both waveform optimisation techniques, Section VIII displays the numerical results, and finally, Section IX concludes the work.
 $Notation$: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors.
@@
 The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$.
-As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16].
+As a starting point, non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16].
 These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI.
-The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
+The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
 $\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},$ (1)
-where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem.
+where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem.
 The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS.
@@
 This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information.
-The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots.
+The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$.
 This formulation allows us to incorporate the communication signals into the radar detection process.
@@
 $y_{\mathrm{c},k,l}=\overbrace{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}}^{\text{Desired $k$-th UE signal}}+\overbrace{\omega_{\mathrm{MN},k}}^{\begin{subarray}{c}\text{Inter-user}\\ \text{interference}\end{subarray}}+\overbrace{\eta_{k}}^{\begin{subarray}{c}\text{Radar interference}\\ \text{and noise}\end{subarray}},$ (2)
-where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN).
+where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN).
 The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$.
-In the next two subsections, we revisit the conventional MRT and ZF designs which are typically used to precode the communication data symbols at the BS.
-These precoding schemes will serve as benchmarks for our designs.
-II-A1 MRT beamforming
-In MRT precoding, the amount of received power at the UEs is maximised without taking into account the imposed interference.
-This approach usually leads to acceptable performance at low SNR values but its performance deteriorates at large SNRs.
-The normalised MRT beamforming matrix is designed for all UEs based on the channel matrix as follows,
-$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\>\mathbf{H}^{*}/\|\mathbf{H}\|_{F}=\mathbf{P}\>[\mathbf{h}_{1}^{*},\ldots,\mathbf{h}_{K}^{*}]/\|\mathbf{H}\|_{F},$ (3)
-where $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
-II-A2 ZF beamforming
-On the contrary, ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power.
+ZF designs which is typically used to precode the communication data symbols at the BS.
+This precoding scheme will serve as a benchmark for our designs.
+The ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power.
 This approach usually results in excellent performance at high SNR values.
@@
 $\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (3)
-where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$ is the ZF beamformer before normalisation.
+where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
 II-B Radar system
@@
 However, it is crucial to acknowledge that the value of $T$ is constrained by the available antennas and UEs serviced by the BS, highlighting the need for efficient waveform design and resource allocation in ISAC systems.
-Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [32, 33].
-The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [34].
+Our focus aligns with scenarios where targets are spatially separated, with each target confined to a distinct radar bin, as per prior studies [35, 36]  [^1] The literature presents algorithms to separate signals from closely spaced or unresolved targets, facilitating accurate target enumeration [37]..
 The total radar return signal at the $l$-th snapshot is shown as follows,
-$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},$ (5)
-where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix of the BS-Target-BS route for the $t$-th target, following a Rayleigh distribution [35].
-The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
-To analyse the radar returns from all beams, a series of matched filters are typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin.
-Because $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target.
-The radar precoding matrix $\mathbf{W}_{r,l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$.
-The received signal for the $t$-th target at the BS, under the binary hypotheses $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ signifies target presence and $\mathcal{H}_{0}$ denotes target absence, can be mathematically represented as follows,
+$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vspace{-0.05 in}$ (4)
+where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix that captures the complete BS-Target-BS path characteristics for the $t$-th target, inherently incorporating propagation effects, target cross-section, and reflection coefficients.
+The target response matrix follows a Rayleigh distribution to model the composite effects of multiple scattering centres inherent in extended targets, thereby characterising the statistical nature of target scattering while preserving the direct path essential for target detection [38, 39].
+The $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot.
+To process the radar returns from all beams, a bank of matched filters is typically applied, each tuned to a specific waveform $s_{\mathrm{r},t}:\forall t=\left\{1,2,\cdots,T\right\}$, which corresponds to a particular radar angular-range-Doppler bin.
+Assuming $s_{\mathrm{r},t}$ is orthogonal to $s_{\mathrm{r},i}$ for all $t\neq i$ using signal space and time domain, the radar returns from distinct beams can be effectively separated, enabling independent detection of each potential target [39, 34].
+The radar precoding matrix $\mathbf{W}_{\mathrm{r},l}$ is designed such that the covariance matrix is $\mathbf{R}_{w}\triangleq\frac{1}{L}\sum_{l=1}^{L}\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}\in\mathbb{C}^{N\times N}$.
+The received signal for the $t$-th target at the BS after match-filtering, under the binary hypothesis testing problem $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ represents the target presence whereas $\mathcal{H}_{0}$ denotes the target absence, is denoted by $\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}$ and can be mathematically represented as follows,
 $\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}\>q+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}\>q+\mathbf{n}_{t,l},$ (5)
-The $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the additive white Gaussian noise vector at the $l$-th snapshot.
-III The communication KLD: ZF and MRT
-In this section, we derive the KLD of the communication system for the MRT and ZF beamforming using the shared deployment antenna configuration.
+where $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot at the output of the matched filter.
+III The communication KLD: ZF and Arbitrary precoding matrix
+In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver.
 The analysis is carried out for the normalised precoders introduced in Section II-A.
@@
 $\mathrm{KLD}_{n\rightarrow m}=\frac{1}{2\ln 2}\left(\mathrm{tr}\left(\Sigma_{n}^{-1}\Sigma_{m}\right)-2+\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)^{T}\right.\\ \;\left.\times\;\Sigma_{n}^{-1}\left(\mathbf{\mu}_{n}-\mathbf{\mu}_{m}\right)+\ln\frac{\left|\Sigma_{n}\right|}{\left|\Sigma_{m}\right|}\right).$ (6)
-III-A MRT beamforming
-For the MRT precoding scheme, the analysis encompasses the desired signal, IUI, radar interference, and noise components.
-The examination starts with the IUI term, which is the second term in (2), given by,
-$\omega_{\mathrm{MN},k}=d_{\mathrm{c},k}^{-\zeta/2}\mathbf{h}_{k}^{T}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}=\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\|\mathbf{H}\|_{F}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (8)
-Since the elements of $\mathbf{h}_{k}$ and $\mathbf{h}_{i}$ $\forall i\neq k$ are i.i.d.
-complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$, the term $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ is the inner product of two independent complex Gaussian vectors.
-Thus, the central limit theorem (CLT) can be invoked to show that the distribution of $\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}$ tends to be complex Gaussian with zero mean and variance $N\sigma_{h}^{4}$ for a large number of antennas $N$.
-Meanwhile, the normalisation term $\|\mathbf{H}\|_{F}^{2}=\sum_{i=1}^{K}\|\mathbf{h}_{i}\|^{2}$ follows a chi-squared distribution with $2NK$ degrees of freedom, scaled by $\sigma_{h}^{2}/2$.
-Consequently, the exact distribution of $\omega_{\mathrm{MN},k}$ is a ratio of complex normal variables to the square root of a chi-squared variable.
-This ratio distribution does not have a simple closed-form expression.
-However, we can approximate its behaviour for large $N$ and $K$, by using the law of large numbers, $\|\mathbf{H}\|_{F}^{2}\approx NK\sigma_{h}^{2}$.
-Using this approximation, we can write (8) as,
-$\omega_{\mathrm{MN},k}\approx\frac{d_{\mathrm{c},k}^{-\zeta/2}}{\sqrt{NK\sigma_{h}^{2}}}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\mathbf{h}_{k}^{T}\mathbf{h}_{i}^{*}\sqrt{p_{i}}s_{\mathrm{c},i,l}.$ (9)
-$\omega_{\mathrm{MN}}$ is a sum of independent complex normal variables. By invoking CLT, for large $KN$, the sum approaches a complex normal distribution with zero mean and a variance of
-$\sigma_{\omega,\mathrm{MN}}^{2}\approx\frac{d_{\mathrm{c},k}^{-\zeta}}{NK\sigma_{h}^{2}}\sum_{i\neq k}N\sigma_{h}^{4}p_{i}=\frac{d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}}{K}\sum_{i\neq k}p_{i},$ (10)
-The accuracy of the approximated density of $\omega_{\mathrm{MN}}$ is validated by simulation, which compares it with simulated density functions under different values of $N$.
-It can also be shown that the distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$.
-This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
+III-A ZF beamforming
+With ZF precoding, the received signal at the $k$-th user is,
+$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l},$ (7)
+where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix.
+Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI.
+The distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$.
+This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
 $\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (8)
 where $\mathbf{R}_{w,t}\!\!=\!\!\frac{1}{L}\!\sum_{l=1}^{L}\!\!\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}^{H}$.
-Now, the equivalent IUI and radar interference plus noise term $\tilde{\omega}_{\mathrm{MN},k}=\omega_{\mathrm{MN}}+\eta_{k}$ can be approximated as a complex Gaussian random variable with zero mean and variance $\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2}$.
-In order to find the KLD for the MRT case with matrix normalisation, we need to evaluate the expected value of the desired signal power, which is represented by the first term in (2).
-This term, $|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$ can be expressed as follows,
-$|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}=|\mathbf{h}_{k}^{T}\mathbf{h}_{k}^{*}|^{2}/\|\mathbf{H}\|_{F}^{2}=\|\mathbf{h}_{k}\|^{4}/\|\mathbf{H}\|_{F}^{2}.$ (12)
-The expected value of this term can be approximated using the properties of the chi-squared distribution and the approximation $\mathbb{E}[1/X]\approx 1/\mathbb{E}[X]$ for a random variable $X$,
-$\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]\approx p_{k}\mathbb{E}[\|\mathbf{h}_{k}\|^{4}]/\mathbb{E}[\|\mathbf{H}\|_{F}^{2}]=p_{k}(N+1)\sigma_{h}^{2}/K.$ (13)
-Given that $\Sigma_{n}=\Sigma_{m}=(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})\mathbf{I}_{2}$, and $\mathbf{\mu}_{k,m}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(m)}$, and $\mathbf{\mu}_{k,n}=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|]s_{\mathrm{c},k,l}^{(n)}$, substituting these into equation (7), then the KLD can be found as [21, Corollary 1],
-$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>\mathbb{E}[|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}]}{(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2},$ (14)
-where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is constant for a given constellation, for example, for MPSK signalling, $\lambda_{\mathrm{MPSK}}=2\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left(1-\cos\left(\phi_{k,m}-\phi_{k,n}\right)\right)$.
-Substituting (13) in (14), the KLD for MRT precoding with matrix normalisation is,
-$\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{MRT}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N+1)\sigma_{h}^{2}}{K\>(\sigma_{\omega,\mathrm{MN}}^{2}+\sigma_{\eta}^{2})M(M-1)\ln 2}.$ (15)
-III-B ZF beamforming
-With ZF precoding, the received signal at the $k$-th user is,
-$y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}}+n_{k,l},$ (16)
-where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix.
-Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI.
 The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
 $\boldsymbol{\mu}_{k,j}$ $=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]s_{\mathrm{c},k,l}^{(j)}$
-$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[|\mathbf{\tilde{W}}_{\mathrm{c}}|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (17)
+$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (9)
 $\ \Sigma_{n}$ $=\Sigma_{m}=\sigma_{\eta}^{2},$ (10)
@@
 Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$.
-By substituting (17), and (18) in (7), then the KLD becomes,
+By substituting (9), and (10) in (6), and considering all possible pairs of dissimilar symbols [22, Corollary 1], then the KLD becomes,
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}}{M(M-1)\sigma_{\eta}^{2}\ln 2}\mathbb{E}\left[\frac{1}{\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}}\right],$ (11)
+where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ is a constant that depends on the constellation.
 The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
@@
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}(N-K)/M(M-1)\sigma_{\eta}^{2}\ln 2.$ (13)
-III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
+III-B Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
 A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem.
@@
 $=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\quad j\in\{n,m\},$ (15)
-Note that, the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
-$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$ (24)
-Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
-$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
-×$\times\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
-where $\lambda=\sum_{m=1}^{M}\sum_{\begin{subarray}{c}n=1\\ n\neq m\end{subarray}}^{M}\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}$ and $\left|s_{\mathrm{c},k,l}^{(n)}-s_{\mathrm{c},k,l}^{(m)}\right|^{2}=\lambda/(M(M-1))$, substituting it in the KLD expression which becomes,
-$\begin{split}\mathrm{KLD}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (26)
-It should be noted that the second term in this equation, represents the form of the SINR, where the interference is the sum of IUI and radar interference.
+Note that, the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
+$\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},$ (16)
+by substituting the evaluated mean vectors and covariance matrices in (14), and (16) into the KLD expression given in (6), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, which can be represented as follows,
+$\!\!\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}\!=\!\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
+×$\!\!\times\!\Bigg{(}\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg{)}^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg{)}\!.\!\!$ (17)
+Thereafter, by considering all possible pairs of dissimilar symbols, then the KLD expression using $\lambda$ is simplified as,
+$\begin{split}\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (18)
+It should be noted that the second term in this equation, represents a form of the SINR, where the interference is the sum of IUI and radar interference.
 IV Radar system analysis
 IV-A Composite detection and response matrix estimation
-The main objective of typical radar systems is to detect targets and estimate the matrix response for each one.
+The main objective of typical radar systems is to detect the existence of targets and estimate the matrix response for existing targets.
 After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
-$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{q}).$ (27)
-This joint optimisation problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present.
-For the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis can be shown as a multivariate normal density,
-$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
+$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{q}).$ (19)
+This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining ^$\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if presents.
+It can be observed from (5) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
+$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t,l}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
 ×$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
 where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$.
-The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, derived as,
-$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (29)
-As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain,
-$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1}).$ (30)
-The likelihood function is Gaussian, and the maximisation leads to the least squares solution,
-$\hat{\mathbf{H}}_{t}$ $=\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}-\mathbf{H}_{t}\mathbf{x}_{t}\right\|^{2}=\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t}^{H}\left(\mathbf{x}_{t}\mathbf{x}_{t}^{H}\right)^{-1}.$ (31)
-This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain,
+The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, which is derived as,
+$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (21)
+As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem, we obtain,
+$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1}).$ (22)
+Since the PDF of $\mathbf{y}_{\mathrm{r},t,l}$ is multivariate Gaussian, the maximisation leads to the least squares solution given by,
+$\hat{\mathbf{H}}_{t}\!$ $=\!\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}\!-\!\mathbf{H}_{t}\mathbf{x}_{t,l}\right\|^{2}\!=\!\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t,l}^{H}\left(\mathbf{x}_{t,l}\mathbf{x}_{t,l}^{H}\right)^{-1}.$ (23)
+This least squares solution provides an unbiased estimate of the target‐response matrix $\mathbf{H}_{t}$, with accuracy directly tied to the detection probability $P_{D}$.
+When the radar correctly detects a target under $\mathcal{H}_{1}$, the maximum likelihood solution $\hat{\mathbf{H}}_{t}$ is used to estimate $\mathbf{H}_{t}$.
+In contrast, a miss detection (i.e., deciding $\mathcal{H}_{0}$ under $\mathcal{H}_{1}$) forces $\hat{\mathbf{H}}_{t}=\mathbf{0}$, resulting in significant errors.
+The overall unconditional MSE is defined as,
+$\mathrm{MSE}$ $=p(H_{1})\Bigl{[}(1-P_{D})\,\mathrm{MSE}_{0\_1}+P_{D}\,\mathrm{MSE}_{1\_1}\Bigr{]}$
+$+\bigl{[}1-p(H_{1})\bigr{]}\Bigl{[}P_{\mathrm{FA}}\,\mathrm{MSE}_{1\_0}\Bigr{]},$ (24)
+[…]
+Substituting the PDF in (20) into the GLRT, we obtain,
 $\Lambda(\mathbf{y}_{\mathrm{r},t,l})=$
-$\frac{\pi^{N}\sigma_{n}^{2N}\exp\left(-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\right)}{\pi^{N}\det(\hat{\mathbf{R}}_{1,t})\exp\left(-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\right)}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{0}$ (32)
+$\frac{\!\sigma_{n}^{2N}\!\exp\!\left(\!-(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})\!\right)}{\!\det(\hat{\mathbf{R}}_{1,t})\exp\!\left(\!-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\!\right)}\!\!\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\!\tau_{0}$ (25)
 where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold.
-Taking the logarithm and rearranging, we obtain,
-$\ln\Lambda(\mathbf{y}_{\mathrm{r},t,l})$ $=\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}-N\ln\sigma_{n}^{2}+\ln\det(\hat{\mathbf{R}}_{1,t})$
-$\!-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\ln\tau_{0}$ (33)
-The terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ are constant with respect to $\mathbf{y}_{\mathrm{r},t,l}$, so they can be absorbed into the threshold, which further simplifies our GLRT as follows,
+Taking the logarithm of the likelihood ratio and rearranging terms, the final GLRT form is given by,
 $\dot{\Lambda}(\mathbf{y}_{\mathrm{r},t,l})=\mathbf{y}_{\mathrm{r},t,l}^{H}\left(\mathbf{I}_{N}+\hat{\mathbf{R}}_{1,t}^{-1}\right)\mathbf{y}_{\mathrm{r},t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{1}$ (26)
 where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$.
-This expression compares hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$.
-To use all $L$ snapshots with independent signals, we average as,
-$\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}.$ (35)
-where $\tau_{2}$ is the detection threshold considering all snapshots and can be designed per the Neyman-Pearson lemma.
+Here, the constant terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ have been absorbed into the threshold.
+This expression decides the most probable hypothesis, i.e., $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$, by incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$.
+Since the received signals from different snapshots are conditionally independent, (26) can be updated to adapt to $L$ snapshot scenario by taking the average, i.e., $\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}$, with $\tau_{2}$ is the detection threshold considering all snapshots and can be designed according to Neyman-Pearson lemma.
 IV-B The KLD of radar system
@@
 $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
+The average KLD of the whole ISAC system can be formulated as follows,
+$\mathrm{KLD}_{\mathrm{ISAC}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}+\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}.$ (28)
 V Radar Waveform Optimisation
@@
 $\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
-s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (37b)
-$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (37c)
+s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (29b)
+${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (29c)
 $\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{r}},$ (29d)
@@
 This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance.
+The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively.
 The constraint on radar power (29d) is explicit because the radar waveform is our primary optimisation variable.
@@
 $\mathcal{P}_{2}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}$ $\!\!\!\!\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\ln{2}}\Bigg{(}\!\ln\!\left(\mathrm{det}(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N})\!\right)$ (30)
-$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)-N\>\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\Bigg{)}$
+$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\!\sigma_{n}^{2}\!\right)\!-\!N\!\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\!\Bigg{)}$
 s.t. (29b), (29c), (29d), and (29e)
@@
 where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function.
-The gradients of $\mathrm{KLD_{r}}$ and $\mathrm{KLD_{c}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A.
+The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A.
 The final forms are given by,
@@
 $-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big{)}_{i},$ (35)
-$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}=$ $-\frac{2\>d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (44)
+$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}\!=\!$ $-\!\frac{2d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N\!\!-\!\!K)\sigma_{h}^{2}}{M(M\!\!-\!\!1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\!\mathrm{r}}]_{i,l,t}.$ (36)
 The update rule for the projected gradient method with penalty ensures that each iteration moves the solution towards optimality while maintaining feasibility, and can be shown as follows,
@@
 $\mathcal{P}(\mathbf{\dot{W}}_{\mathrm{r}})=\begin{cases}\mathbf{\dot{W}}_{\mathrm{r}}\sqrt{\frac{P_{\mathrm{r}}L}{\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}}},&\text{if }\frac{1}{L}\|\mathbf{\dot{W}}_{\mathrm{r}}\|_{F}^{2}>P_{\mathrm{r}}\\ \mathbf{\dot{W}}_{\mathrm{r}},&\text{otherwise}\end{cases}$ (38)
-The step size $\alpha_{n}$ is determined using a backtracking line search to ensure convergence. We find the smallest non-negative integer $m$ such that,
-$\begin{split}&f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\geq\\ &f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})+c\alpha_{n}\|\mathbf{G}^{(n)}\|_{F}^{2},\end{split}$ (47)
-where $c\in(0,1)$.
-This condition, known as the Armijo condition, ensures a sufficient decrease in the penalised objective function [38].
-The step size is set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, with $\beta\in(0,1)$ as a reduction factor.
-The penalty parameter $\rho_{n}$ is updated as follows,
-$\rho_{n+1}=\begin{cases}\gamma\rho_{n},&\text{if }p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})>0\\ \rho_{n},&\text{otherwise}\end{cases}$ (48)
-where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied.
 Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimisation
@@
 14:  return  $\mathbf{\dot{W}}_{\mathrm{r}}^{*}=\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}$
+The step size $\alpha_{n}$ is determined using a backtracking line search to ensure convergence. We find the smallest non-negative integer $m$ such that,
+$\begin{split}&f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\geq\\ &f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})+c\alpha_{n}\|\mathbf{G}^{(n)}\|_{F}^{2},\end{split}$ (39)
+where $c\in(0,1)$.
+This condition, known as the Armijo condition, ensures a sufficient decrease in the penalised objective function [42].
+The step size is set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, with $\beta\in(0,1)$ as a reduction factor.
+The penalty parameter $\rho_{n}$ is updated as follows,
+$\rho_{n+1}=\begin{cases}\gamma\rho_{n},&\text{if }p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})>0\\ \rho_{n},&\text{otherwise}\end{cases}$ (40)
+where $\gamma>1$ is a constant factor.
+This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied.
 The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied.
@@
 The line search condition can be expressed as,
-$\begin{split}\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\leq&\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})\\ &+c\alpha_{n}\langle\nabla\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}),\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\end{split}$ (49)
+$\begin{split}\!\!\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{(n+1)})\!\!\leq&\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{\!(n)})\!+\!c\alpha_{n}\!\langle\nabla\!\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{\!(n)}),\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}\!\!-\!\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\!\!\!\end{split}$ (41)
 where $\phi(\mathbf{\dot{W}}_{\mathrm{\!r}})\!=\!\!f(\mathbf{\dot{W}}_{\mathrm{\!r}})\!-\!\!\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{\!r}})$ is the penalised objective function.
@@
 The communication subsystem optimisation uses the gradient-assisted IPM [43], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures.
-Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$.
+Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$.
 This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions.
-The communication waveform is optimised while using the conventional identity covariance matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$.
+The communication waveform is optimised while using the conventional identity covariance (CIC) matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$.
 The optimisation problem can be expressed as follows,
-$\mathcal{P}_{3}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (50a)
-s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (50b)
-$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (50c)
+$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (42a)
+s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (42b)
+${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (42c)
 $\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (42d)
@@
 Using the chain rule and properties of matrix derivatives, the gradients can be expressed as follows,
-$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\quad\text{for all }j$ (52)
+$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\quad\quad\quad\quad\forall j$ (44)
 $\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ =$=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (45)
@@
 Next, we need to consider the constraints of our optimisation problem. The power constraint is given by,
-$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.$ (55)
+$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vspace{-0.05 in}$ (47)
 The gradient of this constraint is derived as follows,
-$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.$ (56)
+$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vspace{-0.05 in}$ (48)
 The gradient of the radar KLD constraint with respect to $\mathbf{W}_{\mathrm{c}}$ is derived as follows,
-$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).$ (57)
+$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vspace{-0.03 in}$ (49)
 Algorithm 2 Interior Point Method for Communication Optimisation
-0:  Initial point $\mathbf{W}_{c}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
-0:  Optimal solution $\mathbf{W}_{c}^{*}$
+0:  Initial point $\mathbf{W}_{\mathrm{c}}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
+0:  Optimal solution $\mathbf{W}_{\mathrm{c}}^{*}$
 1:  Initialise $n=0$, $\mu=\mu_{0}$
 2:  while $n<\mathrm{max\_iter}$ do
-3:     Compute $f(\mathbf{W}_{c}^{(n)})$, $\nabla f(\mathbf{W}_{c}^{(n)})$, $g_{i}(\mathbf{W}_{c}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{c}^{(n)})$
+3:     Compute $f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $\nabla f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$
 4:     Compute KKT conditions
@@
 7:     end if
-8:     Solve Newton system for search direction $\Delta\mathbf{W}_{c}$
+8:     Solve Newton system for search direction $\Delta\mathbf{W}_{\mathrm{c}}$
 9:     Perform line search to find step size $\alpha$
-10:     Update $\mathbf{W}_{c}^{(n+1)}=\mathbf{W}_{c}^{(n)}+\alpha\Delta\mathbf{W}_{c}$
+10:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}=\mathbf{W}_{\mathrm{c}}^{(n)}+\alpha\Delta\mathbf{W}_{\mathrm{c}}$
 11:     Update $\mu=\gamma\mu$
@@
 13:  end while
-14:  return  $\mathbf{W}_{c}^{*}=\mathbf{W}_{c}^{(n)}$
+14:  return  $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
 Solving the optimisation problem defined in equations (42a)-(42d) uses the IPM, as detailed in Algorithm 2.
@@
 The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly.
-This continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
-VII Numerical Results
-In this section, we present simulation results for our conventional benchmarks and the two previously introduced optimisation techniques.
-The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout.
-The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets, and $K=3$ UEs.
-The pathloss exponent is $\eta=3$ to model large-scale fading.
+This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
+VII Integrated Waveform Optimisation for ISAC
+Finally, we introduce this unified approach that maximises the whole ISAC system KLD, which ensures improved radar detection and communication reliability while satisfying minimum subsystem KLD constraints and adhering to a predefined power budget.
+The optimisation problem is expressed as,
+$\mathcal{P}_{1}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\quad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (50a)
+s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t\in T,$ (50b)
+$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k\in K,$ (50c)
+$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (50d)
+$P_{\mathrm{T}}=P_{\mathrm{r}}+P_{\mathrm{c}}.$ (50e)
+The total power constraint in (50d) limits the combined radar and communication power to $P_{\mathrm{T}}$.
+To solve this non-convex problem, the ADMM is employed.
+This approach unifies the projected gradient method used for radar waveform optimisation and the IPM technique used for communication waveform optimisation, efficiently handling the coupling between subsystems while decomposing the complex problem into more manageable subproblems.
+This involves introducing auxiliary variables $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ such that $\mathbf{Z}_{\mathrm{r}}=\mathbf{W}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}=\mathbf{W}_{\mathrm{c}}$.
+The augmented Lagrangian for the problem is given as,
+$\mathcal{L}_{\rho}=$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}+\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}$
+$+\langle\mathbf{U}_{\mathrm{r}},\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\rangle+\langle\mathbf{U}_{\mathrm{c}},\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\rangle$
+$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),$ (51)
+where $\mathbf{U}_{\mathrm{r}}$ and $\mathbf{U}_{\mathrm{c}}$ are the dual variables (Lagrange multipliers), and $\rho>0$ is the ADMM penalty parameter.
+The ADMM algorithm proceeds iteratively, updating the primal variables $(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$, auxiliary variables $(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$, and dual variables $(\mathbf{U}_{\mathrm{r}},\mathbf{U}_{\mathrm{c}})$.
+The updates are given as follows,
+$\mathbf{W}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{r}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{r}}^{(n)},$ (52)
+$\mathbf{W}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{c}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{c}}^{(n)},$ (53)
+$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\quad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (54)
+$\mathbf{U}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{r}}^{(n)}+\rho(\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}),$ (55)
+$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).$ (56)
+The auxiliary variables are updated by solving the following constrained subproblem,
+$\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\quad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (57)
+s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t,$ (58)
+$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k,$ (59)
+$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},$ (60)
+where $p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$ is a penalty term that ensures constraints are satisfied.
+Since the subproblem (57) may not have a closed-form solution, it is solved approximately using projected gradient descent.
+The auxiliary variables are iteratively updated as: $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathbf{Z}_{\mathrm{r}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{r}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{r}}}p\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathbf{Z}_{\mathrm{c}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{c}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{c}}}p\right)$, where $\rho_{n}$ is the penalty parameter at iteration $n$, and $\alpha_{n}$ is the step size determined by a backtracking line search.
+The gradients of $f$ with respect to $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ are computed as,
+$\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (61)
+$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).$ (62)
+The closed-form expressions for $\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ and $\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ are given by,
+$\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (63)
+$\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ $=\frac{2\lambda d_{\mathrm{c},k}^{-\zeta}}{M(M-1)\ln 2}\Bigg{[}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},k}}{\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}}$
+$\quad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg{]}.$ (64)
+After each gradient-based update, the auxiliary variables are projected onto the feasible set defined by the power constraint as, $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right)$, where $\mathcal{P}(\cdot)$ denotes the projection operator ensuring that (60) is satisfied.
+This iterative procedure continues until convergence criteria, such as primal and dual residual thresholds, are met.
+Algorithm 3 ADMM for Unified ISAC Waveform optimisation
+0:  Initial variables $\mathbf{W}_{\mathrm{r}}^{(0)}$, $\mathbf{W}_{\mathrm{c}}^{(0)}$, $\mathbf{Z}_{\mathrm{r}}^{(0)}$, $\mathbf{Z}_{\mathrm{c}}^{(0)}$, $\mathbf{U}_{\mathrm{r}}^{(0)}$, $\mathbf{U}_{\mathrm{c}}^{(0)}$, parameters $\rho$, $\alpha_{0}$, $\rho_{0}$, tolerance $\varepsilon$, max iterations $\mathrm{max\_iter}$
+0:  Optimized waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
+1:  Initialize $n=0$
+2:  while $n<\mathrm{max\_iter}$ do
+3:     Update $\mathbf{W}_{\mathrm{r}}^{(n+1)}$ using (52)
+4:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}$ using (53)
+5:     Solve optimisation sub-problem (57) to update $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ using the projected gradient descent
+6:     Project $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ onto feasible set
+7:     Update $\mathbf{U}_{\mathrm{r}}^{(n+1)}$ using (55)
+8:     Update $\mathbf{U}_{\mathrm{c}}^{(n+1)}$ using (56)
+9:     if $\left\|\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right\|_{F}<\varepsilon$ and $\left\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right\|_{F}<\varepsilon$ then
+10:        break
+11:     end if
+12:     $n=n+1$
+13:  end while
+14:  return  $\mathbf{W}_{\mathrm{r}}^{*}=\mathbf{W}_{\mathrm{r}}^{(n)}$, $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
+VIII Complexity Analysis
+VIII-1 Projected Gradient Method with Penalty
+The projected gradient method with adaptive penalties for radar waveform optimisation, defined in Algorithm 1, has an estimated per-iteration complexity as [42, 41],
+$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),$ (65)
+where $I$ is the number of iterations, $E$ denotes the cost of evaluating the objective function and its gradients, $n$ represents the number of optimisation variables in the tensor $\dot{\mathbf{W}}_{\mathrm{r}}$, $m$ is the number of constraints, and $k$ is the number of backtracking line search iterations per step.
+The dominant $n^{3}$ term arises from matrix operations in the projection step.
+VIII-2 Gradient-assisted IPM
+For communication waveform optimisation, we utilise a gradient-assisted Interior Point Method as outlined in Algorithm 2. Its per-iteration complexity is [44],
+$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),$ (66)
+where $n$ denotes the number of optimisation variables in the matrix $\mathbf{W}_{\mathrm{c}}$. Again, while the dominant $n^{3}$ term reflects matrix operations.
+VIII-3 ADMM
+The ADMM algorithm for unified optimisation defined in Algorithm 3, has per-iteration complexity as [45],
+$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),$ (67)
+where $n$ represents the combined number of optimisation variables, and $J$ is the number of inner projected gradient iterations per ADMM step.
+The dominant $n^{3}$ term arises from matrix operations in the projection steps.
+IX Numerical Results
+In this section, we present simulation results for our conventional benchmarks and the three previously introduced optimisation techniques.
+The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section.
+The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed.
+A pathloss exponent of $\eta=3$ is considered to model large-scale fading.
 The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters.
-The channel variance is $\sigma_{h}^{2}=1$.
-VII-A Trade-off for the conventional benchmarks
-The radar subsystem employs a conventional identity covariance matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced MRT and ZF.
-Figure 1: Performance of the communication subsystem for both beamforming schemes, showing BER and achievable $\mathrm{KLD}_{\mathrm{c}}$ versus $P_{\mathrm{R}}/{N}_{o}$ with $P_{\mathrm{r}}=0.5$.
+The channel variance is fixed at $\sigma_{h}^{2}=1$, and the probability of false alarm $P_{FA}$ is set to $10^{-4}$.
+$\overline{C}_{k}$ is set to $210^{-}$ 2 which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits.
+IX-A Trade-off for the conventional benchmark
+The radar subsystem employs a CIC matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced ZF.
+Figure 1: The ISAC system performance at $P_{\mathrm{r}}=0.5$.
+a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
+[…]
+IX-B Radar Waveform KLD-based Optimisation
+The radar waveform KLD-based optimisation algorithm is set with a maximum number of iterations of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$.
+The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
+Figure 2: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using radar waveform optimisation: K-ROP.
+a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-1 shows the communication subsystem performance for both ZF and MRT beamforming.
+2 illustrates the performance of both subsystems using the KLD-based radar waveform optimisation (K-ROP) technique.
 In Fig.
-1.a, the MRT scheme’s performance is illustrated, with KLD on the right $y$-axis and BER on the left.
-The KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ up to about $10$ dB, where an upper bound appears due to radar interference, with IUI contributing less significantly.
-Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, reflecting improved communication performance but also reaches an error floor around $10$ dB.
+2.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ on the right axis and detection probability $P_{D}$ on the left axis versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
+As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance.
+By comparing the results here with the ZF-based design in Fig.
+1, it can be noticed that there is a significant gain can be achieved using K-ROB.
+For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig.
+1, demonstrating improved detection capabilities through K-ROP optimisation technique.
+For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system.
 Fig.
-1.b presents the ZF scheme’s performance, showing similar KLD and BER trends.
-However, the ZF scheme exhibits higher KLD values and lower BER values for the same received SNR compared to MRT, confirming that radar interference is the primary source, as IUI is cancelled using ZF precoding.
-Both schemes display a significant error floor in BER and an upper bound in KLD due to persistent radar interference at $P_{\mathrm{c}}=0.5$.
-This figure effectively demonstrates the utility of KLD in characterising the communication system.
-Figure 2: The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
-Fig.
-2 presents the receiver operating characteristic (ROC) curves for the radar subsystem’s performance across three potential targets at varying $P_{\mathrm{R}}/{N}_{o}$ levels, using ZF precoding for the communication subsystem as shown in Fig.
+2.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ on the right axis and BER on the left axis versus $P_{\mathrm{R}}/N_{o}$.
+For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB suggests it has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced.
+The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted from Fig.
 1.
-The ROC curves plot the probability of detection ($P_{D}$) against the probability of false alarm ($P_{FA}$) for each target.
-At $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are closely aligned, indicating similar detection performance at low SNR.
-However, at $P_{\mathrm{R}}/{N}_{o}=10$ dB, the curves diverge, with $\mathrm{T}_{3}$ showing the best performance, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$.
-This trend continues at $P_{\mathrm{R}}/{N}_{o}=20$ dB, where the separation is even more pronounced.
-Notably, at $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively.
-$\mathrm{T}_{3}$ remains the most detectable, while both $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant detection improvements compared to lower $P_{\mathrm{R}}/{N}_{o}$ scenarios.
-This illustrates that higher SNR enhances the radar subsystem’s capability to detect and differentiate targets effectively.
-Figure 3: a) Radar subsystem performance: achievable $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. b) KLD trade-off between subsystems for ZF and MRT.
-In Fig.
-3.a, the radar subsystem performance is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$.
-As $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ also improves, indicating enhanced radar performance.
-The performance curves for different $P_{\mathrm{r}}$ values are close, suggesting that the system is less sensitive to $P_{\mathrm{r}}$ changes due to the communication signal’s incorporation in target detection using our improved system model.
-Fig.
-3.b presents the trade-off performance between both subsystems using ZF and MRT for communication, along with the conventional identity covariance design for radar.
-The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value going higher from left to right.
-The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ slightly outperforms $\mathrm{KLD_{c}}^{\mathrm{MRT}}$, both exhibiting saturation at higher $P_{\mathrm{R}}/N_{o}$ due to radar interference, while $\mathrm{KLD_{r}}$ steadily increases along the curve.
-Notably, a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits is achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ requires around $25$ dB, implying an SNR gain of about $13$ dB in favor of ZF.
-VII-B Radar Waveform KLD-based Optimisation
-The radar waveform KLD-based optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$.
-The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
-Figure 4: The performance of both subsystems using optimised radar waveform.
-a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
-Fig.
-4 illustrates the performance of both subsystems with the optimised radar waveform.
-In Fig.
-4.a, the achievable $\mathrm{KLD_{r}}$ for each target versus $P_{\mathrm{R}}/N_{o}$ is shown.
-As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance.
-$\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS.
-At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, indicating consistent performance enhancement in the radar system.
-Fig.
-4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$.
-For $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB, $\mathrm{KLD_{c}}$ exhibits minimal improvement, indicating system infeasibility in this low power regime due to the inability to meet minimum KLD requirements for all UEs.
-$\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$.
-At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB suggests the system has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced.
-These results highlight the intricate relationship between radar detection capabilities and communication system performance, illustrating how both subsystems benefit once the feasibility threshold is surpassed, with the performance reflecting the respective distances from the BS.
-Figure 5: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
-Fig.
-5.a presents the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
-At $P_{\mathrm{R}}/{N}_{o}=20$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.909,0.804,0.935\}$ in the non-optimal case, demonstrating improved detection capabilities through optimisation.
-Fig.
-5.b shows the BER performance for the three UEs versus $P_{\mathrm{R}}/N_{o}$.
-As $P_{\mathrm{R}}/N_{o}$ increases, BER decreases for all UEs, with $\mathrm{UE}_{3}$ achieving the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
-The communication subsystem significantly benefits from radar waveform optimisation, as evidenced by the non-optimal case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig.
-1, while optimised waveforms reduce $P_{e}$ below $10^{-5}$ by $13$ dB, effectively mitigating radar interference.
-Figure 6: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
-Fig.
-6.a illustrates the trade-off performance between both subsystems using the optimised radar waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values.
-As $P_{\mathrm{R}}/N_{o}$ increases, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems.
-This highlights the robustness of the optimisation technique in balancing radar and communication performance.
-In comparison, the non-optimal approach in Fig.
-3.b saturates at around $11$ bits, indicating an increasing and favourable trade-off for the optimal way.
-Fig.
-6.b presents the average CPU time of the radar waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$.
-The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes, indicating consistent computational efficiency across varying $P_{\mathrm{R}}/N_{o}$ values, which is beneficial for practical implementation in diverse channel conditions.
-VII-C Communication Waveform KLD-based Optimisation
+These results highlight the intricate relationship between radar detection capabilities and communication system performance.
+IX-C Communication Waveform KLD-based Optimisation
 The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
@@
 The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
+Figure 3: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using communication waveform optimisation: K-COP.
+a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-7 illustrates the performance of the radar and communication subsystems using the optimised communication waveform.
+3 illustrates the performance of both subsystems using the KLD-based communication waveform optimisation (K-COP) technique.
+In Fig.
+3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ (right axis) and detection probability $P_{D}$ (left axis) versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
+For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform.
+At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig.
+1.
+K-COP demonstrates improved detection capabilities over non-optimisation but remains below the radar waveform optimisation performance achieved using K-ROP in Fig.
+2, which is $\{98,89.2,99.9\}\%$.
 Fig.
-7.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for power allocations of $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$.
-As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$.
-For example, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$, increasing to $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The linear improvement in $\mathrm{KLD_{r}}$ beyond $15$ dB indicates stable radar performance despite the suboptimal waveform.
+3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ (right axis) and BER (left axis) versus $P_{\mathrm{R}}/N_{o}$.
+For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
+For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig.
+1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs.
+K-COP optimisation technique provides significant performance gains over the non-optimised ZF scenario, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec.
+IX-B due to radar interference.
+The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
+IX-D Integrated Waveform KLD-based Optimisation for ISAC
+The integrated waveform optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, $\rho=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$.
+The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
+Figure 4: The ISAC system performance using the integrated waveform optimisation.
+a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-7.b depicts the communication subsystem performance with $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$.
-$\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
-At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$, rising to $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem thresholds with communication enhancements across different power allocations.
-Figure 7: The performance of both subsystems using optimised communication waveform.
-a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
+4 illustrates the ISAC system performance under the KLD-based integrated waveform optimisation (K-IOP).
+In Fig.
+4.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ (left axis) and $\mathrm{KLD}_{\mathrm{r}}$ (right axis) for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$.
+As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP.
+For instance, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, the achievable $\mathrm{KLD}_{\mathrm{r}}$ values for $\{\mathrm{T}_{1},\mathrm{T}_{2},\mathrm{T}_{3}\}$ are approximately $\{38.87,32.77,41.85\}$ bits, with corresponding detection probabilities $\{0.94,0.89,0.97\}$.
+Performance continues to rise at higher $\mathit{P}_{\mathrm{R}}/N_{o}$, achieving near-unity detection across all targets once $\mathit{P}_{\mathrm{R}}/N_{o}$ exceeds 22 dB.
+In Fig.
+4.b, the communication subsystem performance is depicted via the BER (left axis) and $\mathrm{KLD}_{\mathrm{c}}$ (right axis) for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$.
+Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER.
+For example, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, $\mathrm{KLD}_{\mathrm{c}}$ values of $\{77.19,28.13,260.53\}$ bits are observed for $\{\mathrm{UE}_{1},\mathrm{UE}_{2},\mathrm{UE}_{3}\}$, with corresponding BERs of approximately $\{5.1\times 10^{-9},1.7\times 10^{-4},\lesssim 10^{-10}\}$.
+Beyond 14-16 dB, the BER for all UEs drops rapidly– often reaching effectively error-free communication– and $\mathrm{KLD}_{\mathrm{c}}$ grows significantly as the interference from the radar subsystem is jointly minimised.
+Comparing these K-IOP results with the non-optimised ZF baseline (cf.
+Fig.1) or the separate (radar-only or communication-only) waveform approaches in Sec.
+IX-BIX-C, it is evident that the integrated design offers substantially improved performance simultaneously for radar detection and communication reliability.
+This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
+Figure 5: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
 Fig.
-8.a shows the ROC curves for the three targets at different $P_{\mathrm{R}}/N_{o}$ levels.
-At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.217,0.109,0.273\}$ for the non-optimal case.
-These results indicate that communication optimisation enhances detection capabilities at lower $P_{\mathrm{R}}/N_{o}$ compared to non-optimisation, as shown in Fig.
-2.
+5 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for K-ROP, K-COP, and K-IOP optimisation techniques.
+At low $P_{\mathrm{R}}/N_{o}$ values (00-$10$ dB), K-IOP immediately achieves lower MSE than both K-ROP and K-COP.
+As $P_{\mathrm{R}}/N_{o}$ increases beyond 10 dB, the techniques begin to show more distinct performance.
+K-ROP demonstrates faster MSE decay, particularly in the $15$-$20$ dB range, while K-COP exhibits a steadier improvement but converges slightly more slowly.
+The K-IOP continues to outperform both separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values.
+Eventually, all three techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels.
+Notably, this MSE performance strongly correlates with the detection capabilities in Figs.
+3.a, 2.a, and 4, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$.
+Conversely, missed detections or false alarms degrade the estimation by introducing noise-based or zero-valued entries.
+Figure 6: Average CPU time (seconds) versus antenna array size $N$ for both optimisation techniques.
 Fig.
-8.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$.
-The BER decreases with increasing $P_{\mathrm{R}}/N_{o}$, reflecting improved communication performance.
-$\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
-In the non-optimal scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB (Fig.
-1), while with optimisation, $P_{e}$ drops below $10^{-4}$ by 15 dB across all UEs.
-Although optimised waveforms provide a performance gain over the non-optimal method, the gain is less pronounced than with optimised radar waveforms due to radar interference.
-Figure 8: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
-Fig.
-9.a illustrates the trade-off between both subsystems using the optimised communication waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values.
-As $P_{\mathrm{R}}/N_{o}$ increases from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems.
-This highlights the robustness of the optimisation technique in balancing radar and communication performance.
-In contrast, the non-optimal case in Fig.
-3.b shows $\mathrm{KLD_{c}}$ saturating around 11 bits.
-Fig.
-9.b presents the average CPU time for the communication waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$.
-The computational complexity stabilises after $15$ dB as the ISAC system transitions from an infeasible region, where minimum thresholds cannot be met for all targets and UEs, to a feasible region, with average CPU time fluctuating between $0.02$ and $0.04$ seconds.
-Figure 9: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
-VIII Conclusion
+6 presents the average CPU time versus antenna array size $N$, comparing all optimisation techniques.
+The K-ROP consistently demonstrates the lowest computational cost, with CPU times increasing from $0.001$ s at $N=10$ to $7.98$s at $N=100$.
+Meanwhile, K-COP exhibits moderate requirements, ranging from $0.04$ s to $14.07$s across the same $N$.
+By contrast, the K-IOP incurs the highest computational load, growing from $2.44$ s at $N=10$ to $37.69$s at $N=100$.
+For instance, at $N=50$, K-ROP requires only $2.80$s, whereas K-COP needs $3.90$s, and the K-IOP demands $19.96$s.
+These results indicate that while the radar-only and communication-only algorithms remain relatively efficient as $N$ increases, the K-IOP entails modestly higher computational complexity, while offering much higher performance.
+X Conclusion
 This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems.
-The two considered optimisation techniques, namely, KLD-based radar waveform optimisation and KLD-based communication precoding optimisation, demonstrated significant improvements over non-optimised scenarios.
-Radar waveform KLD-based optimisation yielded substantial enhancements in both target detection and communication performance, as evidenced by the ROC curves and reduced BER across all UEs.
-Communication KLD-based precoding optimisation, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs.
-Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimisation demonstrating stable computational efficiency.
-Moreover, trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances the overall ISAC system performance.
-These findings provide valuable insights for designing and optimising shared deployment ISAC systems, offering a promising approach to balance sensing and communication requirements in future wireless networks.
-The proposed framework provides the means for a holistic ISAC design, paving the way for more efficient and versatile wireless communications and sensing technologies.
+The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP for ISAC approach, each demonstrated significant improvements over the non-optimised scenario.
+K-ROP yielded substantial enhancements in both target detection and communication performance while maintaining efficient computational requirements, as evidenced by the numerical results.
+K-COP, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels.
+Both techniques exhibited robust performance across varying SNR values.
+In addition, the K-IOP effectively unifies these advantages, consistently outperforming the two separate approaches in terms of detection probability, BER, and estimation accuracy.
+By jointly designing radar and communication waveforms, the K-IOP approach leverages the synergy between sensing and data transmission, thereby achieving higher overall KLD and a more balanced performance for future ISAC networks.
+Moreover, the numerical results revealed that exploiting the communication signal as part of the radar subsystem significantly boosts ISAC system capabilities, underscoring the importance of holistic design in shared-antenna deployments.
+These findings provide valuable insights for designing and optimising ISAC systems, offering a promising pathway to balance sensing and communication requirements in next-generation wireless networks.
+The proposed framework offers the flexibility to adapt system resources for diverse operating regimes, paving the way for more efficient and versatile wireless communications and sensing technologies.
 Appendix A Gradient Derivations
 This appendix provides detailed derivations for the gradients used in our optimisation problem.
-A-A Gradients for Radar-Waveform Optimisation
 The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27).
 Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$.
-Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, we obtain,
-$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}=\mathbf{R}_{2,t}^{-1}$ (58)
-For the trace term, applying $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ yields,
-$\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}(\sigma_{n}^{2}\mathbf{I}_{N}))}{\partial\mathbf{R}_{2,t}}=-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}$ (59)
-Applying the chain rule to (36), and substituting (58) and (59), we get,
-$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}$ (60)
+Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
+$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.$ (68)
+Applying the chain rule to (27), and substituting (68) we get,
+$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},$ (69)
 where $\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{H}_{t}^{H}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$.
 Substituting this into (69) and simplifying, we obtain,
-$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})$ (61)
+$\!\!\!\!\!\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!$ (70)
 Finally, the gradient of $f$ with respect to $\mathbf{W}_{\mathrm{r},t}$, derived from (31) and (70), is,
-$[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}(\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})_{i}$ (62)
+$\!\!\!\!\!\![\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!$ (71)
 For $\mathrm{KLD}_{\mathrm{c},k}$ gradient calculations, we consider how changes in $\mathbf{W}_{\mathrm{r}}$ affect $\sigma_{\eta}^{2}$, where it is defined as follows,
@@
 The derivative of $\sigma_{\eta}^{2}$ with respect to $\mathbf{W}_{\mathrm{r},l}$ is,
-$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}$ (64)
+$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.$ (73)
 Computing $\partial\mathrm{KLD}_{\mathrm{c},k}/\partial\sigma_{\eta}^{2}$,
-$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}$ (65)
+$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.$ (74)
 Utilising the chain rule and (73), (74), we obtain,
-$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}$ (66)
+$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.$ (75)
 Therefore, the gradient of $\mathrm{KLD}_{\mathrm{c},k}$ with respect to $\mathbf{\dot{W}}_{\mathrm{r}}$ is,
-$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}$ (67)
-Equations (62) and (67) provide the mathematical foundation for the gradients used in our optimization algorithm.
+$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (76)
+Equations (71) and (76) provide the mathematical foundation for the gradients used in our optimisation algorithm.

### v5

@@
 The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms.
-The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby strengthening the trade-off between sensing and communication objectives and achieving higher overall performance than the individual radar- or communication-only techniques.
+The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby synergising sensing and communication objectives and achieving higher overall performance than either radar- or communication-only techniques.
 Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem.
-Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming and identity-covariance radar baselines, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
+Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming, identity-covariance radar baselines, and traditional optimisation approaches, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
 Index Terms:
@@
 Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications.
-ISAC aims to synergistically merge sensing and communication functionalities hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13].
+ISAC aims to synergistically merge sensing and communication functionalities, hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13].
 This integration can be achieved through various approaches.
@@
 This disparity complicates the holistic assessment and optimisation of ISAC systems.
-To this end, some unified performance measures have been introduced in the literature, such as the mutual information [18, 19, 20], and Kullback-Leibler divergence, also known as the relative entropy, which we recently introduced [21, 22, 23, 24].
+To this end, some unified performance measures have been introduced in the literature, such as mutual information [18, 19, 20, 21], and Kullback-Leibler divergence (KLD), also known as the relative entropy, which we recently introduced [22, 23, 24, 25].
+Unlike mutual information approaches that assume capacity-achieving codes and focus on information transmission [26, 27], KLD directly measures distinguishability for practical finite-length systems, offers direct performance insights through Stein’s lemma for hypothesis testing, and captures the impact of practical modulation schemes on system performance, aligning naturally with ISAC’s dual objective.
 I-A Literature review
+While the application of KLD in sensing systems is well-established [19, 28], its potential in characterising communication system performance has recently emerged as a powerful tool for ISAC design.
+For instance, in our previous works [23, 25], we employed KLD to analyse the performance trade-off in ISAC systems with separated antenna deployment, relating radar detection probability $P_{D}$ and communication BER to the achievable KLD.
+This connection stems from the fact that KLD is intrinsically tied to the maximum-likelihood detection, widely used in radar detection and equally central to communication receivers, whose errors are quantified by BER.
+Consequently, KLD can quantify the detection capability at both the radar and communication ends in a similar manner, making it a powerful and unified design tool for ISAC systems.
+While in [22] we proposed a low-complexity unified objective function based on KLD, the aim was focused on optimising network resources for both separated and shared deployment scenarios.
+Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [29].
+Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target ISAC system using the promising KLD framework.
+The fundamental strength of KLD lies in its ability to provide a unified measure that captures the key detection processes in both subsystems.
+Unlike traditional metrics that operate in disparate mathematical spaces—probability of detection for radar and BER for communication, KLD enables direct comparison and joint optimisation.
+Moreover, KLD offers analytical tractability through closed-form gradient expressions, avoiding the computational complexity associated with Q-functions and error function derivatives in traditional formulations.
 Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance.
@@
 Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario.
-Additionally, this work investigates the trade-off between trade-offs between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [22].
+Additionally, this work investigates the synergies between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [23].
 Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model.
@@
   • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together.
-This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, leading to enhanced performance in terms of radar detection accuracy and communication reliability compared with the separate waveform methods.
-By capturing and exploiting the synergy between sensing and data transmission, the unified technique achieves higher overall performance in shared-antenna ISAC deployments.
+This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, yielding enhanced performance in radar detection accuracy and communication reliability compared with the separate waveform methods.
+By capturing and exploiting the synergy between sensing and data transmission, the unified approach achieves superior performance in shared-antenna ISAC deployments.
   • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation.
 This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection, which offers new insights into the underlying synergy between radar and communication waveforms and quantifies how their joint optimisation can enhance overall ISAC capabilities.
-  • Comprehensive Complexity Assessment: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility.
+  • Complexity Evaluation: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility.
 This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
-Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline.
-The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, with modest radar performance gains.
+  • Benchmark Analysis: We compare the proposed KLD-based strategies against the conventional zero-forcing (ZF) beamforming for the communication subsystem and the identity covariance design for the radar subsystem.
+In addition, we introduce two optimisation frameworks based on the BER and probability of detection to further benchmark our methods.
+These comparisons highlight the performance gains, computational complexity trade-offs, and practical relevance of our KLD-based approach in a range of ISAC scenarios.
+Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline, as well as the BER-based and detection-based designs in terms of computational cost.
+The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, yielding modest gains in radar performance.
 Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency.
-The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering even higher overall performance in terms of sensing and communication signals.
-The numerical results highlight how leveraging communication signals for radar detection significantly enhances overall system performance, especially in scenarios with constrained dedicated radar resources.
-These findings emphasise the practical significance of KLD-based designs for shared-antenna ISAC systems, provide important adaptability and insights for balancing sensing and communication requirements in future wireless networks.
+The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering higher overall performance in terms of sensing and communication signals.
+KLD-based formulations also facilitate a structured, scalable, and computationally efficient optimisation framework—unlike complex BER-based and detection-based designs, as evidenced by our numerical results.
+Notably, utilising the communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources.
+These highlight the practical significance of KLD-based waveform design and provide key insights for developing efficient shared deployment ISAC systems that balance sensing and communication requirements.
 I-C Paper organisation
 The paper is structured as follows.
-Section II presents the system model.
-Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system.
-In Section V, the radar waveform optimisation is investigated using the KLD.
-Section VI, the communication waveform optimisation is investigated using the KLD.
-Section VII shows the complexity analysis for both waveform optimisation techniques, Section VIII displays the numerical results, and finally, Section IX concludes the work.
-$Notation$: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors.
+Sec.II presents the system model.
+Sec.III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system.
+In Sec.V, the radar waveform optimisation is investigated.
+Sec.VI, the communication waveform optimisation is investigated.
+Sec.VII, introduces the integrated waveform optimisation.
+Sec.VIII shows the complexity analysis for all waveform optimisation techniques, Sec.IX displays the numerical results, and finally, Sec.X concludes the work.
+NotationNotation: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors.
 Superscripts $(\cdot)$$^{*}$, ${(\cdot)}^{T}$, and ${(\cdot)}^{H}$, denote the conjugate, transpose, and Hermitian transpose, respectively.
@@
 We consider an $N$ antenna MIMO-BS.
-The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner.
+The antennas are utilised for detecting a maximum number of T{T} targets and serving $K$ number of single-antenna communication UEs in the downlink direction in a shared deployment manner, reflecting the emerging paradigm where next-generation wireless infrastructure will inherently support dual-functional operation for both communication and sensing services [38].
 The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties.
 The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$.
-As a starting point, non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16].
+As a starting point, a non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16].
 These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI.
 The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
-$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},$ (1)
+$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},\vskip-3.61371pt$ (1)
 where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem.
-The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS.
+The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets, which reflect the received $\mathbf{x}_{l}$ back to the BS.
 Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target.
@@
 This formulation allows us to incorporate the communication signals into the radar detection process.
-It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
+It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and $\|.\|_{F}$ is the Frobenius norm.
 II-A Communication System
@@
 The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$.
-ZF designs which is typically used to precode the communication data symbols at the BS.
-This precoding scheme will serve as a benchmark for our designs.
-The ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power.
-This approach usually results in excellent performance at high SNR values.
-The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
-$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (3)
-where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
+We assume that channel state information (CSI) is obtained through standard pilot-based estimation procedures widely adopted in modern communication systems with established techniques such as MMSE and ML estimators [39, 40].
 II-B Radar system
-The radar system is designed for high adaptability, allowing for real-time adjustments in detection capabilities across $L$ frame snapshots.
-$T$ represents the radar beams emitted during a specific detection frame.
-Using MIMO radar technology, multiple beams can be generated simultaneously using orthogonal signals [33, 34], where $T$ signifies the upper limit of detectable targets in a single frame.
+The radar system is designed for high adaptability, allowing real-time adjustments in detection capabilities across L\!L\!
+frame snapshots.
+TT\!
+denotes the radar beams emitted during a given detection frame.
+Using MIMO radar technology, multiple beams can be formed simultaneously via orthogonal signals [41, 42], where $T$ is the upper limit of detectable targets in one frame.
 The system explores multiple angular-range-Doppler bins over successive frames, increasing target identification.
@@
 The total radar return signal at the $l$-th snapshot is shown as follows,
-$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vspace{-0.05 in}$ (4)
+$\mathbf{y}_{\mathrm{r},l}=\sum_{t=1}^{T}{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}}+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{n}_{l},\vskip-3.61371pt$ (3)
 where $\mathbf{H}_{t}\in\mathbb{C}^{N\times N}$ is the target response matrix that captures the complete BS-Target-BS path characteristics for the $t$-th target, inherently incorporating propagation effects, target cross-section, and reflection coefficients.
@@
 The received signal for the $t$-th target at the BS after match-filtering, under the binary hypothesis testing problem $\mathcal{H}_{q}\forall q\in\{0,1\}$, where $\mathcal{H}_{1}$ represents the target presence whereas $\mathcal{H}_{0}$ denotes the target absence, is denoted by $\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}$ and can be mathematically represented as follows,
-$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}\>q+\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}\>q+\mathbf{n}_{t,l},$ (5)
+$\mathbf{y}_{\mathrm{r},t,l|\mathcal{H}_{q}}=\overbrace{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{w}_{\mathrm{r},t,l}}^{\begin{subarray}{c}\text{Radar waveform}\\ \text{ radar return}\end{subarray}}\>q+\overbrace{\mathbf{H}_{t}^{T}\>d_{\mathrm{r},t}^{\zeta/2}\>\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}}^{\begin{subarray}{c}\text{Comms.
+waveform}\\ \text{ radar return}\end{subarray}}q+\overbrace{\mathbf{n}_{t,l}}^{\text{Noise}},$ (4)
 where $\mathbf{n}_{t,l}\sim\mathcal{CN}(\mathbf{0},\sigma_{n}^{2}\mathbf{I}_{N})$ is the AWGN vector at the $l$-th snapshot at the output of the matched filter.
 III The communication KLD: ZF and Arbitrary precoding matrix
-In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver.
+In this section, we derive the KLD of the communication system for ZF beamforming with shared deployment antenna configuration, motivated by its direct connection to the maximum likelihood detection process and tractable analytical expressions for system optimisation, where the KLD quantifies the distinguishability between different transmitted symbols by measuring the information divergence of their received signal distributions, capturing the fundamental ability to differentiate communication symbols at the receiver.
 The analysis is carried out for the normalised precoders introduced in Section II-A.
@@
 III-A ZF beamforming
+ZF designs are typically used to precode the communication data symbols at the BS.
+This scheme will serve as a benchmark for our designs.
+ZF eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power.
+This usually results in excellent performance at high SNR values.
+The normalised ZF beamforming matrix is also designed based on the channel matrix, and can be represented as follows,
+$\mathbf{W}_{\mathrm{c}}=\mathbf{P}\mathbf{\tilde{W}}_{\mathrm{c}}/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}$ (6)
+where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.
 With ZF precoding, the received signal at the $k$-th user is,
 $y_{k,l}=\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k}+\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l},$ (7)
-where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of the normalised ZF precoding matrix.
-Note, that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI.
-The distribution of $\eta_{k}$ is well approximated by a complex Gaussian distribution using CLT, particularly for large $N$.
-This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of,
-$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}P_{\mathrm{r}}\sigma_{h}^{2}+\sigma_{n}^{2}$ (8)
-where $\mathbf{R}_{w,t}\!\!=\!\!\frac{1}{L}\!\sum_{l=1}^{L}\!\!\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}\mathbf{\tilde{w}}_{\!\mathrm{r},t,l}^{H}$.
+where $\mathbf{w}_{\mathrm{c},k}$ is the $k$-th column of $\mathbf{W}_{\mathrm{c}}$.
+It is worth noting that the term $\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}=0$ as the ZF precoder eliminates the IUI.
+The radar interference term $\mathbf{h}_{k}^{T}d_{\mathrm{c},k}^{-\zeta/2}\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}$ consists of $NT$ independent random variables arising from the random radar precoding.
+For moderate to large $NT$ values, CLT ensures that this sum converges to a Gaussian distribution.
+Combined with the inherently Gaussian noise $n_{k,l}$, the total interference-plus-noise $\eta_{k}$ is well-approximated as $\mathcal{CN}(0,\sigma_{\eta}^{2})$, particularly for large $N$.
+This approximation yields a mean of $\mathbb{E}[\eta_{k}]=0$, and a variance of
+$\sigma_{\eta}^{2}$ $=d_{\mathrm{c},k}^{-\zeta}\mathbb{E}[\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r},l}\|^{2}]+\sigma_{n}^{2}=d_{\mathrm{c},k}^{-\zeta}\sigma_{h}^{2}\mathrm{Tr}(\mathbf{R}_{w})+\sigma_{n}^{2}$ (8)
 The mean vectors and covariance matrices for the $n$-th and $m$-th symbols are given by,
 $\boldsymbol{\mu}_{k,j}$ $=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]s_{\mathrm{c},k,l}^{(j)}$
-$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\quad j\in\{n,m\},$ (9)
+$=d_{\mathrm{c},k}^{-\zeta/2}\mathbb{E}\left[\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\right]s_{\mathrm{c},k,l}^{(j)},\qquad j\in\{n,m\},$ (9)
 $\ \Sigma_{n}$ $=\Sigma_{m}=\sigma_{\eta}^{2},$ (10)
 where $\mathbb{E}[\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}]=1$ due to the normalisation of the ZF precoding matrix.
-Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter $N-K+1$ and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$.
+Since $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}=\mathrm{tr}((\mathbf{H}\mathbf{H}^{H})^{-1})$, under flat fading channel, $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}$ follows a Gamma distribution with shape parameter NK+1N-K+1 and scale parameter $1$, denoted as $\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{-2}\sim\mathrm{Gamma}(L_{\mathrm{G}},1)$, where $L_{\mathrm{G}}=N-K+1$.
 Let $x\triangleq\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$.
-By substituting (9), and (10) in (6), and considering all possible pairs of dissimilar symbols [22, Corollary 1], then the KLD becomes,
+By substituting (9), and (10) in (5), and considering all possible pairs of dissimilar symbols [23, Corollary 1], then the KLD becomes,
 $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}\>p_{k}}{M(M-1)\sigma_{\eta}^{2}\ln 2}\mathbb{E}\left[\frac{1}{\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}}\right],$ (11)
@@
 The expectation of $1/\|\mathbf{\tilde{W}}_{\mathrm{c}}\|_{F}^{2}$ can be calculated using the PDF of Gamma distribution as follows,
-[…]
+$\begin{split}\mathbb{E}\!\left[\|\mathbf{\tilde{W}}_{\!\mathrm{c}}\|_{F}^{-2}\right]\!\!&=\!\!\!\int_{0}^{\infty}\!\!\frac{1}{x}\cdot\frac{1}{\Gamma(L_{\mathrm{G}})}x^{L_{\mathrm{G}}-1}e^{-x}\!dx\!=\!N\!\!-\!\!K.\end{split}$ (12)
 Substituting this result into (11), we get the final $\mathrm{KLD}_{\mathrm{c},k}^{\mathrm{ZF}}$,
@@
 A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem.
-Considering the received signal representation in (2), the conditional mean vectors and variance are,
-$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\$ (14)
+By noting that the received signal representation $y_{\mathrm{c},k,l}$ in (2) consists of a desired term, (K1)(K-1) statistically independent inter-user symbols and $NT$ independent radar/noise terms, as well as, by considering CLT, the resulting sum is well approximated by a proper complex Gaussian for a typical MIMO-ISAC system which employs large antenna arrays.
+We therefore take $y_{\mathrm{c},k,l}^{(j)}\sim\mathcal{CN}(\mu_{k,j},\Sigma_{\mathrm{c},j}),\ \ j\in\{n,m\}$, where the conditional mean vectors and variance are,
+$\mu_{k,j}=\mathbb{E}[y_{\mathrm{c},k,l}^{(j)}]=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},k}s_{\mathrm{c},k,l}^{(j)},\!\quad j\in\{n,m\},\\$ (14)
 $\operatorname{Var}(y_{\mathrm{c},k,l}^{(j)})$ $=\mathbb{E}[|y_{\mathrm{c},k,l}^{(j)}-\boldsymbol{\mu}_{k,n}|^{2}]$
-$=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\quad j\in\{n,m\},$ (15)
-Note that, the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
+$=\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\!\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2},\!\!\!\qquad j\in\{n,m\}.$ (15)
+Note the variance of the received signal for a given pair of symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
 $\Sigma_{\mathrm{c},n}=\Sigma_{\mathrm{c},m}=d_{\mathrm{c},k}^{-\zeta}\>\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2},$ (16)
@@
 $\!\!\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}\!=\!\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg(\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
-×$\!\!\times\!\Bigg{(}\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg{)}^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg{)}\!.\!\!$ (17)
-Thereafter, by considering all possible pairs of dissimilar symbols, then the KLD expression using $\lambda$ is simplified as,
+$\!\!\times\!\Bigg(\!d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}\Bigg)^{\!\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}\!-\!s_{\mathrm{c},k,l}^{(n)})\!\right)\!\!\Bigg)\!.\!\!$ (17)
+Thereafter, by considering all possible pairs of dissimilar symbols, the KLD expression using $\lambda$ is simplified as,
 $\begin{split}\mathrm{KLD}^{m\rightarrow n}_{\mathrm{c},k}&=\frac{\lambda\>d_{\mathrm{c},k}^{-\zeta}}{2M(M-1)\ln 2}\!\!\times\!\!\frac{\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\|^{2}}{d_{\mathrm{c},k}^{-\zeta}\!\sum_{i=1,i\neq k}^{K}\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}\!+\!\sigma_{\eta}^{2}}.\end{split}$ (18)
-It should be noted that the second term in this equation, represents a form of the SINR, where the interference is the sum of IUI and radar interference.
+Where the interference is the sum of IUI and radar interference.
+It should be noted that the second term in this equation represents a form of the SINR, confirming KLD’s alignment with conventional metrics while uniquely enabling modulation-aware distinguishability and synergistic interference exploitation through its unified hypothesis-testing foundation.
 IV Radar system analysis
 IV-A Composite detection and response matrix estimation
-The main objective of typical radar systems is to detect the existence of targets and estimate the matrix response for existing targets.
+The main objective of radar systems is to detect the existence of targets and estimate the matrix response for existing targets.
 After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
 $\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{q}).$ (19)
-This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining ^$\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if presents.
+This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining $\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if present.
 It can be observed from (4) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
 $f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t,l}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
-×$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
+$\times\exp\left(-\mathbf{y}_{\mathrm{r},t,l}^{H}\left[(1-q)\sigma_{n}^{-2}+q\,\det(\mathbf{R}_{1,t})\right]\mathbf{y}_{\mathrm{r},t,l}\right),$ (20)
 where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$.
@@
 IV-B The KLD of radar system
-[…]
+As the number of snapshots $L$ approaches infinity, the law of large numbers ensures that the estimated target response matrix converges in probability to the true value, i.e., $\hat{\mathbf{H}}_{t}\xrightarrow{p}\mathbf{H}_{t}$.
+This allows us to use $\mathbf{H}_{t}$ in place of $\hat{\mathbf{H}}_{t}$ for large $L$.
+The received radar signal $\mathbf{y}_{r,t,l}$ becomes a linear transformation of the Gaussian noise $\mathbf{n}_{t,l}$ plus deterministic signal components, where we can proceed with a more tractable derivation of the KLD for the radar subsystem.
 Using this property, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (20), and (5) as follows,
@@
 $\mathrm{KLD}_{\mathrm{ISAC}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}+\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}.$ (28)
+This formulation represents the aggregate information-theoretic performance measure across both subsystems, where the averaging within each subsystem ensures balanced treatment regardless of the number of users $K$ or targets $T$, where $\mathrm{KLD}_{\mathrm{ISAC}}$ provides a unified optimisation objective that captures the fundamental goal of maximising information extraction from shared ISAC resources.
 V Radar Waveform Optimisation
-In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
-$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\quad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
+The KLD-based optimisation framework provides analytically tractable KLD expressions, and their gradients as shown in Sec.
+III, Sec.
+IV-B, and Appendix A.
+This computational efficiency, combined with the unified performance characterisation, makes KLD particularly suitable for real-time ISAC applications.
+In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem.
+The optimisation problem is,
+$\mathcal{P}_{1}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\qquad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (29a)
 s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (29b)
@@
 The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively.
-The constraint on radar power (29d) is explicit because the radar waveform is our primary optimisation variable.
+For example, for the radar subsystem, a KLD threshold of $\underline{A}_{t}=10$ bits corresponds to a detection probability of approximately $P_{D}\approx 0.25$, representing a baseline detection capability.
+For the communication subsystem, a KLD threshold of $\underline{B}_{k}=10$ bits ensures a BER of approximately $10^{-2}$ for QPSK modulation, which is acceptable for many practical applications.
+These thresholds can be adjusted based on specific system requirements—higher KLD values would yield improved performance at the cost of more stringent resource allocation.
+The mathematical relationships between KLD and these conventional metrics, established in [23], allow system designers to set appropriate thresholds based on application needs.
+The constraint on radar power (29d) is explicit, where fixed power allocation is used between subsystems, because the radar waveform is our primary optimisation variable.
 The communication power is implicitly considered through the ZF beamforming scheme and the constraint (29c) on communication performance.
@@
 $\begin{split}p(\mathbf{\dot{W}}_{\mathrm{r}})=&\sum_{t=1}^{T}\left(\max(0,\underline{A}_{t}-\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}})^{2}\right)\\ &+\sum_{k=1}^{K}\left(\max(0,\underline{B}_{k}-\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}}})^{2}\right).\end{split}$ (32)
-The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix I.A.
-The final equation is given by,
+The gradient of the objective function, $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})\in\mathbb{C}^{N\times L\times T}$, is a gradient tensor computed element-wise, as detailed in Appendix A The final equation is given by,
 $\begin{split}[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=&\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}\Big(\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ &-\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\Big)_{i}.\end{split}$ (33)
@@
 where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function.
-The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A.
+The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix A.
 The final forms are given by,
@@
 If constraints are consistently violated, the increasing penalty will emphasise constraint adherence.
+V-A Radar waveform detection probability optimisation
+To incorporate the detection performance directly into the optimisation framework, we derive the probability of detection $P_{D,t}$ from the GLRT in Sec.
+IV.A, where $\mathbf{Q}_{t}=\mathbf{I}_{N}+\mathbf{R}_{1,t}^{-1}$.
+For large $L$ and by invoking the CLT, the test statistic converges to a Gaussian distribution under both hypotheses.
+Under $\mathcal{H}_{0}$ where $\mathbf{y}_{\mathrm{r},t,l}=\mathbf{n}_{t,l}$, the mean and variance are $\mu_{0}=\sigma_{n}^{2}\text{tr}(\mathbf{Q}_{t})$ and $\sigma_{0}^{2}=\sigma_{n}^{4}\text{tr}(\mathbf{Q}_{t}^{2})/L$.
+Under $\mathcal{H}_{1}$ where $\mathbf{y}_{\mathrm{r},t,l}=d_{\mathrm{r},t}^{\zeta/2}\mathbf{H}_{t}\mathbf{x}_{t,l}+\mathbf{n}_{t,l}$, we have $\mu_{1}=\mu_{0}+\Delta\mu$ and $\sigma_{1}^{2}=\sigma_{0}^{2}+2\sigma_{n}^{2}\Delta\mu/L$, where $\Delta\mu=d_{\mathrm{r},t}^{\zeta}\text{tr}(\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{Q}_{t})$.
+For a fixed false alarm rate $P_{FA}$, the Neyman-Pearson threshold is,
+$\tau_{2}=\mu_{0}+Q^{-1}(P_{FA})\sigma_{0}.$ (42)
+Substituting (42) into the detection probability expression $P_{D,t}=Q((\tau_{2}-\mu_{1})/\sigma_{1})$ yields [23],
+$P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}=Q\left(\frac{Q^{-1}(P_{FA})\sigma_{0}-\Delta\mu}{\sigma_{1}}\right),$ (43)
+where $Q(\cdot)$ is the complementary cumulative distribution function of the standard normal distribution.
+The detection probability in (43) depends on the waveform design through $\mathbf{R}_{t}$, which affects both $\mathbf{Q}_{t}$ and $\Delta\mu$.
+We form a baseline radar waveform optimisation by using the detection probability directly in the problem defined in (44) below.
+$\mathcal{P}_{3}:\max_{\mathbf{\dot{W}}_{\mathrm{r}}}\qquad\!\!$ $\frac{1}{T}\sum_{t=1}^{T}P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ (44a)
+[…]
+$P_{D,t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}\leq\overline{D}_{t}\,\forall t\in T,$ (44b)
+where the communication subsystem uses ZF beamforming with precoding matrix $\mathbf{W}_{\mathrm{c}}$ and $\overline{D}_{t}$ is the upper limit of the $P_{D}$ of the $t$-th target.
+This $P_{D,t}$ objective depends on the $Q(\cdot)$ function and its inverse, which introduces additional computational complexity in the optimisation process due to its non-elementary nature and the challenges in evaluating its derivatives efficiently.
+This distinction illustrates the enhanced analytical tractability of the proposed KLD-based approach.
+This problem is inherently non-convex and nonlinear due to the $Q$-function and its constraint.
+Using $\nabla_{\mathbf{W}_{\mathrm{r},t}}\,\mathrm{P}_{D,t}$ derived in Appendix B, and the gradients derived previously, we employ the same projected gradient method with penalty as in Algorithm 1, to handle the nonconvexities in (44).
 VI Communication optimisation
@@
 The optimisation problem can be expressed as follows,
-$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (42a)
-s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (42b)
-${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (42c)
-$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (42d)
-$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.$ (42e)
-Note, that the radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow.
+$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\qquad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (45a)
+[…]
+$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}\leq P_{\mathrm{c}},$ (45b)
+$P_{\mathrm{r}}=P_{\mathrm{T}}-P_{\mathrm{c}}.\vskip-3.61371pt$ (45c)
+Similar to the radar waveform optimisation, fixed power allocation is used between subsystems, where radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow.
 To solve this optimisation problem, we must derive the gradients of the objective function and the constraints.
@@
 Using the chain rule and properties of matrix derivatives, the gradients can be expressed as follows,
-$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\quad\quad\quad\quad\forall j$ (44)
-$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ =$=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (45)
+$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\qquad\qquad\qquad\qquad\forall j$ (47)
+$\nabla{\mathbf{w}_{\mathrm{c},j}}\mathrm{KLD}_{\mathrm{c},k}$ $=\begin{cases}C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}}{b_{k}},&\text{if }j=k\\ -C\cdot\frac{2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\cdot a_{k}}{b_{k}^{2}},&\text{if }j\neq k\end{cases}$ (48)
 where $C=\frac{\lambda}{2M(M-1)\ln 2}$, and $b_{k}=d_{\mathrm{c},k}^{\zeta}\sum_{i=1,i\neq k}^{K}\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}$.
@@
 Next, we need to consider the constraints of our optimisation problem. The power constraint is given by,
-$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vspace{-0.05 in}$ (47)
+$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vskip-3.61371pt$ (50)
 The gradient of this constraint is derived as follows,
-$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vspace{-0.05 in}$ (48)
+$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vskip-3.61371pt$ (51)
 The gradient of the radar KLD constraint with respect to $\mathbf{W}_{\mathrm{c}}$ is derived as follows,
-$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vspace{-0.03 in}$ (49)
+$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vskip-2.168pt$ (52)
 Algorithm 2 Interior Point Method for Communication Optimisation
@@
 14:return $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
-Solving the optimisation problem defined in equations (42a)-(42d) uses the IPM, as detailed in Algorithm 2.
+Solving the optimisation problem defined in equations (45a)-(45b) uses the IPM, as detailed in Algorithm 2.
 This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$.
@@
 This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
+VI-A Communication waveform BER-optimisation
+We formulate another baseline design that minimises the instantaneous BER with respect to $\mathbf{W}_{\mathrm{c}}$, while ensuring a minimum radar KLD per target.
+Where the instantaneous BER of user $k$ is approximated by (54).
+The optimisation problem is,
+$\mathcal{P}_{5}:\min_{\mathbf{W}_{\mathrm{c}}}\qquad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{BER}_{k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (53a)
+[…]
+$\mathrm{BER}_{k}\;\leq\;\overline{C}_{k},\forall k\in K,$ (53b)
+where the radar subsystem precoder $\mathbf{W}_{\mathrm{r}}$ uses the identity covariance $\mathbf{R}_{w}=\mathbf{I}_{N}$.
+As such, it is important to define the BER, where for $M$-ary constellations, BER of user $k$ is approximated by [52],
+$\mathrm{BER}_{k}~\approx~\frac{\lambda_{\mathrm{c}}}{M\,\log_{2}M}\,Q\!\Bigl(\sqrt{\beta_{\mathrm{c}}\,\mathrm{SINR}_{k}}\Bigr),$ (54)
+where $\lambda_{\mathrm{c}}$ and $\beta_{\mathrm{c}}$ are constellation-dependent constants (e.g., $\lambda_{\mathrm{c}}=4$, $\beta_{\mathrm{c}}=2$ for QPSK).
+Unlike the unified closed-form gradients in our KLD-based design, this BER objective depends on the $Q(\cdot)$ function in (54), which introduces additional computational complexity in the optimisation process due to its non-elementary nature and the challenges in evaluating its derivatives efficiently.
+This distinction illustrates the enhanced analytical tractability of the proposed KLD-based approach.
+This problem is inherently non-convex and nonlinear due to the $Q$-function in (54) and $\mathrm{KLD}_{\mathrm{r},t}$ constraint.
+Using $\nabla_{\mathbf{W}_{\mathrm{c}}}\,\mathrm{BER}_{k}$ derived in Appendix C, and the constraint gradients derived in (50)-(52), we employ the same gradient-assisted IPM as in Algorithm 2, to handle the nonconvexities in (53), leveraging the reduced dimensionality of $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$.
 VII Integrated Waveform Optimisation for ISAC
@@
 The optimisation problem is expressed as,
-$\mathcal{P}_{1}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\quad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (50a)
-s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t\in T,$ (50b)
-$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k\in K,$ (50c)
-$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (50d)
-$P_{\mathrm{T}}=P_{\mathrm{r}}+P_{\mathrm{c}}.$ (50e)
-The total power constraint in (50d) limits the combined radar and communication power to $P_{\mathrm{T}}$.
+$\mathcal{P}_{6}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\qquad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (55a)
+[…]
+$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (55b)
+$P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}.$ (55c)
+The total power constraint in (55b) limits the combined radar and communication power to $P_{\mathrm{T}}$, and optimises the power allocation between the two subsystems.
 To solve this non-convex problem, the ADMM is employed.
@@
 $+\langle\mathbf{U}_{\mathrm{r}},\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\rangle+\langle\mathbf{U}_{\mathrm{c}},\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\rangle$
-$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),$ (51)
+$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),\vskip-3.61371pt$ (56)
 where $\mathbf{U}_{\mathrm{r}}$ and $\mathbf{U}_{\mathrm{c}}$ are the dual variables (Lagrange multipliers), and $\rho>0$ is the ADMM penalty parameter.
@@
 $\mathbf{W}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{c}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{c}}^{(n)},$ (58)
-$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\quad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (54)
+$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\qquad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (59)
 $\mathbf{U}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{r}}^{(n)}+\rho(\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}),$ (60)
-$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).$ (56)
+$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).\vskip-3.61371pt$ (61)
 The auxiliary variables are updated by solving the following constrained subproblem,
-$\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\quad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (57)
-s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t,$ (58)
-$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k,$ (59)
-$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},$ (60)
+$\vskip-3.61371pt\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\qquad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (62)
+s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\qquad\forall t,$ (63)
+$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\qquad\forall k,$ (64)
+$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},\vskip-3.61371pt$ (65)
 where $p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$ is a penalty term that ensures constraints are satisfied.
@@
 The gradients of $f$ with respect to $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ are computed as,
-$\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (61)
-$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).$ (62)
+$\vskip-3.61371pt\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (66)
+$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).\vskip-3.61371pt$ (67)
 The closed-form expressions for $\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ and $\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ are given by,
-$\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (63)
+$\vskip-3.61371pt\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (68)
 $\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ $=\frac{2\lambda d_{\mathrm{c},k}^{-\zeta}}{M(M-1)\ln 2}\Bigg[\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},k}}{\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}}$
-$\quad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg{]}.$ (64)
+$\qquad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg].\vskip-3.61371pt$ (69)
 After each gradient-based update, the auxiliary variables are projected onto the feasible set defined by the power constraint as, $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right)$, where $\mathcal{P}(\cdot)$ denotes the projection operator ensuring that (65) is satisfied.
@@
 0: Initial variables $\mathbf{W}_{\mathrm{r}}^{(0)}$, $\mathbf{W}_{\mathrm{c}}^{(0)}$, $\mathbf{Z}_{\mathrm{r}}^{(0)}$, $\mathbf{Z}_{\mathrm{c}}^{(0)}$, $\mathbf{U}_{\mathrm{r}}^{(0)}$, $\mathbf{U}_{\mathrm{c}}^{(0)}$, parameters $\rho$, $\alpha_{0}$, $\rho_{0}$, tolerance $\varepsilon$, max iterations $\mathrm{max\_iter}$
-0:  Optimized waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
-1:  Initialize $n=0$
+0: Optimised waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
+1: Initialise $n=0$
 2:while $n<\mathrm{max\_iter}$ do
@@
 The projected gradient method with adaptive penalties for radar waveform optimisation, defined in Algorithm 1, has an estimated per-iteration complexity as [50, 49],
-$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),$ (65)
+$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),\vskip-3.61371pt$ (70)
 where $I$ is the number of iterations, $E$ denotes the cost of evaluating the objective function and its gradients, $n$ represents the number of optimisation variables in the tensor $\dot{\mathbf{W}}_{\mathrm{r}}$, $m$ is the number of constraints, and $k$ is the number of backtracking line search iterations per step.
@@
 For communication waveform optimisation, we utilise a gradient-assisted Interior Point Method as outlined in Algorithm 2. Its per-iteration complexity is [53],
-$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),$ (66)
+$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),\vskip-3.61371pt$ (71)
 where $n$ denotes the number of optimisation variables in the matrix $\mathbf{W}_{\mathrm{c}}$. Again, while the dominant $n^{3}$ term reflects matrix operations.
@@
 The ADMM algorithm for unified optimisation defined in Algorithm 3, has per-iteration complexity as [54],
-$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),$ (67)
+$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),\vskip-3.61371pt$ (72)
 where $n$ represents the combined number of optimisation variables, and $J$ is the number of inner projected gradient iterations per ADMM step.
@@
 The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed.
-A pathloss exponent of $\eta=3$ is considered to model large-scale fading.
+A pathloss exponent of $\zeta=3$ is considered to model large-scale fading.
 The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters.
-The channel variance is fixed at $\sigma_{h}^{2}=1$, and the probability of false alarm $P_{FA}$ is set to $10^{-4}$.
-$\overline{C}_{k}$ is set to $210^{-}$ 2 which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits.
-IX-A Trade-off for the conventional benchmark
+The channel variance is fixed at $\sigma_{h}^{2}=1$, the probability of false alarm $P_{FA}$ is set to $10^{-4}$, $\overline{C}_{k}$ is set to $10^{-}2$ which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits, $N_{o}$ is the noise power spectral density, and $P_{\mathrm{R}}$ is the total received power.
+IX-A Conventional benchmark
 The radar subsystem employs a CIC matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced ZF.
-Figure 1: The ISAC system performance at $P_{\mathrm{r}}=0.5$.
+Figure 1: ISAC performance at $P_{\mathrm{r}}=0.5$.
 a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
 b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
-[…]
+Fig.1 presents the ISAC system performance, analysing both its radar and communication subsystems at $P_{\mathrm{c}}=0.5$.
+The analysis encompasses the radar subsystem’s target detection capabilities and the communication subsystem’s data transmission quality.
+Fig.1.a presents the performance of the radar subsystem, illustrated in terms of the $\mathrm{KLD_{r}}$ and the probability of detection $P_{D}$ against $P_{\mathrm{R}}/N_{o}$ for three potential targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$.
+The curves for all three targets exhibit a consistent upward trend with increasing $P_{\mathrm{R}}/N_{o}$, indicating that higher $P_{\mathrm{R}}/N_{o}$ levels enhance the radar subsystem’s capability to distinguish and detect targets.
+It can be observed from the figure that at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\mathrm{KLD_{\mathrm{r}}}$ values for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are approximately $\{8.43,\,6.77,\,9.1\}$ bits, respectively, while the corresponding detection probabilities $P_{D}$ reach $\{0.197,\,0.143,\,0.221\}$.
+By the time $P_{\mathrm{R}}/N_{o}$ increases to $20$ dB, the $\mathrm{KLD_{\mathrm{r}}}$ improves to around $\{19.87,\,17.59,\,20.71\}$ bits, and $P_{D}$ increases to $\{0.771,\,0.627,\,0.803\}$.
+Throughout these $P_{\mathrm{R}}/N_{o}$ values, $\mathrm{T}_{3}$ remains the most detectable target, consistently demonstrating the highest $\mathrm{KLD_{\mathrm{r}}}$ and $P_{D}$, followed by $\mathrm{T}_{1}$ and then $\mathrm{T}_{2}$.
+The consistent rise in both $\mathrm{KLD_{\mathrm{r}}}$ and $P_{D}$ with increasing $P_{\mathrm{R}}/N_{o}$ demonstrates the strong correlation between the two measures, confirming the validity of utilising $\mathrm{KLD_{\mathrm{r}}}$ as a reliable measure of radar performance, as well as a reliable design tool.
+Higher $\mathrm{KLD_{\mathrm{r}}}$ values are directly translated to improved detection probabilities, reinforcing that $\mathrm{KLD_{\mathrm{r}}}$ effectively captures the system’s ability to distinguish targets.
+Fig.1.b shows the performance of the communication subsystem under the ZF precoding scheme, where the $\mathrm{KLD_{\mathrm{c}}}$ and the BER are plotted versus $P_{\mathrm{R}}/N_{o}$ for three single-antenna UEs.
+It can be seen that, as $P_{\mathrm{R}}/N_{o}$ increases, the KLD for all UEs rises, reflecting enhanced signal distinguishability.
+Correspondingly, the BER decreases, indicating better communication quality.
+Around $10$ dB, however, both an upper bound in KLD and an error floor in BER become evident, highlighting the persistent radar interference as the dominant performance limiting factor once IUI has been largely cancelled by ZF at $P_{\mathrm{c}}=0.5$.
+Thereby, this validates the use of KLD as an effective tool for characterising and designing communication systems since higher KLD values systematically align with lower BER.
+Figure 2: Radar subsystem CIC design performance comparison between communication-assisted and non-assisted configurations at $P_{\mathrm{r}}=\{0.1,0.9\}$: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+Fig.
+2 compares communication-assisted and non-assisted radar configurations for $P_{\mathrm{r}}=0.9$ and $P_{\mathrm{r}}=0.1$ for target $\mathrm{T}_{1}$.
+The assisted configuration exploits communication signals as an additional information source for radar detection, while the non-assisted baseline considers only dedicated radar signals.
+Importantly, the communication subsystem performance remains unaffected by the choice of assisted or non-assisted radar operation, as the exploitation occurs solely at the radar receiver through processing the reflected signals.
+For $P_{\mathrm{r}}=0.1$, the communication-assisted configuration demonstrates substantial performance gains.
+At $P_{\mathrm{R}}/N_{o}=20$ dB, the assisted system achieves $\mathrm{KLD}_{\mathrm{r}}=19.489$ bits compared to $3.842$ bits for the non-assisted case, representing a remarkable improvement.
+The corresponding detection probabilities reach $P_{D}=0.749$ and $0.086$, respectively.
+This significant enhancement stems from the high communication power $P_{\mathrm{c}}=0.9$ providing substantial additional signal energy that, when reflected from target $\mathrm{T}_{1}$, contributes to the detection process.
+At $P_{\mathrm{r}}=0.9$, both configurations show improved performance, with the assisted system maintaining a consistent advantage.
+At $P_{\mathrm{R}}/N_{o}=20$ dB, the assisted configuration achieves $\mathrm{KLD}_{\mathrm{r}}=19.873$ bits versus $17.795$ bits for the non-assisted scenario, yielding a modest but good improvement.
+The detection probabilities are $P_{D}=0.771$ and $0.694$, respectively.
+The reduced relative gain is expected as the dedicated radar signal already provides substantial detection capability, diminishing the proportional contribution from communication signals.
+The consistent performance enhancement across all SNR levels confirms that our communication-assisted approach effectively exploits the inherent synergy between communication and radar signals.
 IX-B Radar Waveform KLD-based Optimisation
@@
 The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
-Figure 2: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using radar waveform optimisation: K-ROP.
-a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
-b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
+Figure 3: ISAC performance at $P_{\mathrm{r}}=0.5$ using two radar waveform optimisations: K-ROP and D-ROP.
+a) Radar subsystem: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-2 illustrates the performance of both subsystems using the KLD-based radar waveform optimisation (K-ROP) technique.
+3 illustrates the performance of both subsystems using both KLD-based radar waveform optimisation (K-ROP) and detection probability radar waveform optimisation (D-ROP) techniques.
 In Fig.
-2.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ on the right axis and detection probability $P_{D}$ on the left axis versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
+3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ and detection probability $P_{D}$ versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
 As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance.
-By comparing the results here with the ZF-based design in Fig.
-1, it can be noticed that there is a significant gain can be achieved using K-ROB.
+The results in this figure also show that a slight improvement can be achieved using D-ROP compared to K-ROP, at the expense of slightly worse communication performance due to increased interference.
+Moreover, by comparing the results here with the ZF-based design in Fig.
+1, it can be observed that a significant gain can be achieved using K-ROP and D-ROP.
 For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig.
-1, demonstrating improved detection capabilities through K-ROP optimisation technique.
+1, while D-ROP achieves $\{97.4,90.7,99.99\}\%$, demonstrating improved detection capabilities by both optimisation techniques with D-ROP showing a slight improvement, at the cost of higher complexity.
 For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS.
 At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+While for D-ROP, $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ approximately $\{12.73,8.7,14.4\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB, increasing to $\{51.2,42.9,55.41\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
 The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system.
+The continued increase in $\mathrm{KLD}_{r}$ even as $P_{D}$ approaches unity demonstrates a key advantage of the KLD.
+While probability metrics saturate at their bounds, making it difficult to distinguish between detection probabilities closer to $1$, the unbounded KLD clearly differentiates these performance levels.
+This sensitivity is crucial for high-reliability applications and provides better numerical properties for optimisation algorithms.
 Fig.
-2.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ on the right axis and BER on the left axis versus $P_{\mathrm{R}}/N_{o}$.
+3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ and BER versus $P_{\mathrm{R}}/N_{o}$.
 For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$.
-At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
-The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB suggests it has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced.
-The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted from Fig.
+At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, respectively, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB.
+D-ROP demonstrates similar trends with slightly lower performance, achieving $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values of $\{49.62,18.07,167.48\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB and $\{496.2,180.84,1674.79\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB for both techniques suggests they have reached their optimisation feasibility points, significantly enhancing the communication performance as the radar interference is effectively reduced.
+The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, while D-ROP reaches this threshold at $14$ dB, both significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted in Fig.
 1.
-These results highlight the intricate relationship between radar detection capabilities and communication system performance.
+These results highlight the intricate relationship between radar detection capabilities and communication performance, with K-ROP demonstrating slightly superior performance over D-ROP.
 IX-C Communication Waveform KLD-based Optimisation
-The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
+The communication waveform KLD-based optimisation is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$.
 The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$.
 The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
-Figure 3: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using communication waveform optimisation: K-COP.
-a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
-b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
+Figure 4: ISAC performance at $P_{\mathrm{r}}=0.5$ using two communication waveform optimisations: K-COP and B-COP.
+a) Radar subsystem: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
+b) Communication subsystem: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
 Fig.
-3 illustrates the performance of both subsystems using the KLD-based communication waveform optimisation (K-COP) technique.
+4 illustrates the performance of both subsystems using both KLD-based communication waveform optimisation (K-COP) and BER-based communication waveform optimisation (B-COP) techniques.
 In Fig.
-3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ (right axis) and detection probability $P_{D}$ (left axis) versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
+4.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ and detection probability $P_{D}$ versus $P_{\mathrm{R}}/N_{o}$ are shown for each target.
 For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$.
 At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+B-COP shows slightly lower $\mathrm{KLD_{\mathrm{r}}}$ performance, with values $\{9.31,7.48,10.04\}$ bits at $P_{\mathrm{R}}/N_{o}=10$ dB, rising to $\{21.93,19.42,22.86\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
 The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform.
-At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig.
+At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while B-COP achieves $\{68.5,58.2,72.3\}\%$, compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig.
 1.
-K-COP demonstrates improved detection capabilities over non-optimisation but remains below the radar waveform optimisation performance achieved using K-ROP in Fig.
-2, which is $\{98,89.2,99.9\}\%$.
+Both techniques demonstrate improved detection capabilities over the non-optimised scenario but remain below the radar waveform optimisation performance achieved using K-ROP in Fig.
+3.
 Fig.
-3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ (right axis) and BER (left axis) versus $P_{\mathrm{R}}/N_{o}$.
+4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ and BER versus $P_{\mathrm{R}}/N_{o}$.
 For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$.
 For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB.
+Interestingly, B-COP demonstrates marginally better communication performance.
+The BER performance reflects this trend, with B-COP achieving slightly better performance than K-COP.
 In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig.
 1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs.
-K-COP optimisation technique provides significant performance gains over the non-optimised ZF scenario, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec.
-IX-B due to radar interference.
-The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
+B-COP achieves this threshold marginally earlier at $13.5$ dB.
+Both optimisation techniques provide significant performance gains over the non-optimised ZF, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec.
+IX-B due to radar interference, where K-COP’s fixed omnidirectional radar CIC matrix design creates an interference ceiling that limits achievable communication performance regardless of precoding sophistication.
+The communication performance shows less improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
 IX-D Integrated Waveform KLD-based Optimisation for ISAC
@@
 The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
-Figure 4: The ISAC system performance using the integrated waveform optimisation.
+Figure 5: ISAC performance using the integrated waveform optimisation.
 a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$.
@@
 In Fig.
-4.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ (left axis) and $\mathrm{KLD}_{\mathrm{r}}$ (right axis) for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$.
+5.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$.
 As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP.
@@
 In Fig.
-4.b, the communication subsystem performance is depicted via the BER (left axis) and $\mathrm{KLD}_{\mathrm{c}}$ (right axis) for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$.
+5.b, the communication subsystem performance is depicted via the BER and $\mathrm{KLD}_{\mathrm{c}}$ for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$.
 Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER.
@@
 This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
-Figure 5: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
+Figure 6: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$.
 Fig.
-5 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for K-ROP, K-COP, and K-IOP optimisation techniques.
-At low $P_{\mathrm{R}}/N_{o}$ values (00-$10$ dB), K-IOP immediately achieves lower MSE than both K-ROP and K-COP.
+6 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
+At low $P_{\mathrm{R}}/N_{o}$ values (0-$10$ dB), K-IOP immediately achieves lower MSE than all of the other techniques.
 As $P_{\mathrm{R}}/N_{o}$ increases beyond $10$ dB, the techniques begin to show more distinct performance.
-K-ROP demonstrates faster MSE decay, particularly in the $15$-$20$ dB range, while K-COP exhibits a steadier improvement but converges slightly more slowly.
-The K-IOP continues to outperform both separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values.
-Eventually, all three techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels.
-Notably, this MSE performance strongly correlates with the detection capabilities in Figs.
-3.a, 2.a, and 4, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$.
-Conversely, missed detections or false alarms degrade the estimation by introducing noise-based or zero-valued entries.
-Figure 6: Average CPU time (seconds) versus antenna array size $N$ for both optimisation techniques.
+K-ROP and D-ROP demonstrate faster MSE decay, particularly in the 1515-2020 dB range, while K-COP and B-COP exhibit a steadier improvement but converge slightly slower.
+The K-IOP continues to outperform all separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values.
+Eventually, all techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels.
+This MSE performance strongly correlates with the detection capabilities in Figs.
+4.a, 3.a, and 5, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$.
+Conversely, missed detections or false alarms degrade the estimation by introducing noise-based estimates.
+[…]
 Fig.
-6 presents the average CPU time versus antenna array size $N$, comparing all optimisation techniques.
-The K-ROP consistently demonstrates the lowest computational cost, with CPU times increasing from $0.001$ s at $N=10$ to $7.98$s at $N=100$.
-Meanwhile, K-COP exhibits moderate requirements, ranging from $0.04$ s to $14.07$s across the same $N$.
-By contrast, the K-IOP incurs the highest computational load, growing from $2.44$ s at $N=10$ to $37.69$s at $N=100$.
-For instance, at $N=50$, K-ROP requires only $2.80$s, whereas K-COP needs $3.90$s, and the K-IOP demands $19.96$s.
-These results indicate that while the radar-only and communication-only algorithms remain relatively efficient as $N$ increases, the K-IOP entails modestly higher computational complexity, while offering much higher performance.
+7.a presents the average CPU time versus antenna array size $N$ for all optimisation techniques.
+The K-COP demonstrates the lowest computational cost, increasing from $0.04$s at $N=10$ to $7.98$s at $N=100$.
+On the other hand, K-ROP exhibits moderate requirements, ranging from $0.001$s to $14.07$s across the same $N$ range, and D-ROP shows higher CPU time growth, reaching $21.63$s at $N=100$.
+The communication-oriented B-COP requires $32.35$s at $N=100$, while K-IOP incurs the highest computational load at $37.69$s.
+These results confirm that while the radar-only optimisations maintain relatively low complexity, the joint K-IOP technique and B-COP require additional processing time, with K-IOP achieving superior ISAC performance despite its higher computational cost.
+Fig.
+7.b illustrates the convergence behaviour of the $\mathrm{KLD_{ISAC}}$ metric versus the iteration count at $P_{\mathrm{R}}/N_{o}=18$ dB.
+As can be observed from the figure, K-IOP demonstrates the most rapid convergence, achieving substantial gains within the first $100$ iterations and reaching $74\%$ of its final value by iteration $158$.
+K-ROP also shows efficient convergence, stabilising at $95\%$ of its final value by iteration $135$, and D-ROP achieves stabilisation at $95\%$ of its final value by iteration $320$.
+The communication-focused technique K-COP converges quickly but saturates early, with K-COP reaching $97\%$ of its final value by iteration $90$, while B-COP exhibits a more gradual convergence pattern, requiring approximately $400$ iterations to reach its steady-state value.
+This analysis reveals that K-IOP efficiently balances rapid convergence with joint radar-communication optimisation, making it particularly suitable for real-time ISAC applications.
+Figure 8: Achievable KLD under estimation errors for KLD-based optimisations, with $\sigma_{e}^{2}\!\in\!\{10^{-6},10^{-3},10^{-1}\}$.
+a) $\mathrm{KLD_{r}}$ $P_{\mathrm{R}}/N_{o}$.
+b) $\mathrm{KLD_{c}}$ vs $P_{\mathrm{R}}/N_{o}$.
+Fig.
+8 illustrates the effect of imperfect channel and radar target-response estimates on the KLD obtained by the proposed optimisation techniques.
+For clarity, this is shown for $\mathrm{UE}_{1}$, and $\mathrm{T}_{1}$.
+Fig.
+8.a, presents $\mathrm{KLD_{r}}$ versus $P_{\mathrm{R}}/N_{o}$, and Fig.
+8.b presents $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for three error-variance levels.
+For each optimisation technique, the quasi-perfect case ($\sigma_{e}^{2}=10^{-6}$) yields the highest KLD; increasing the error variance to $10^{-3}$ noticeably flattens the curves, and at $10^{-1}$ the gains from all schemes almost disappear.
+As the error variance rises from $10^{-6}$ to $10^{-1}$, both the radar and communication KLD curves flatten, producing a clear error floor at high $P_{\mathrm{R}}/N_{o}$.
+The observed degradation is therefore a predictable consequence of reduced CSI fidelity, underlining the importance of accurate estimation.
 X Conclusion
 This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems.
-The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP for ISAC approach, each demonstrated significant improvements over the non-optimised scenario.
-K-ROP yielded substantial enhancements in both target detection and communication performance while maintaining efficient computational requirements, as evidenced by the numerical results.
-K-COP, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels.
-Both techniques exhibited robust performance across varying SNR values.
-In addition, the K-IOP effectively unifies these advantages, consistently outperforming the two separate approaches in terms of detection probability, BER, and estimation accuracy.
-By jointly designing radar and communication waveforms, the K-IOP approach leverages the synergy between sensing and data transmission, thereby achieving higher overall KLD and a more balanced performance for future ISAC networks.
-Moreover, the numerical results revealed that exploiting the communication signal as part of the radar subsystem significantly boosts ISAC system capabilities, underscoring the importance of holistic design in shared-antenna deployments.
-These findings provide valuable insights for designing and optimising ISAC systems, offering a promising pathway to balance sensing and communication requirements in next-generation wireless networks.
-The proposed framework offers the flexibility to adapt system resources for diverse operating regimes, paving the way for more efficient and versatile wireless communications and sensing technologies.
-Appendix A Gradient Derivations
-This appendix provides detailed derivations for the gradients used in our optimisation problem.
+The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP, each demonstrated significant improvements over both the non-optimised scenario and optimisation benchmarks (D-ROP and B-COP).
+K-ROP and D-ROP yielded substantial enhancements in target detection and communication performance while K-ROP maintained superior computational efficiency compared to D-ROP, validating the analytical tractability advantages of KLD over traditional Q-function-based BER formulations.
+K-COP showed notable improvements in the communication subsystem with modest radar gains, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels, and demonstrated more efficient optimisation than B-COP due to KLD’s tractable gradient expressions.
+The K-IOP approach effectively unified these advantages, consistently outperforming all separate approaches and benchmarks in detection probability, BER, and estimation accuracy.
+By jointly designing radar and communication waveforms, K-IOP leverages the synergy between sensing and data transmission to achieve higher overall KLD and more balanced performance.
+The numerical results revealed that exploiting communication signals within the radar subsystem significantly boosts ISAC capabilities, underscoring the importance of holistic design in shared-antenna deployments.
+These findings provide valuable insights for next-generation ISAC systems, offering a promising framework that adapts system resources for diverse operating regimes while paving the way for more efficient and versatile wireless communications and sensing technologies.
+Future work will explore the synergistic interplay between KLD and machine learning in ISAC, extend the framework to emerging architectures including cell-free MIMO and full-duplex systems, and develop practical interference exploitation strategies.
+This appendix provides detailed derivations for the gradients used in the optimisation problems.
+-A Gradients for Radar-Waveform Optimisation
 The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27).
@@
 Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
-$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.$ (68)
+$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.\vskip-3.61371pt$ (73)
 Applying the chain rule to (27), and substituting (73) we get,
-$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},$ (69)
+$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},\vskip-3.61371pt$ (74)
 where $\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{H}_{t}^{H}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}$.
 Substituting this into (74) and simplifying, we obtain,
-$\!\!\!\!\!\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!$ (70)
+$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!\vskip-3.61371pt$ (75)
 Finally, the gradient of $f$ with respect to $\mathbf{W}_{\mathrm{r},t}$, derived from (31) and (75), is,
-$\!\!\!\!\!\![\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!$ (71)
+$\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!\vskip-3.61371pt$ (76)
 For $\mathrm{KLD}_{\mathrm{c},k}$ gradient calculations, we consider how changes in $\mathbf{W}_{\mathrm{r}}$ affect $\sigma_{\eta}^{2}$, where it is defined as follows,
-$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}$ (72)
+$\sigma_{\eta}^{2}=d_{\mathrm{c},k}^{\zeta}\>\sigma_{h}^{2}\mathrm{Tr}(\frac{1}{L}\sum_{l=1}^{L}{\mathbf{W}_{\mathrm{r},l}\mathbf{W}_{\mathrm{r},l}^{H}})+\sigma_{n}^{2}\vskip-3.61371pt$ (77)
 The derivative of $\sigma_{\eta}^{2}$ with respect to $\mathbf{W}_{\mathrm{r},l}$ is,
-$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.$ (73)
+$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.\vskip-3.61371pt$ (78)
 Computing $\partial\mathrm{KLD}_{\mathrm{c},k}/\partial\sigma_{\eta}^{2}$,
-$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.$ (74)
+$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.\vskip-3.61371pt$ (79)
 Utilising the chain rule and (78), (79), we obtain,
-$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.$ (75)
+$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.\vskip-3.61371pt$ (80)
 Therefore, the gradient of $\mathrm{KLD}_{\mathrm{c},k}$ with respect to $\mathbf{\dot{W}}_{\mathrm{r}}$ is,
-$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (76)
+$\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.\vskip-4.33601pt$ (81)
 Equations (76) and (81) provide the mathematical foundation for the gradients used in our optimisation algorithm.
+-B Gradient of $P_{D,t}$ w.r.t. $\mathbf{W}_{\mathrm{r}}$
+To enable gradient-based optimisation of $P_{D,t}$, we derive $\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}$.
+Using the chain rule with $P_{D,t}=Q(z_{t})$ where $z_{t}\!=\!(\!Q^{-1}(P_{FA})\sigma_{0}\!-\!\Delta\mu\!)/\!\sigma_{1}$ from (43), and $\phi(\!z)\!=\!\frac{1}{\sqrt{\!2\pi}}e^{\!-z^{2}/2}$ being the standard normal PDF, we have,
+$\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}=-\phi(z_{t})\nabla_{\mathbf{W}_{\mathrm{r},t}}z_{t}.\vskip-4.33601pt$ (82)
+where $\nabla_{\!\mathbf{W}_{\mathrm{r},t}}z_{t}\!\!=\!\frac{1}{\sigma_{1}}\left(\!Q^{-1}\!(P_{FA})\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\sigma_{0}\!-\!\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\Delta\mu\!-\!z_{t}\nabla_{\!\mathbf{W}_{\mathrm{r},t}}\!\sigma_{1}\!\right)$.
+Using the matrix derivative identity $\nabla\!\mathbf{A}^{\!-1}\!=\!-\mathbf{A}^{\!-1}\!(\!\nabla\!\mathbf{A}\!)\mathbf{A}^{\!-1}\!$ obtaining $\mathbf{Q}_{t}$ gradient,
+$\nabla_{\mathbf{W}_{\mathrm{r},t}}\mathbf{Q}_{t}\!=\!-\frac{2d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\mathrm{r},t}\!\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}\!\mathbf{R}_{1,t}^{-1}.\vskip-4.33601pt$ (83)
+The gradient of $\Delta\mu\!$ using the cyclic property of trace and (83) is,
+$\nabla_{\mathbf{W}_{\mathrm{r},t}}\Delta\mu\!=\!\frac{2d_{\mathrm{r},t}^{\zeta}}{L}\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}[\mathbf{Q}_{t}\!-\!d_{\mathrm{r},t}^{\zeta}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{R}_{1,t}^{-1}]\mathbf{H}_{t}.\vskip-3.61371pt$ (84)
+For $\sigma_{0}$ and $\sigma_{1}$, the gradients follow a similar form. The final gradient can be expressed in matrix form as,
+$\nabla_{\mathbf{W}_{\mathrm{r},t}}P_{D,t}=-\frac{2d_{\mathrm{r},t}^{\zeta}\phi(z_{t})}{L\sigma_{1}}\mathbf{W}_{\mathrm{r},t}^{H}\mathbf{H}_{t}^{H}\mathbf{M}_{t}\mathbf{H}_{t},\vskip-4.33601pt$ (85)
+where $\mathbf{M}_{t}=\mathbf{Q}_{t}-d_{\mathrm{r},t}^{\zeta}\mathbf{R}_{1,t}^{-1}\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}\mathbf{R}_{1,t}^{-1}+\alpha_{t}\mathbf{Q}_{t}\mathbf{R}_{1,t}^{-1}\mathbf{Q}_{t}\mathbf{R}_{1,t}^{-1}$ with $\alpha_{t}\!=\!(Q^{-1}(\!P_{FA}\!)\sigma_{n}^{2})/(2\sigma_{0}\sigma_{1}\!\sqrt{L})\!-\!z_{t}\sigma_{n}^{2}/\!\sigma_{0}\!\sqrt{L}$.
+-C Gradients of BER w.r.t $\mathbf{W}_{\mathrm{c}}$
+Using the chain rule, the gradient of $\mathrm{BER}_{k}$ w.r.t. $\mathbf{W}_{\mathrm{c}}$ is:
+$\vskip-3.61371pt\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{BER}_{k}=\frac{\partial\mathrm{BER}_{k}}{\partial\mathrm{SINR}_{k}}\cdot\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k},\vskip-4.33601pt$ (86)
+where the partial derivative is,
+$\frac{\partial\mathrm{BER}_{k}}{\partial\mathrm{SINR}_{k}}=-\frac{\lambda_{\mathrm{c}}\sqrt{\beta_{\mathrm{c}}}}{2M\log_{2}M\sqrt{2\pi\mathrm{SINR}_{k}}}e^{-\beta_{c}\mathrm{SINR}_{k}/2},\vskip-4.33601pt$ (87)
+and $\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k}$ is derived.
+Let $\mathcal{D}=d_{\mathrm{c},k}^{-\zeta}\sum_{i\neq k}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}|^{2}+d_{\mathrm{c},k}^{-\zeta}\|\mathbf{h}_{k}^{T}\mathbf{W}_{\mathrm{r}}\|^{2}+\sigma_{n}^{2}$.
+Then,
+$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{SINR}_{k}$ $=\frac{2d_{\mathrm{c},k}^{-\zeta}}{\mathcal{D}}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}\mathbf{e}_{k}^{T}$
+$-\!\frac{2d_{\mathrm{c},k}^{-2\zeta}}{\mathcal{D}^{2}}\|\mathbf{h}_{k}^{T}\!\mathbf{w}_{\mathrm{c},k}\|^{2}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{W}_{\!\mathrm{c}}(\mathbf{I}\!-\!\mathbf{e}_{k}\mathbf{e}_{k}^{T}),\vskip-4.33601pt$ (88)
+where $\mathbf{e}_{k}$ $k$-th standard basis vector. Substituting (87) and (88) into (86), we get the gradient $\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{BER}_{k}$.